### v1

# Quantum accessible information and classical entropy inequalities

Abstract
Computing accessible information for an ensemble of quantum states is a basic problem in quantum information theory. The recently obtained optimality criterion [7] when applied to specific ensembles of states leads to nontrivial tight entropy inequalities that are discrete relatives of the famous log-Sobolev inequality. In this light, the hypothesis of globally information-optimal measurement for an ensemble of equiangular equiprobable states [2] is reconsidered and the corresponding entropy inequalities are proposed. We prove these inequalities in the cases of state ensembles corresponding to acute or flat pyramids, thus providing the first proof of the hypotheses concerning globally information-optimal observables put forward in the paper [2].
1
Introduction
The problem of accessible information – the maximal information that can be extracted from measurements over an ensemble of quantum states – is a basic problem in quantum information and communication theory, coming back to its origin, and important in particular for applications in quantum cryptography, see e.g. the survey [14]. The problem was solved in some special (notably, symmetric) cases but in general remains open. In this paper we describe a general approach following [7] which gives a criterion for optimality and ties up the problem with entropy inequalities – the tight lower bounds for the classical Shannon information. It turns out that the optimality criterion [7] when applied to specific ensembles of states leads to new nontrivial entropy inequalities that are discrete “distant relatives” of the famous log-Sobolev inequality. In this light, the hypothesis of an information-optimal measurement for an ensemble of equiangular equiprobable states – quantum pyramids formulated and numerically substantiated by Englert and Řeháček in [2] – is reconsidered in the present paper. In the case of acute pyramids the corresponding parametric family of entropy inequalities (23) is proposed in the section 4.1. The inequalities suggest the new tight lower bounds for the Shannon entropy of a discrete probability distribution as compared e.g. to the known bound from [11], theorem 3.11.
The proof of the inequality (23), given in section 5, was obtained by the second author as a response to the conjecture put forward by the first author. Via the optimality criterion of section 2, it also implies the validity of a hypothesis concerning accessible information for state ensembles of acute quantum pyramids, formulated and numerically substantiated in [2] (corollary 1 to theorem 2).
In the section 4.4 the case of obtuse pyramids is considered to get further insight and to formulate the hypotheses concerning the corresponding entropy inequalities (60), (64). Our approach allows also to derive the nonlinear equations (50) (resp. (62)) serving to determine the critical values of the parameter $t$ of the optimal observable corresponding to transition from moderately to strongly acute (resp, obtuse) pyramids.
In the section 4.5 we derive and prove the tight entropy inequality (66) implying the global optimality of the hypothetical observable from [2] (proposition 2 and corollary 2). The findings in sections 4.4, 4.5 also emphasize that for obtuse and flat pyramids in the dimensions $m\geq 7$ the sharp “square root” observable (SRM) outperforms the “unambiguous discrimination” observable (UDM) and is the optimal one in the flat case (tentatively also in the obtuse case).
2
Accessible information: a criterion for optimality
Let $\mathcal{H}$ be a finite dimensional Hilbert space, $\mathfrak{S}=\mathfrak{S}(\mathcal{H})$ the convex set of quantum states (density operators on $\mathcal{H}$).  Ensemble of quantum states is a collection =$\mathcal{E}=\left\{\pi_{j},\rho_{j}\right\}_{j=1,\dots,m}$ where $\rho_{j}\in\mathfrak{S,}$ and $\pi=\left\{\pi_{j}\right\}$ is a probability distribution. Let =$\mathcal{M}=\left\{M_{k}\right\}_{k=1,\dots,n}$ be an observable i.e. a collection of operators on $\mathcal{H}$ such that $M_{k}\geq 0,\,\sum_{k=1}^{n}M_{k}=I$ (unit operator). The joint distribution of “input” $j$ and “output” $k$ is then
$p_{jk}=\pi_{j}\mathrm{Tr\,}\,\rho_{j}M_{k}$
and the Shannon mutual information between $j$ and $k$ is (throughout the paper $\log$ denotes the binary logarithm)
$I(\mathcal{E},\mathcal{M})=\sum_{j,k}p_{jk}\log\frac{p_{jk}}{\pi_{j}p_{k}},$
with
$p_{k}=\sum_{j}p_{jk}=\mathrm{Tr\,}\,\overline{\rho}M_{k},$
where $\overline{\rho}=\sum_{j}\pi_{j}\rho_{j}$ is the average state of the ensemble $\mathcal{E}.$
Accessible information of the ensemble is defined as
$A(\mathcal{E})=\sup_{\mathcal{M}}I(\mathcal{E},\mathcal{M}).$
It was shown [1] that the supremum is attained on an observable of the form $M_{k}=\left|\varphi_{k}\rangle\langle\varphi_{k}\right|,\,k=1,\dots,n$ with $d\leq n\leq d^{2}$ and linearly independent $M_{k}$. In particular, the supremum can be replaced with the maximum.
Introducing the dual ensemble =$\mathcal{E}^{\prime}=\left\{p_{k},\sigma_{k}\right\}$ with
$\sigma_{k}=p_{k}^{-1}\overline{\rho}^{1/2}M_{k}\overline{\rho}^{1/2}=\left|\phi_{k}\rangle\langle\phi_{k}\right|,$
where
$|\phi_{k}\rangle=p_{k}^{-1/2}\overline{\rho}^{1/2}|\varphi_{k}\rangle,$
and observable $\mathcal{M}^{\prime}$ with the components [^1] In the case $\overline{\rho}$ is degenerate, one should use generalized inverse in the following formula.
$M_{j}^{\prime}=\overline{\rho}^{-1/2}\left(\pi_{j}\rho_{j}\right)\overline{\rho}^{-1/2}$
we have the same average state
$\overline{\sigma}=\sum_{k}p_{k}\sigma_{k}=\overline{\rho}$
and the new expression for accessible information:
$A(\mathcal{E})=\max_{\mathcal{E}^{\prime}:\overline{\sigma}=\overline{\rho}}I(\mathcal{E}^{\prime},\mathcal{M}^{\prime})$ (1)
with
$I(\mathcal{E}^{\prime},\mathcal{M}^{\prime})$ $=$ $I(\mathcal{E},\mathcal{M})=-\sum_{j}\pi_{j}\log\pi_{j}+\sum_{k,j}p_{k}p(j|k)\log p(j|k)$
$=$ $H(\pi)-\sum_{k}\mathrm{Tr\,}\,\left(p_{k}\sigma_{k}\right)K\left(\sigma_{k}\right),$
where $p(j|k)=\mathrm{Tr\,}\,\sigma_{k}M_{j}^{\prime},$
$K\left(\sigma\right)=-\sum_{j}M_{j}^{\prime}\log\mathrm{Tr\,}\,\sigma M_{j}^{\prime}.$ (2)
Therefore
$A(\mathcal{E})=H(\pi)-\min_{\mathcal{E}^{\prime}:\overline{\sigma}=\overline{\rho}}\sum_{k}\mathrm{Tr\,}\,\left(p_{k}\sigma_{k}\right)K\left(\sigma_{k}\right)$ (3)
Remark 1.
On the other hand, (1) is also the value of $\overline{\rho}-$constrained classical capacity $C(\mathcal{M}^{\prime},\overline{\rho})$ of the measurement channel corresponding to the observable $\mathcal{M}^{\prime}$ [8]. If $H$ is the Hamiltonian and $E$ is the energy limit then the energy-constrained capacity is
$C(\mathcal{M}^{\prime},E)=\max_{\overline{\rho}:\mathrm{Tr\,}\overline{\rho}H\leq E}C(\mathcal{M}^{\prime},\overline{\rho})$
and the full classical capacity is just $C(\mathcal{M}^{\prime})=\max_{\overline{\rho}}C(\mathcal{M}^{\prime},\overline{\rho}).$
Theorem 1.
The minimization problem
$\min_{\mathcal{E}^{\prime}:\overline{\sigma}=\overline{\rho}}\sum_{k}\mathrm{Tr\,}\,\left(p_{k}\sigma_{k}\right)K\left(\sigma_{k}\right)$ (4)
has the dual problem
$\max\left\{\mathrm{Tr\,}\overline{\rho}\Lambda:\Lambda^{\ast}=\Lambda,\,\Lambda\leq K(\sigma)\text{ for all }\sigma\in\mathfrak{S}\right\}.$ (5)
The following statements are equivalent:
Ensemble $\mathcal{E}_{0}^{\prime}=\left\{p_{k}^{0},\sigma_{k}^{0}\right\}$ with $\sigma_{k}^{0}=\left|\phi_{k}^{0}\rangle\langle\phi_{k}^{0}\right|$ is the solution of the problem (4); $\Lambda_{0}$ is the solution of the problem (5);
a.
$\Lambda_{0}\leq K(\sigma)$ for all $\sigma\in\mathfrak{S};$
b.
$\left[K(\sigma)-\Lambda_{0}\right]\,|\phi_{k}^{0}\rangle=0,\,k=1,\dots,n$.
a.
The entropy inequality
$-\sum_{j}\left\langle\psi|M_{j}^{\prime}|\psi\right\rangle\log\left\langle\psi|M_{j}^{\prime}|\psi\right\rangle\mathrm{\,}\,\geq\left\langle\psi|\Lambda_{0}|\psi\right\rangle,$ (6)
holds for all unit vectors $\psi\in\mathcal{H};$
b.
The unit vectors $|\phi_{k}^{0}\rangle$ turn (6) into equality:
$-\sum_{j}\left\langle\phi_{k}^{0}|M_{j}^{\prime}|\phi_{k}^{0}\right\rangle\log\left\langle\phi_{k}^{0}|M_{j}^{\prime}|\phi_{k}^{0}\right\rangle=\left\langle\phi_{k}^{0}|\Lambda_{0}|\phi_{k}^{0}\right\rangle,\quad k=1,\dots.n.$ (7)
The value of the accessible information (3) is
$A(\mathcal{E})=H(\pi)-\mathrm{Tr\,}\overline{\rho}\Lambda_{0}.$
Proof.
We introduce the POVM on $\mathfrak{S}=\mathfrak{S}(\mathcal{H})$ as
$\Pi(d\sigma)=\sum_{k}\left(p_{k}\sigma_{k}\right)\delta_{\sigma_{k}}(d\sigma),$
where $\delta_{\sigma_{k}}(d\sigma)$ is the $\delta-$measure concentrated in the point $\sigma_{k}\in\mathfrak{S.}$ then the problem (4) can be written as [^2] We refer to [7] for rigorous general definition of the traced integral and formulations of the optimality conditions.
$\mathrm{Tr}\int_{\mathfrak{S}}\,K(\sigma)\Pi(d\sigma)$$\longrightarrow$ $\min$ (8)
[…]
$\Pi(\mathfrak{S})$ $=$ $\overline{\rho},$
which is similar to the Bayes estimation problem [6].This is a convex programming problem for which the dual problem is (5) and the optimality conditions are (ii.a) and
$\int_{B}\left[K(\sigma)-\Lambda_{0}\right]\Pi_{0}(d\sigma)=0,\quad B\subset\mathfrak{S}(\mathcal{H}),$
see [7] for detail. In terms of the optimal ensemble $\mathcal{E}_{0}^{\prime}=\left\{p_{k}^{0},\sigma_{k}^{0}\right\}$ the last condition can be rewritten as
$\left[K(\sigma_{k}^{0})-\Lambda_{0}\right]\,\sigma_{k}^{0}=0,\quad(p_{k}^{0}>0).$ (9)
Note that by summing over $k$ we obtain
$\sum_{k}K(\sigma_{k}^{0})\,p_{k}^{0}\sigma_{k}^{0}\,=\Lambda_{0}\overline{\rho},$ (10)
where $\Lambda_{0}$ must be Hermitean operator.
By taking into account that $\sigma_{k}^{0}=\left|\phi_{k}^{0}\rangle\langle\phi_{k}^{0}\right|,$ we can rewrite (9) as
$\left[K(\sigma_{k}^{0})-\Lambda_{0}\right]|\phi_{k}^{0}\rangle=0$ (11)
and (10) as
$\sum_{k}K(\sigma_{k}^{0})\,p_{k}^{0}\left|\phi_{k}^{0}\rangle\langle\phi_{k}^{0}\right|=\Lambda_{0}\overline{\rho}.$ (12)
Taking into account (2), the condition (ii.a) becomes
$-\sum_{j}\left\langle\psi|M_{j}^{\prime}|\psi\right\rangle\log\mathrm{Tr\,}\,\sigma M_{j}^{\prime}\geq\left\langle\psi|\Lambda_{0}|\psi\right\rangle,\quad\psi\in\mathcal{H},$
for all $\sigma\in\mathfrak{S.}$ Without loss of generality we can assume that $\left\|\psi\right\|=1$ here. Then for fixed $\psi$ the minimal value of the lefthand side is
$-\sum_{j}\left\langle\psi|M_{j}^{\prime}|\psi\right\rangle\log\left\langle\psi|M_{j}^{\prime}|\psi\right\rangle$
since the difference is just the relative entropy of two probability distributions.
Thus we arrive to the entropy inequality (6) as an equivalent of the condition (ii.a). Further, by taking into account that $K(\sigma_{k}^{0})-\Lambda_{0}\geq 0,$ the condition (ii.b) in the form (11) is the same as
$\langle\phi_{k}^{0}|K(\sigma_{k}^{0})|\phi_{k}^{0}\rangle=\langle\phi_{k}^{0}|\Lambda_{0}|\phi_{k}^{0}\rangle$
which via (2) means that the unit vectors $|\phi_{k}^{0}\rangle$, satisfying the constraint $\sum_{k}p_{k}^{0}\left|\phi_{k}^{0}\rangle\langle\phi_{k}^{0}\right|=\overline{\rho},$ turn (6) into the equality (7). ∎
The optimality conditions can be useful in two ways:
  1. 1.
    for a problem which has a hypothetical solution, the conditions can be used to verify its actual optimality. One inserts the conjectured $p_{k}^{0},\sigma_{k}^{0}$ into (9) to compute Hermitean $\Lambda_{0}$. Then the main difficulty will be proof of the entropy inequality (6). In the continuous variable version, this was our strategy in the solution of the related Gaussian optimizers conjecture for the capacity of Gaussian measurement channels [8], with the new generalizations of the famous log-Sobolev inequality as the entropy inequality.
  2. 2.
    for a problem which has a known solution, we get an entropy inequality (6) as a consequence of the optimality. Also one can obtain an alternative proof of the optimality if one can find an independent proof of the entropy inequality. In the next section we illustrate this point with a simple example resulting in the yet nontrivial tight lower bounds for the binary Shannon entropy.
3 Example: two pure states
[…]
$A(\mathcal{E})=1-h\left(\frac{1+\sin\alpha}{2}\right)$ (13)
is attained by the optimal observable =$\mathcal{M=}\left\{M_{+},M_{-}\right\}$ where $M_{\pm}=\left|e_{\pm}\rangle\langle e_{\pm}\right|$ and $|e_{\pm}\rangle=\frac{1}{\sqrt{2}}\left(|e_{0}\rangle\pm|e_{1}\rangle\right)$ is an orthonormal basis [^3] Strictly speaking, the paper [10] contains the check of local optimality.. Here
$h(t)=-t\log t-(1-t)\log(1-t)$
is the classical binary entropy.
Let us apply theorem 1 giving a proof of the global optimality and, as a byproduct, a new entropy inequality. The average state of the ensemble $\mathcal{E}$ is
$\overline{\rho}=\frac{1}{2}\left|\psi_{+}\rangle\langle\psi_{+}\right|+\frac{1}{2}\left|\psi_{-}\rangle\langle\psi_{-}\right|=\cos^{2}\alpha/2\left|e_{0}\rangle\langle e_{0}\right|+\sin^{2}\alpha/2\left|e_{1}\rangle\langle e_{1}\right|$
with
$\overline{\rho}^{1/2}=\cos\alpha/2\left|e_{0}\rangle\langle e_{0}\right|+\sin\alpha/2\left|e_{1}\rangle\langle e_{1}\right|$ (14)
so that
$\frac{1}{\sqrt{2}}\overline{\rho}^{-1/2}|\psi_{\pm}\rangle=|e_{\pm}\rangle$ (15)
and
$M_{\pm}^{\prime}=\frac{1}{2}\overline{\rho}^{-1/2}\rho_{\pm}\overline{\rho}^{-1/2}=\left|e_{\pm}\rangle\langle e_{\pm}\right|,$ (16)
which is also the conjectured optimal measurement. Therefore the conjectured optimal ensemble $\mathcal{E}_{0}^{\prime}$ coincides with $\mathcal{E}$  because
$\sigma_{\pm}^{0}=2\overline{\rho}^{1/2}\left|e_{\pm}\rangle\langle e_{\pm}\right|\overline{\rho}^{1/2}=\left|\psi_{\pm}\rangle\langle\psi_{\pm}\right|,\quad\pi_{\pm}^{0}=\frac{1}{2},$
 so that $|\phi_{\pm}^{0}\rangle=|\psi_{\pm}\rangle.$ Let us check the optimality conditions for the ensemble $\mathcal{E}^{0}=\mathcal{E}$.
Introducing the important parameter
$p=\frac{1+\sin\alpha}{2},\quad\frac{1}{2}\leq p\leq 1,$ (17)
we have
$\left\langle e_{\pm}|\psi_{\pm}\right\rangle=\sqrt{p},\quad\left\langle e_{\pm}|\psi_{\mp}\right\rangle=\sqrt{1-p}.$
Incerting $\sigma_{\pm}^{0}=\left|\psi_{\pm}\rangle\langle\psi_{\pm}\right|$ into
$K(\sigma)=-\sum_{\pm}\left|e_{\pm}\rangle\langle e_{\pm}\right|\log\left\langle e_{\pm}|\sigma|e_{\pm}\right\rangle$
we obtain
$K(\sigma_{+}^{0})$ $=$ $-\left|e_{+}\rangle\langle e_{+}\right|\log p-\left|e_{-}\rangle\langle e_{-}\right|\log(1-p),$
$K(\sigma_{-}^{0})$ $=$ $-\left|e_{+}\rangle\langle e_{+}\right|\log(1-p)-\left|e_{-}\rangle\langle e_{-}\right|\log p.$
Multiplying (12) by $\overline{\rho}^{-1/2}$ from the right and taking into account (15) we obtain
$\Lambda_{0}\overline{\rho}^{1/2}$ $=$ $-\frac{1}{\sqrt{2}}\left(\left|e_{+}\rangle\langle e_{+}\right|+\left|e_{-}\rangle\langle e_{-}\right|\right)\sqrt{p}\log p$
$-\frac{1}{\sqrt{2}}\left(\left|e_{+}\rangle\langle e_{-}\right|+\left|e_{-}\rangle\langle e_{+}\right|\right)\sqrt{1-p}\log(1-p)$
$=$ $-\frac{1}{\sqrt{2}}\left|e_{0}\rangle\langle e_{0}\right|\left(\sqrt{p}\log p+\sqrt{1-p}\log(1-p)\right)$
$-\frac{1}{\sqrt{2}}\left|e_{1}\rangle\langle e_{1}\right|\left(\sqrt{p}\log p-\sqrt{1-p}\log(1-p)\right).$
Taking into account (14) and noticing that
$\cos\alpha/2$ $=$ $\frac{1}{\sqrt{2}}\left(\sqrt{p}+\sqrt{1-p}\right)$
$\sin\alpha/2$ $=$ $\frac{1}{\sqrt{2}}\left(\sqrt{p}-\sqrt{1-p}\right),$
we obtain
$\Lambda_{0}=\lambda_{0}(p)\left|e_{0}\rangle\langle e_{0}\right|-\lambda_{1}(p)\left|e_{1}\rangle\langle e_{1}\right|,$ (18)
where
$\lambda_{0}(p)=-\frac{\sqrt{p}\log p+\sqrt{1-p}\log\left(1-p\right)}{\sqrt{p}+\sqrt{1-p}},$
$\lambda_{1}(p)=\frac{\sqrt{p}\log p-\sqrt{1-p}\log\left(1-p\right)}{\sqrt{p}-\sqrt{1-p}},$
provided $p>1/2$.
Then the conditions (ii.b) are fulfilled while (ii.a) reduces to (6) which via (16), (18) amounts to the following parametric family of tight lower bounds for the classical binary entropy  [^4] See the dynamical plot of the inequality at https://www.desmos.com/Calculator/11ikk1i8fb
Proposition 1.
For $p>1/2$ and $0\leq t\leq 1$ (in our case $t=|\left\langle e_{+}|\psi\right\rangle|^{2},1-t=|\left\langle e_{-}|\psi\right\rangle|^{2}$),
$h(t)\geq(\lambda_{0}(p)+\lambda_{1}(p))\sqrt{t(1-t)}-\frac{\lambda_{1}(p)-\lambda_{0}(p)}{2},$ (19)
where
$\frac{\lambda_{0}(p)+\lambda_{1}(p)}{2}=\sqrt{p\left(1-p\right)}\frac{\log p-\log\left(1-p\right)}{2p-1}\equiv\mu_{0}(p),$
$\frac{\lambda_{1}(p)-\lambda_{0}(p)}{2}=\frac{p\log p-\left(1-p\right)\log\left(1-p\right)}{2p-1}.$
The inequality (19) becomes equality for $t=p$ and $t=1-p.$
In these relations $p>1/2;$ taking the limit $p\rightarrow 1/2$ the inequality (19) becomes
$h(t)\geq 2\log e\sqrt{t(1-t)}-\log\frac{e}{2}.$ (20)
The value $p=1/2$ corresponds to the trivial case $\rho_{+}=\rho_{-}$ with $A(\mathcal{E})=0$. It is remarkable, however, that in the limit $p\rightarrow 1/2$ we get the nontrivial entropy inequality (20). It is equivalent to an inequality which was a starting point in the original derivation of logarithmic Sobolev inequality by L. Gross [3].
Thus (19), (20) can be regarded as discrete “distant relatives” of the log-Sobolev inequality. In the paper [8] we elaborated continuous variable version of the optimality conditions and applied it in the bosonic Gaussian case; the corresponding entropy inequality turned out to be a generalization of the genuine log-Sobolev inequality.
Proof.
For convenience of differentiation we pass to natural logarithms. Denote by $g(t,p)$ the righthand side of (19), then
$\frac{d}{dt}h(t)$ $=$ $-\ln t+\ln(1-t),$ (21)
$\frac{d}{dt}g(t,p)$ $=$ $\mu_{0}(p)\frac{1-2t}{\sqrt{t\left(1-t\right)}}.$ (22)
Taking into account the symmetry of the functions with respect to the point $t=1/2,$ we can restrict to the interval $t\in[1/2,1].$ One can check that $h(p)=g(p,p)$ and $\frac{d}{dt}|_{t=p}h(t)=\frac{d}{dt}|_{t=p}g(t,p)$ which means that $h(t)$ is the envelop of the family $g(t,p);\,p\geq 1/2,$ with $t=p$ as the tangential points. By using (21), (22) we obtain the key relation
$\frac{d}{dt}\left[h(t)-g(t,p)\right]=\left[\mu_{0}(p)-\mu_{0}(t)\right]\frac{2t-1}{\sqrt{t\left(1-t\right)}}.$
The function
$\mu_{0}(t)=\sqrt{t\left(1-t\right)}\frac{\ln t-\ln\left(1-t\right)}{2t-1}$
is monotonously decreasing on $[1/2,1],$ which can be seen by showing its derivative is negative on $(1/2,1]$ (for this it is convenient to make a monotonic change of variable $u=\sqrt{t/\left(1-t\right)}$).
Thus $h(t)-g(t,p)=0$ for $t=p,$ and $\frac{d}{dt}\left[h(t)-g(t,p)\right]$ has the same sign as $t-p.$ Hence $h(t)-g(t,p)\geq 0.$
Theorem 1 then implies the global optimality of the observable =$\mathcal{M=}\left\{M_{+},M_{-}\right\}.$ Also
$\mathrm{Tr\,}\overline{\rho}\Lambda_{0}=\lambda_{0}(p)\cos^{2}\alpha/2-\lambda_{1}(p)\sin^{2}\alpha/2=h(p)$
in accordance with (13).
4 Multidimensional case
4.1 The lower bound for the Shannon entropy: formulation and discussion
The multidimensional generalization of (19) to be derived in what follows is:
Theorem 2.
For integer $m\geq 2$ and $p>\frac{m-1}{m}$
$-\sum_{j=1}^{m}t_{j}\log t_{j}\geq\mu_{0}(p)\left(\sum_{j=1}^{m}\sqrt{t_{j}}\right)^{2}-\mu_{1}(p),$ (23)
where $t_{j}\geq 0,\sum_{j=1}^{m}t_{j}=1$ and
$\mu_{0}(p)=\frac{\sqrt{\frac{p\left(1-p\right)}{m-1}}\left(\log p-\log\frac{1-p}{m-1}\right)}{\left(\sqrt{p}+\left(m-1\right)\sqrt{\frac{1-p}{m-1}}\right)\left(\sqrt{p}-\sqrt{\frac{1-p}{m-1}}\right)},$ (24)
$\mu_{1}(p)=\frac{\sqrt{p}\log p-\sqrt{\frac{1-p}{m-1}}\log\frac{1-p}{m-1}}{\left(\sqrt{p}-\sqrt{\frac{1-p}{m-1}}\right)}.$ (25)
This lower bound is tight: the graphs of the functions of $t_{1},\dots,t_{m}$ in the left and in the right sides of (23) are tangential at the points obtained by permutations of $t_{1}=p,t_{k}=\frac{1-p}{m-1},k=2,\dots,m$.
The inequality (23) arises in connection with the problem of accessible information for ensemble of $m$ equiprobable equiangular pure states discussed in the next subsection. The proof of (23) will be given in section 5. Note that in the case $m=2$ the inequality (23) reduces to (19).
For $m\geq 3$ the value $p=\frac{m-1}{m}$ is allowed giving the inequality
$-\sum_{j=1}^{m}t_{j}\log t_{j}$$\geq$ $\frac{\log\left(m-1\right)}{m-2}\left(\sum_{j=1}^{m}\sqrt{t_{j}}\right)^{2}-\left[\frac{m\log\left(m-1\right)}{m-2}-\log m\right]\quad$ (26)
$=$ $\log m-\frac{m\log\left(m-1\right)}{m-2}\left[1-B(P;P_{U})^{2}\right],$ (27)
where
$B(P;P_{U})=\sum_{j=1}^{m}\sqrt{t_{j}/m}$
is the Bhattacharyya coefficient between the probability distribution $P=\left(t_{1},\dots,t_{m}\right)$ and the uniform distribution $P_{U}=\left(1/m,\dots,1/m\right).$
The inequality (27) then gives a lower bound for the Shannon entropy in terms of a “distance” between the distribution $P$ and the uniform distribution $P_{U}$, related to the Bhattacharyya coefficient. It can be compared with the lower bound of theorem 3.11 in [11]:
$-\sum_{j=1}^{m}t_{j}\log t_{j}\geq\log m-\frac{m\log m}{2\left(m-1\right)}D(P;P_{U}),$ (28)
where $D(P;P_{U})=\sum_{j=1}^{m}\left|t_{j}-1/m\right|$ is the total variation distance. There is well-known relation for any two distributions $P,P^{\prime}$:
$2\left(1-B(P;P^{\prime})\right)\leq D(P;P^{\prime})\leq 2\sqrt{1-B(P;P^{\prime})^{2}},$
which allows to obtain from (27) a degraded version of (28) and vice versa. For $P$ close to $P_{U},$ the bound (27) is substantially better than (28) while close to degenerate distributions with $B(P;P_{U})\approx\sqrt{1/m}$ the right-hand side of (27) becomes negative which never happens for (28).
Similarly to (28), the inequality (27) (as well as (23)) can be rewritten as a tight upper bound for the classical relative entropy (the Kullback-Leibler divergence)
$H(P||P_{U})\leq\frac{m\log\left(m-1\right)}{m-2}\left[1-B(P;P_{U})^{2}\right].$
4.2 The conjecture: “quantum pyramids”
Let $|e_{j}\rangle,\,j=1,\dots,m,$ be an orthonormal basis in $m-$dimensional Hilbert space.
$|e_{0}\rangle=\frac{1}{\sqrt{m}}\sum_{j=1}^{m}|e_{j}\rangle,\quad\langle e_{0}|e_{j}\rangle\equiv\frac{1}{\sqrt{m}}.$
Following Englert and Řeháček  [2] we consider the system of equiangular vectors – the quantum pyramid:
$|\psi_{j}\rangle=\sqrt{mr_{1}}|e_{j}\rangle+\left(\sqrt{r_{0}}-\sqrt{r_{1}}\right)|e_{0}\rangle,\quad j=1,\dots,m,$ (29)
$r_{0},\,r_{1}\geq 0,\quad r_{0}+(m-1)r_{1}=1.$ (30)
Alternative parametrization was given in [4], section VI.2, where this configuration was related to “orthogonal signals” in communication theory:
$|\psi_{j}\rangle=a|e_{j}\rangle+b\sum_{k\neq j}|e_{k}\rangle,\quad j=1,\dots,m,$ (31)
$a\geq 0,\quad a^{2}+(m-1)b^{2}=1.$ (32)
The relation between the two parametrizations is
$a=\sqrt{mr_{1}}+\frac{\sqrt{r_{0}}-\sqrt{r_{1}}}{\sqrt{m}},\quad b=\frac{\sqrt{r_{0}}-\sqrt{r_{1}}}{\sqrt{m}}.$ (33)
The cosine between the edges is $\langle\psi_{j}|\psi_{k}\rangle=r_{0}-r_{1}\,,\,j\neq k,$ thus the case $r_{0}>r_{1}\,(b>0)$ corresponds to the acute, $r_{0}=r_{1}\,(b=0)$ – to the orthogonal, and $r_{0}<r_{1}\,(b<0)$ – to the obtuse pyramids. The extreme cases are $r_{1}=0\,\left(a=b=\sqrt{\frac{1}{m}}\right)$ – the ort $|\psi_{j}\rangle\equiv|e_{0}\rangle,$ and $r_{0}=0\,\left(a=\sqrt{\frac{m-1}{m}},b=-\sqrt{\frac{1}{m\left(m-1\right)}}\right)$ – the flat pyramid.
We will consider the ensemble =$\mathcal{E}=\left\{\pi_{j},\rho_{j}\right\}_{j=1,\dots,m}$ with $\pi_{j}=\frac{1}{m},\,\rho_{j}=|\psi_{j}\rangle\langle\psi_{j}|$ with the average state
$\bar{\rho}\equiv\frac{1}{m}\sum_{k\neq j}^{m}|\psi_{j}\rangle\langle\psi_{j}|=r_{1}I+\left(r_{0}-r_{1}\right)|e_{0}\rangle\langle e_{0}|.$ (34)
Remark 2.
Operators of the form $c_{1}I+c_{0}|e_{0}\rangle\langle e_{0}|$ constitute an Abelian algebra.
One can then easily check that
$\bar{\rho}^{1/2}=\sqrt{r_{1}}I+\left(\sqrt{r_{0}}-\sqrt{r_{1}}\right)|e_{0}\rangle\langle e_{0}|$ (35)
and
$|\psi_{j}\rangle=\sqrt{m}\bar{\rho}^{1/2}|e_{j}\rangle.$ (36)
In what follows we consider the solution for the accessible information $A(\mathcal{E})$ and its optimizing observable conjectured and supported numerically in [2]. In the rest of this section and in the next section we consider acute pyramids for which $r_{0}\geq r_{1}$ or, equivalently, $b\geq 0.$ According to [2] the conjectured optimal observable belongs to the family depending on the parameter $0\leq t\leq 1$ and has the form =$\mathcal{M}=\left\{M_{k};k=0,1,\dots,m\right\}$ where $M_{k}=|\tilde{e}_{k}\rangle\langle\tilde{e}_{k}|$
$|\tilde{e}_{k}\rangle=|e_{k}\rangle+\frac{t-1}{\sqrt{m}}|e_{0}\rangle;\quad k=1,\dots,m,$ (37)
$|\tilde{e}_{0}\rangle=\sqrt{1-t^{2}}|e_{0}\rangle.$
The optimal value of the parameter is
$t_{a}(m)=\left\{\begin{array}[]{cc}1,&r_{0}\leq\frac{4(m-1)}{m^{2}}\\ \frac{2(m-1)}{m-2}\sqrt{\frac{r_{1}}{r_{0}}},&r_{0}>\frac{4(m-1)}{m^{2}}\end{array}\right.,$ (38)
giving the accessible information (eq. (36) in [2])
$A(\mathcal{E})=\left\{\begin{array}[]{cc}I^{(SRM)},&r_{0}\leq\frac{4(m-1)}{m-2}\\ \frac{m\left(1-r_{0}\right)}{m-2}\log\sqrt{\frac{r_{1}}{r_{0}}},&r_{0}>\frac{4(m-1)}{m-2}\end{array}\right.,$ (39)
where
$I^{(SRM)}$ $=$ $\frac{1}{m}\left(\sqrt{r_{0}}+\left(m-1\right)\sqrt{r_{1}}\right)^{2}\log\left(\sqrt{r_{0}}+\left(m-1\right)\sqrt{r_{1}}\right)^{2}$ (40)
-- $\frac{m-1}{m}\left(\sqrt{r_{0}}-\sqrt{r_{1}}\right)^{2}\log\left(\sqrt{r_{0}}-\sqrt{r_{1}}\right)^{2}$
$=$ $\log m+a^{2}\log a^{2}+\left(m-1\right)b^{2}\log b^{2}.$ (41)
Note that in the case $t_{a}(m)=1$ the conjectured information-optimal observable $M_{k}=|{e}_{k}\rangle\langle{e}_{k}|,\,k=1,\dots,m,$ is optimal for the Bayes discrimination problem for the ensemble $\mathcal{E}$ [4], sec. VI.2. It is called square-root measurement (SRM) since the system $\left\{|{e}_{k}\rangle\right\}$ is obtained from $\left\{|{\psi}_{k}\rangle\right\}$ by applying inverse square root of its Gram matrix, cf. (36).
4.3 Checking the optimality conditions: the case of acute pyramids
Let us apply our approach of section 2. By using (36) the dual observable $\mathcal{M}^{\prime}$ has the components
$M_{j}^{\prime}=\overline{\rho}^{-1/2}\left(\frac{1}{m}|\psi_{j}\rangle\langle\psi_{j}|\right)\overline{\rho}^{-1/2}=|e_{j}\rangle\langle e_{j}|,$ (42)
hence
$K\left(\sigma\right)=-\sum_{j}|e_{j}\rangle\langle e_{j}|\,\log\mathrm{\,}\,\langle e_{j}|\sigma|e_{j}\rangle.$
From (37), (35) we obtain the dual ensemble =$\mathcal{E}^{\prime}=\left\{p_{k},\sigma_{k}\right\}$ with
$\sigma_{k}=p_{k}^{-1}|\tilde{\phi}_{k}\rangle\langle\tilde{\phi}_{k}|\,,\quad p_{k}=\langle\tilde{\phi}_{k}|\tilde{\phi}_{k}\rangle$
and
$|\tilde{\phi}_{k}\rangle$ $=$ $\overline{\rho}^{1/2}|\tilde{e}_{k}\rangle=\sqrt{r_{1}}|e_{k}\rangle+\frac{\sqrt{r_{0}}t-\sqrt{r_{1}}}{\sqrt{m}}|e_{0}\rangle;$
$p_{k}$ $=$ $r_{1}+\frac{r_{0}t^{2}-r_{1}}{m};\quad k=1,\dots,m;$
$|\tilde{\phi}_{0}\rangle=\overline{\rho}^{1/2}|\tilde{e}_{0}\rangle=\sqrt{r_{0}\left(1-t^{2}\right)}|e_{0}\rangle;\quad p_{0}=r_{0}\left(1-t^{2}\right).$
Substitution gives
$K\left(\sigma_{k}\right)=-\sum_{j}|e_{j}\rangle\langle e_{j}|\,\log p_{k}^{-1}\mathrm{\,}\,\left|\langle e_{j}|\tilde{\phi}_{k}\rangle\right|^{2},$
hence
$K\left(\sigma_{k}\right)=\left\{\begin{array}[]{cc}-\sum_{j}|e_{j}\rangle\langle e_{j}|\,\log p_{k}^{-1}\mathrm{\,}\,\left|\sqrt{r_{1}}\delta_{jk}+\frac{\sqrt{r_{0}}t-\sqrt{r_{1}}}{m}\right|^{2},&k\geq 1\\ -\sum_{j}|e_{j}\rangle\langle e_{j}|\,\log\frac{1}{m}=\left(\log m\right)I,&k=0\end{array}\right..$
Having in mind checking the optimality condition (ii.b) we compute
$K\left(\sigma_{k}\right)|\tilde{\phi}_{k}\rangle$ $=$ $-\sum_{j}|e_{j}\rangle\left(\sqrt{r_{1}}\delta_{jk}+\frac{\sqrt{r_{0}}t-\sqrt{r_{1}}}{m}\right)\,\log\,\frac{\left|\sqrt{r_{1}}\delta_{jk}+\frac{\sqrt{r_{0}}t-\sqrt{r_{1}}}{m}\right|^{2}}{r_{1}+\frac{r_{0}t^{2}-r_{1}}{m}}$ (43)
$=$ $A(t)|e_{k}\rangle+B(t)\sqrt{m}|e_{0}\rangle;\quad k=1,\dots,m,$
where
$A(t)$ $=$ $-\left(\sqrt{r_{1}}+\frac{\sqrt{r_{0}}t-\sqrt{r_{1}}}{m}\right)\,\log\,\frac{\left|\sqrt{r_{1}}+\frac{\sqrt{r_{0}}t-\sqrt{r_{1}}}{m}\right|^{2}}{r_{1}+\frac{r_{0}t^{2}-r_{1}}{m}}-B(t),$ (44)
$B(t)$ $=$ $-\left(\frac{\sqrt{r_{0}}t-\sqrt{r_{1}}}{m}\right)\,\log\,\frac{\left|\frac{\sqrt{r_{0}}t-\sqrt{r_{1}}}{m}\right|^{2}}{r_{1}+\frac{r_{0}t^{2}-r_{1}}{m}}$ (45)
and
$K\left(\sigma_{0}\right)|\tilde{\phi}_{0}\rangle=\sqrt{r_{0}\left(1-t^{2}\right)}\log m\,|e_{0}\rangle.$
Motivated by the symmetry of the problem (discrete rotations around the axis $|e_{0}\rangle$) and by the remark 2 in previous section we look for $\Lambda_{0}$ in the form
$\Lambda_{0}=c_{1}I+c_{0}|e_{0}\rangle\langle e_{0}|.$ (46)
Then
$\Lambda_{0}|\tilde{\phi}_{k}\rangle$ $=$ $c_{1}\sqrt{r_{1}}|e_{k}\rangle+\frac{\left(c_{0}+c_{1}\right)\sqrt{r_{0}}t-c_{1}\sqrt{r_{1}}}{\sqrt{m}}|e_{0}\rangle;\,k\geq 1,$
$\Lambda_{0}|\tilde{\phi}_{0}\rangle$ $=$ $\left(c_{0}+c_{1}\right)\sqrt{r_{0}\left(1-t^{2}\right)}|e_{0}\rangle.$
Comparing coefficients in the condition (ii.b): $\Lambda_{0}|\tilde{\phi}_{k}\rangle=K\left(\sigma_{k}\right)|\tilde{\phi}_{k}\rangle$, we obtain the system of equations
$c_{1}\sqrt{r_{1}}$ $=$ $A(t),$
$\left(c_{0}+c_{1}\right)\sqrt{r_{0}}t-c_{1}\sqrt{r_{1}}$ $=$ $mB(t),$ (47)
$\left(c_{0}+c_{1}\right)\sqrt{1-t^{2}}$ $=$ $\log m\sqrt{1-t^{2}}.$
Two cases are possible: $t=1$ and $t<1.$ In the first case the third equation disappears while the first two turn into
$c_{1}\sqrt{r_{1}}$ $=$ $\frac{b\log b^{2}-a\log a^{2}}{\sqrt{m}},$
$\left(c_{0}+c_{1}\right)\sqrt{r_{0}}-c_{1}\sqrt{r_{1}}$ $=$ $-\sqrt{m}b\log b^{2},$
where we have used (44), (45) and (33). Taking into account the relations (33), we have $\sqrt{mr_{1}}=a-b,$ and the solution is
$c_{1}$ $=$ $\frac{b\log b^{2}-a\log a^{2}}{a-b},$
$c_{0}$ $=$ $m\frac{ab\left(\log a^{2}-\log b^{2}\right)}{\left(a-b\right)\left[a+\left(m-1\right)b\right]}.$ (48)
In the case $t<1$ the third equation in the system (47) reduces to $c_{0}+c_{1}=\log m.$ Eliminating $c_{0},c_{1}$ one obtains the equation for $t:$
$A(t)+mB(t)=\left(\log m\right)\sqrt{r_{0}}t.$ (49)
Substituting the expressions (44), (45) for $A(t),B(t)$ and introducing the parameter
$\tau=\sqrt{\frac{r_{0}}{r_{1}}}t,$
we can reduce (49) to the form
$\tau\log\left(m-1+\tau^{2}\right)=\frac{m-1+\tau}{m}\log\left(m-1+\tau\right)^{2}+\frac{m-1}{m}\left(\tau-1\right)\log\left(\tau-1\right)^{2}.$ (50)
Denoting $s=\tau^{-1}$ equation (50) can be transformed into
$f(s)$$\equiv$ $\ln\left[1+\left(m-1\right)s^{2}\right]$
-- $0\displaystyle 2\frac{1+\left(m-1\right)s}{m}\ln\left[1+\left(m-1\right)s\right]-2\frac{m-1}{m}\left(1-s\right)\ln\left|1-s\right|=$ 0 .
At the point $s=0$ (corresponding to $\tau\rightarrow\infty$) this function and its first and second derivatives vanish, while $f^{\prime\prime\prime}(s)=2\left(m-1\right)\left(m-2\right)>0$ hence $f$ is increasing and positive for small $s$. But $f(1)=-\ln m<0,$ hence there is a (unique) root in the interval $(0,1],$ namely $s=\frac{m-2}{2\left(m-1\right)}.$ There are no roots for $s>1$ because the function is decreasing.
Thus the value $\tau_{a}(m)=\frac{2(m-1)}{m-2}$ suggested by (38) is the unique positive solution of (49). Substituting this value in the first equation of (47) and using the third equation we obtain
$c_{1}=\log m-\frac{m\log\left(m-1\right)}{m-2},\quad c_{0}=\frac{m\log\left(m-1\right)}{m-2},$ (51)
the values which do not depend on the parameters of the pyramid $r_{0},r_{1}.$ The criterion for the applicability of this second case is $t_{a}(m)\equiv\sqrt{\frac{r_{1}}{r_{0}}}\tau_{a}(m)=\sqrt{\frac{r_{1}}{r_{0}}}\frac{2(m-1)}{m-2}<1$ which, via (30), is equivalent to $r_{0}>\frac{4(m-1)}{m^{2}}$, and, via (33), to $a^{2}<\frac{m-1}{m}.$
Correspondingly, the condition for applicability of the solution (48) is $a^{2}\geq\frac{m-1}{m}.$ By introducing the parameter $p=a^{2}$ we have $b=\sqrt{\frac{1-p}{m-1}},$ leading to
$\Lambda_{0}=c_{1}I+c_{0}|e_{0}\rangle\langle e_{0}|=-\mu_{1}(p)I+\mu_{0}(p)m|e_{0}\rangle\langle e_{0}|$
with $\mu_{1}(p),\mu_{0}(p)$ given by (25), (24) under the condition that $p\geq\frac{m-1}{m}$. Introducing complex numbers $z_{j}=\langle e_{j}|\psi\rangle,$ so that $\sum_{j=1}^{m}\left|z_{j}\right|^{2}=1,$ the optimality condition (ii.a) amounts to the inequality
$-\sum_{j=1}^{m}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}\geq\mu_{0}(p)\left|\sum_{j=1}^{m}z_{j}\right|^{2}-\mu_{1}(p).$ (52)
Since $\mu_{0}(p)\geq 0,$ this is equivalent to (23) where $t_{j}=\left|z_{j}\right|^{2}$
Remarkably, inserting the limiting value $p=\frac{m-1}{m}$ we obtain for $m>2$
$\mu_{1}=\frac{m\log\left(m-1\right)}{m-2}-\log m,\quad\mu_{0}=\frac{\log\left(m-1\right)}{m-2},$
which correspond to the values $c_{0},c_{1},$ given by (51) and obtained by solving (47) in the case $t_{a}(m)<1.$
Finally, from (46) and (34)
$\mathrm{Tr\,}\Lambda_{0}\bar{\rho}=c_{1}+c_{0}r_{0}$
and
$A(\mathcal{E})=\log m-\mathrm{Tr\,}\Lambda_{0}\bar{\rho}=\log m+p\log p+\left(1-p\right)\log\frac{1-p}{m-1}$ (53)
agrees with the expression (39). Thus we have the following corollary to the theorem 2:
Corollary 1.
For ensemble of of acute pyramids the information-optimal observable is SRM given by the equations (37), (38) as conjectured in [2]. The accessible information is given by the expression (53).
4.4 The case of obtuse pyramids
When $r_{0}<r_{1}$ $(b<0)$ the pyramid is obtuse, with extreme flat case $r_{0}=0$ when all the state vectors $|\psi_{j}\rangle$ lie in the hyperplane orthogonal to $|e_{0}\rangle.$ The hypothetical optimal observable is then drastically different. To describe it let us first consider the flat pyramid in the case $m=3$ where $|\psi_{j}\rangle;j=1,2,3,$ are essentially the “trine” of equiangular unit vectors in two-dimensional real plane, cf. [5]. Then the optimal observable is =$\mathcal{M}=\left\{M_{k};k=1,2,3\right\},$ where $M_{k}=\frac{2}{3}|\psi_{k}^{\perp}\rangle\langle\psi_{k}^{\perp}|,$ and $|\psi_{k}^{\perp}\rangle$ are the unit vectors in the plane such that $\langle\psi_{k}^{\perp}|\psi_{k}\rangle=0;k=1,2,3$ [12]. In other words, $\mathcal{M}$ is “unambiguous discrimination” measurement (MUD) for the states $\rho_{j}=|\psi_{j}\rangle\langle\psi_{j}|,\,j=1,2,3.$ The analysis based on our optimality criterion shows that the conditions (ii) of the theorem are fulfilled with $\Lambda_{0}=I,$ with the condition (ii.a) equivalent to the entropy inequality
$-\sum_{j=1}^{3}t_{j}\log t_{j}\geq 1;\quad t_{j}=\frac{2}{3}\cos^{2}\left(\alpha+\frac{2\pi j}{3}\right),$ (54)
saturated for $\alpha=\pi/2+\pi k$.
Most interesting is the case of “nearly flat” trine when $06140<r_{0}<\bar{r}_{0}\approx 0,$ 0614 . Then, as shown by P. Shor in [13], the conjectured optimal observable has 6 outcomes and is a statistical mixture of the MUD described above and the SRM $M_{k}=|{e}_{k}\rangle\langle{e}_{k}|,\,k=1,2,3.$ The SRM becomes optimal for obtuse pyramids with $\bar{r}_{0}\leq r_{0}\leq r_{1}=1/3.$
The generalization of this picture to the case $m\geq 3$ based on extensive numerical studies was suggested and elaborated in [2]. Let us apply our optimality criterion to derive the corresponding entropy inequalities and to obtain additional insight. We consider the ensemble =$\mathcal{E}=\left\{\pi_{j},\rho_{j}\right\}_{j=1,\dots,m}$ with $\pi_{j}=\frac{1}{m},\,\rho_{j}=|\psi_{j}\rangle\langle\psi_{j}|,$ where $|\psi_{j}\rangle$ are given by (29) with $r_{0}<r_{1}$ $(b<0)$ The conjectured optimal observable $\mathcal{M}$ [2] has $m+\frac{m(m-1)}{2}=\frac{m(m+1)}{2}$ components of the form:
$M_{k}$ $=$ $t^{-2}|\tilde{e}_{k}\rangle\langle\tilde{e}_{k}|;\quad k=1,\dots,m;$ (55)
$M_{rs}$ $=$ $\frac{2}{m}\left(1-t^{-2}\right)|rs\rangle\langle rs|;\quad 1\leq r<s\leq m,$
where $t\geq 1,$the vectors $|\tilde{e}_{k}\rangle$ are given by (37), and
$|rs\rangle=\frac{1}{\sqrt{2mr_{1}}}\left(|\psi_{r}\rangle-|\psi_{s}\rangle\right)=\frac{1}{\sqrt{2}}\left(|e_{r}\rangle-|e_{s}\rangle\right).$
Therefore the dual ensemble $\mathcal{E}^{\prime}$ consists of the states (we use the same notation as for similar objects in the acute case):
$\sigma_{k}=p_{k}^{-1}|\tilde{\phi}_{k}\rangle\langle\tilde{\phi}_{k}|\,,\quad p_{k}=\langle\tilde{\phi}_{k}|\tilde{\phi}_{k}\rangle,$
where
$|\tilde{\phi}_{k}\rangle$ $=$ $t^{-1}\overline{\rho}^{1/2}|\tilde{e}_{k}\rangle=t^{-1}\left(\sqrt{r_{1}}|e_{k}\rangle+\frac{\sqrt{r_{0}}t-\sqrt{r_{1}}}{\sqrt{m}}|e_{0}\rangle\right);$
$p_{k}$ $=$ $t^{-2}\left(r_{1}+\frac{r_{0}t^{2}-r_{1}}{m}\right);\quad k=1,\dots,m;$
and
$\sigma_{rs}=p_{rs}^{-1}|\tilde{\phi}_{rs}\rangle\langle\tilde{\phi}_{rs}|\,,\quad p_{rs}=\langle\tilde{\phi}_{rs}|\tilde{\phi}_{rs}\rangle,$
where
$|\tilde{\phi}_{rs}\rangle$ $=$ $\sqrt{\frac{2}{m}\left(1-t^{-2}\right)}\,\overline{\rho}^{1/2}|rs\rangle=\sqrt{\frac{r_{1}}{m}\left(1-t^{-2}\right)}\,\left(|e_{r}\rangle-|e_{s}\rangle\right);$
$p_{rs}$ $=$ $\frac{2r_{1}}{m}\left(1-t^{-2}\right).$
To check the optimality condition (ii.b) we proceed in parallel with the acute case, and compute
$K\left(\sigma_{k}\right)|\tilde{\phi}_{k}\rangle=t^{-1}\left(A(t)|e_{k}\rangle+B(t)\sqrt{m}|e_{0}\rangle\right);\quad k=1,\dots,m,$
where $A(t),B(t)$ are given by (44), (45). Also
$K\left(\sigma_{rs}\right)|\tilde{\phi}_{rs}\rangle=-\sum_{j}|e_{j}\rangle\langle e_{j}|\tilde{\phi}_{rs}\rangle\,\log p_{rs}^{-1}\mathrm{\,}\,\left|\langle e_{j}|\tilde{\phi}_{rs}\rangle\right|^{2},$ (56)
with
$\langle e_{j}|\tilde{\phi}_{rs}\rangle$ $=$ $\sqrt{\frac{r_{1}}{m}\left(1-t^{-2}\right)}\left(\delta_{jr}-\delta_{js}\right),$
$\left|\langle e_{j}|\tilde{\phi}_{rs}\rangle\right|^{2}$ $=$ $\frac{r_{1}}{m}\left(1-t^{-2}\right)\left(\delta_{jr}+\delta_{js}\right).$
Substituting this into (56) results in
$K\left(\sigma_{rs}\right)|\tilde{\phi}_{rs}\rangle=|\tilde{\phi}_{rs}\rangle,\quad 1\leq r<s\leq m.$ (57)
Looking again for $\Lambda_{0}=c_{1}I+c_{0}|e_{0}\rangle\langle e_{0}|,$ and comparing coefficients in the condition (ii.b): $\Lambda_{0}|\tilde{\phi}_{k}\rangle=K\left(\sigma_{k}\right)|\tilde{\phi}_{k}\rangle$, we obtain the first two equations of the system (47), namely
$c_{1}\sqrt{r_{1}}$ $=$ $A(t),$ (58)
$\left(c_{0}+c_{1}\right)\sqrt{r_{0}}t-c_{1}\sqrt{r_{1}}$ $=$ $mB(t),$
while $\Lambda_{0}|\tilde{\phi}_{rs}\rangle=K\left(\sigma_{rs}\right)|\tilde{\phi}_{rs}\rangle$ via (57) amounts to
$\left(1-t^{-2}\right)\left(c_{1}-1\right)=0,$ (59)
since $\langle e_{0}|\tilde{\phi}_{rs}\rangle\equiv 0.$
Again, two cases are possible: $t=1$ and $t>1.$ In the first case the third equation disappears while the first two lead to the same formulas (48) as in the acute case, however with $a=\sqrt{p\text{ }}$, but $b=-\sqrt{\frac{1-p}{m-1}}.$ Then, denoting $z_{j}=\langle e_{j}|\psi\rangle$, the condition (ii.a) brings us to the inequality:
For complex $z_{j}$ , $\sum_{j=1}^{m}\left|z_{j}\right|^{2}=1,\,$
$-\sum_{j=1}^{m}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}\geq\tilde{\mu}_{0}(p)\left|\sum_{j=1}^{m}z_{j}\right|^{2}-\tilde{\mu}_{1}(p),$ (60)
where
$\tilde{\mu}_{0}(p)=\frac{\sqrt{\frac{p\left(1-p\right)}{m-1}}\left(\log p-\log\frac{1-p}{m-1}\right)}{\left(\left(m-1\right)\sqrt{\frac{1-p}{m-1}}-\sqrt{p}\right)\left(\sqrt{p}+\sqrt{\frac{1-p}{m-1}}\right)},$
$\tilde{\mu}_{1}(p)=\frac{\sqrt{p}\log p+\sqrt{\frac{1-p}{m-1}}\log\frac{1-p}{m-1}}{\left(\sqrt{p}+\sqrt{\frac{1-p}{m-1}}\right)}.$
The equality in (60) is attained for $z_{1}=\sqrt{p},\,z_{j}=-\sqrt{\frac{1-p}{m-1}},\,j=2,\dots,m,$ and all the permutations of these $z_{j}$.
Note that here $\tilde{\mu}_{0}(p)\leq 0,$ which prevents us from passing to the probabilities $t_{j}=\left|z_{j}\right|^{2}.$ In (60) the parameter $p$ varies in the range $[\max\left\{\frac{m-1}{m},p(m)\right\},1],$ where $p(m)$ is found below. The accessible information in this case is given by the same expression (40) as in the case of moderately acute pyramids.
If $t>1,$ then (59) imply $c_{1}=1$ and equations (58) reduce to
$\sqrt{r_{1}}$ $=$ $A(t),$
$c_{0}$ $=$ $\frac{m}{\sqrt{r_{0}}t}B(t)+\frac{1}{t}\sqrt{\frac{r_{1}}{r_{0}}}-1.$ (61)
Here the first equation serves for determination of $t$ (compatible with the condition $t>1$). By using (44) it can be transformed to
$\log\frac{m\left(m-1+\tau^{2}\right)}{2}-\frac{m-1+\tau}{m}\log\left(m-1+\tau\right)^{2}-\frac{\left(1-\tau\right)}{m}\log\left(1-\tau\right)^{2}=0,$ (62)
where $\tau=\sqrt{\frac{r_{0}}{r_{1}}}t.$
Denoting $g(\tau)$ the function in equation (62), we have
$g(0)=\log\frac{m}{2}-\frac{m-2}{m}\log\left(m-1\right)\left\{\begin{array}[]{cc}>0,&m\leq 6;\\ <0,&m\geq 7,\end{array}\right..$
The function is decreasing in the interval $[0,1]$ and $g(1)=-\log 2<0$, hence it has a root $\tau_{o}(m)$ in this interval if and only if $m\leq 6.$ For $\tau>1$ the function is negative and has no roots.
The condition $t>1$ amounts to $\tau_{o}(m)>\sqrt{\frac{r_{0}}{r_{1}}},$ which via (30), is equivalent to
$r_{0}<\left(1+\frac{m-1}{\tau_{o}(m)^{2}}\right)^{-1}\equiv r_{0}(m),$
and, via (33), to
$p\equiv a^{2}<\left[1+\left(m-1\right)\left(1-\frac{m}{1-\tau_{o}(m)}\right)^{-2}\right]^{-1}\equiv p(m).$ (63)
The numerically found values:
$m$ 2 3 4 5 6 7
$\tau_{o}(m)$ 0 0,3616 0,25866 0,15195 0,04781 [-0,05286]
$mr_{0}(m)$ 0 0,1841 0,08726 0,02869 0,00274
$p(m)$ 0,5 0,8725 0,8656 0,85699 0,84896 0,8417
$\frac{m-1}{m}$ 0,5 0,6(6) 0,75 0,8 0.83(3) 0,8571
Here $mr_{0}(m)$ show a good agreement with the values found empirically for $m\leq 6$ in [2], eq. (39). In the case $m=7$, as well as for greater values of $m$, the solution of the equation (62) becomes negative and also $p(m)<\frac{m-1}{m}.$ This makes plausible the hypothesis that for $m=7,\dots$ the optimal observable is SRM, and the corresponding entropy inequality is (60) for all $p\geq\frac{m-1}{m}.$
Then, for $m\leq 6$, under the condition (63) the equation (59) implies $c_{1}=1$ while from (61)
$c_{0}=\left(\tau_{o}(m)^{-1}-1\right)\log\frac{2\left(\tau_{o}(m)-1\right)^{2}}{m\left(m-1+\tau_{o}(m)^{2}\right)}.$
The condition (ii.a) then amounts to the entropy inequality (60) with $\tilde{\mu}_{0}=c_{0}/m,\quad\tilde{\mu}_{1}=-1$ i.e.
$-\sum_{j=1}^{m}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}\geq\frac{\left(\tau_{o}(m)^{-1}-1\right)}{m}\log\frac{2\left(\tau_{o}(m)-1\right)^{2}}{m\left(m-1+\tau_{o}(m)^{2}\right)}\left|\sum_{j=1}^{m}z_{j}\right|^{2}+1.$ (64)
The equality is attained if $z_{1}=-z_{2}=1/\sqrt{2},z_{j}=0$ for $j\geq 3$ and for all permutations of these $z_{j}.$
In this case the accessible information is
$A(\mathcal{E})=\log m-\mathrm{Tr\,}\Lambda_{0}\bar{\rho}=\log\frac{m}{2}-r_{0}\left(\tau_{o}(m)^{-1}-1\right)\log\frac{2\left(\tau_{o}(m)-1\right)^{2}}{m\left(m-1+\tau_{o}(m)^{2}\right)}.$
4.5 The case of flat pyramid
The considerations above formally do not include the case of flat pyramid with $r_{0}=0,r_{1}=1/(m-1).$ Then the average state of the ensemble $\mathcal{E}$
$\overline{\rho}=\frac{1}{m-1}\left(I-|0\rangle\langle 0|\right)$
is degenerate and we should use generalized inverse in the formula (42), resulting in
$M_{j}^{\prime}=|\tilde{\psi}_{j}\rangle\langle\tilde{\psi}_{j}|,\quad j=1,\dots,m,$
where $|\tilde{\psi}_{j}\rangle=|e_{j}\rangle-\sqrt{\frac{1}{m}}|e_{0}\rangle.$ Denoting $z_{j}=\langle e_{j}|\psi\rangle$ we have
$\sum_{j=1}^{m}\left|z_{j}\right|^{2}=1,\quad\sum_{j=1}^{m}z_{j}=0.$ (65)
Indeed, the unit vector $|\psi\rangle$ lies in the hyperplane $\langle e_{0}|\psi\rangle=0$, hence the second equality follows.
Arguing as in the previous section, we suggest that the hypothetical optimal observable is given by (55) with $t=1$ for $m\geq 7$ and $t=\infty$ for $m\leq 6$. The corresponding tight entropy inequality is
$-\sum_{j=1}^{m}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}\geq\left\{\begin{array}[]{cc}1,&m\leq 6;\\ \log m-\frac{m-2}{m}\log(m-1),&m\geq 7,\end{array}\right.$ (66)
for complex $z_{j}$ satisfying (65). Indeed, for $m\leq 6$ we can use the argument leading to (64) with $\sum_{j=1}^{m}z_{j}=0.$ On the other hand, for $m\geq 7$ we should use (60) with $\sum_{j=1}^{m}z_{j}=0$ and $p=\frac{m-1}{m}.$ Then the first term in the righthandside vanish, while $\tilde{\mu}_{1}(p)$ gets the value
$-\frac{m-1}{m}\log\frac{m-1}{m}-\frac{1}{m}\log\frac{1}{m(m-1)}=\log m-\frac{m-2}{m}\log(m-1),$ (67)
which is greater than 1 for $m\leq 6$, but less for $m\geq 7$.
Proposition 2.
The tight entropy inequality (66) holds for real (or complex) $z_{j}$ satisfying the conditions (65).
In the case $m\leq 6$ the equality in (66) is attained for $z_{1}=-z_{2}=1/\sqrt{2},\,z_{j}=0$ for $j\geq 3;$ in the case $m\geq 7$ – for $z_{1}=\sqrt{\frac{m-1}{m}},\,z_{j}=-\sqrt{\frac{1}{\left(m-1\right)m}},\,j\geq 2,$ (and for all permutations of such $z_{j}).$
Proof.
By lemma 1 below we can restrict to the case of real $z_{j}$. Introducing the Lagrange function
$L\left(z_{1},\dots,z_{m}\right)=-\sum_{j=1}^{m}z_{j}^{2}\log z_{j}^{2}-\lambda\sum_{j=1}^{m}z_{j}^{2}-\mu\sum_{j=1}^{m}z_{j},$
which is continuously differentiable, we have the necessary condition for a minimizer of the lefthand side of (66)
$-\frac{\partial L}{\partial z_{j}}=2z_{j}\log ez_{j}^{2}+2\lambda z_{j}+\mu=0,\quad j=1,\dots,m.$
Case 1. The minimizer has a zero component $z_{j}$. Then $\mu=0$ and the equation $02z\log ez^{2}+2\lambda z=$ 0 has the roots $0,c,-c.$ Therefore the minimizer should have the form $(\underset{k\text{ times}}{\underbrace{c,-c,\dots,c,-c,}}\,0,\dots,0),$ where $12kc^{2}=$ 1 giving the value $-2kc^{2}\log c^{2}=\log 2k$ which has the minimum equal to 1 for $k=1.$
Case 2. The minimizer has no zero components. The equation $02z\log ez^{2}+2\lambda z+\mu=$ 0 has at most three real, nonzero roots $c_{1},c_{2},c_{3}$, which enter the minimizer with the multiplicities $k_{1},k_{2},$ $k_{3},$ such that $k_{1}+k_{2}+k_{3}=m.$ Then consider the Lagrange function
$\tilde{L}\left(c_{1},c_{2},c_{3};k_{1},k_{2},k_{3}\right)=-\sum_{j=1}^{3}k_{j}c_{j}^{2}\log c_{j}^{2}-\lambda\sum_{j=1}^{3}k_{j}c_{j}^{2}-\mu\sum_{j=1}^{3}k_{j}c_{j}-\nu\sum_{j=1}^{3}k_{j}.$
The necessary condition is
$-\frac{\partial\tilde{L}}{\partial c_{j}}=2k_{j}c_{j}\log ec_{j}^{2}+2\lambda k_{j}c_{j}+\mu k_{j}=0,$
$-\frac{\partial\tilde{L}}{\partial k_{j}}=c_{j}^{2}\log ec_{j}^{2}+\lambda c_{j}^{2}+\mu c_{j}+\nu=0,\quad j=1,2,3.$
Taking into account that $c_{j}\neq 0,k_{j}\geq 1,$ we can multiply the first equation by $c_{j}/\left(2k_{j}\right)$ without lost of information and substract the second equation, resulting in
$c_{j}^{2}-\left(\mu/2\right)c_{j}-\nu=0,$
which has at most two different roots, say $c_{1},c_{2}$, entering the minimizer with the multiplicities $k_{1},k_{2},$ such that $k_{1}+k_{2}=m.$ Then $k_{1}c_{1}+k_{2}c_{2}=0,k_{1}c_{1}^{2}+k_{2}c_{2}^{2}=1$ whence $c_{1}=\sqrt{\frac{k_{2}}{k_{1}m}},c_{1}=-\sqrt{\frac{k_{1}}{k_{2}m}}$ and the value of the lefthandside in (66) is
$-k_{1}c_{1}^{2}\log c_{1}^{2}-k_{2}c_{2}^{2}\log c_{2}^{2}=-\frac{k_{2}}{m}\log\frac{k_{2}}{k_{1}m}-\frac{k_{1}}{m}\log\frac{k_{1}}{k_{2}m}.$
This is a concave function of $k_{1}=m-k_{2}$ symmetric on the interval $[1,m-1]$ hence it attains the minimum at the edges of the interval, which is equal to (67). Since (67) is greater than 1 for $m\leq 6$, but less for $m\geq 7$, we arrive to (66). ∎
The theorem 1 then implies
Corollary 2.
For the ensemble of flat pyramid, information-optimal observable is given by (55) with $t=1$ for $m\geq 7$ and $t=\infty$ for $m\leq 6$.
The corresponding accessible information is [^5] Cf. [2], second line in eq. (33).
$A(\mathcal{E})=\log m-\mathrm{Tr\,}\Lambda_{0}\bar{\rho}=\left\{\begin{array}[]{cc}\log\frac{m}{2},&m\leq 6;\\ \frac{m-2}{m}\log(m-1),&m\geq 7.\end{array}\right.$
For $m=3$ the inequality (66) reduces to (54). Indeed, as shown below, in (66) one can restrict to real $z_{j}$. Then $|\psi\rangle$ is a vector in the real plane $\langle e_{0}|\psi\rangle=0$ and
$z_{j}=\langle e_{j}|\psi\rangle=\langle\tilde{\psi}_{j}|\psi\rangle=\|\tilde{\psi}_{j}\|\,\cos\left(\alpha+\frac{2\pi j}{3}\right),\quad j=1,2,3,$
where $\||\tilde{\psi}_{j}\rangle\|=\||e_{j}\rangle-\sqrt{1/3}|e_{0}\rangle\|=\sqrt{2/3}$, hence (66) is the same as (54).
Lemma 1.
If the inequality (66) holds for all real $z_{j}$ satisfying the conditions (65), then it holds for all complex $z_{j}$ satisfying the same conditions.
Proof.
Let $z_{j}=x_{j}+iy_{j}$ for real $x_{j},y_{j}$, then (65) amounts to
$\sum_{j=1}^{m}x_{j}^{2}+\sum_{j=1}^{m}y_{j}^{2}=1,\quad\sum_{j=1}^{m}x_{j}=0,\,\sum_{j=1}^{m}y_{j}=0.$ (68)
[…]
$\sum_{j=1}^{m}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}$ $=$ $\sum_{j=1}^{m}f\left(P\frac{x_{j}^{2}}{P}+(1-P)\frac{y_{j}^{2}}{1-P}\right)$
$\leq$ $P\sum_{j=1}^{m}f\left(\frac{x_{j}^{2}}{P}\right)+(1-P)\sum_{j=1}^{m}f\left(\frac{y_{j}^{2}}{1-P}\right)$
$=$ $P\sum_{j=1}^{m}\frac{x_{j}^{2}}{P}\log\left(\frac{x_{j}^{2}}{P}\right)$
[…]
because
$\sum_{j=1}^{m}\frac{x_{j}^{2}}{P}=\sum_{j=1}^{m}\frac{y_{j}^{2}}{1-P}=1,\quad\sum_{j=1}^{m}\frac{x_{j}}{\sqrt{P}}=\sum_{j=1}^{m}\frac{y_{j}}{\sqrt{1-P}}=0,$
and by assumption (66) holds for real $z_{j}$. ∎
5 Proof of theorem 2
Let $m\geq 2,\ p\geq\dfrac{m-1}{m}$ and let $\mu_{0}(p),\mu_{1}(p)$ be given by (24), (25).
For $p\in[(m-1)/m,1]$ and a probability distribution $(t_{1},...,t_{m})$, it holds
$F(t):=H(t)-\Bigl{(}\mu_{0}(p)\Bigl{(}\sum\limits_{k=1}^{m}\sqrt{t_{k}}\Bigl{)}^{2}-\mu_{1}(p)\Bigl{)}\geq 0,$ (69)
where $H(t)=-\sum\limits_{k=1}^{m}t_{k}\log t_{k}$ is the Shannon entropy.
Proof.
On the boundary of the simplex of probability distributions the function $F(t)$ cannot reach a minimum, since the function $s\mapsto u(t_{*}+sa)$,
$u(s)=-\sum\limits_{i=1}^{m}(t_{*,i}+sa_{i})\log(t_{*,i}+sa_{i})-\mu_{0}(p)\Bigl{(}\sum\limits_{i=1}^{m}\sqrt{t_{*,i}+sa_{i}}\Bigl{)}+\mu_{1}(p)$
for the boundary point $t_{*}$ and the transversal direction $a$ has the derivative $\infty$- ∞ at zero:
$u^{\prime}(0)=\lim\limits_{t\to t_{*}}\Bigg{[}-\sum\limits_{i\in I_{0}}a_{i}\log t_{i}-C\Bigl{(}\sum\limits_{i\in I_{0}}\dfrac{a_{i}}{\sqrt{t_{i}}}\Bigl{)}+O(1)\Bigg{]},$
where $I_{0}=\{1\leq j\leq m;\ t_{*,j}=0\}$, $C>0$ is some constant. In this case, $a_{i}>0,\ i\in I_{0}$ (direction inside the simplex), so $u^{\prime}(0)=-\infty$.
It is necessary to check the inequality (69) at the critical points (see Lemma 2), which leads to the inequality
$\mu_{0}(p,1/m)\Bigl{(}\sqrt{q(k/m)}+\sqrt{(1-q)(1-k/m)}\Bigl{)}^{2}-\mu_{1}(p,1-1/m)\quad$
$\leq$ $-q\log\dfrac{q}{k/m}-(1-q)\log\dfrac{1-q}{1-k/m},$
where $\mu_{0}(p,\alpha)$ and $\mu_{1}(p,\alpha)$ are defined in (70) and (71) below. Lemma 4 states the validity of this inequality. ∎
Consider the functions
$\mu_{0}(p,\alpha)=\dfrac{\sqrt{\dfrac{p(1-p)}{\alpha(1-\alpha)}}\Bigl{(}\log\dfrac{p}{\alpha}-\log\dfrac{1-p}{1-\alpha}\Big{)}}{\Bigl{(}\sqrt{\alpha p}+\sqrt{(1-\alpha)(1-p)}\Bigl{)}\Bigl{(}\sqrt{\dfrac{p}{\alpha}}-\sqrt{\dfrac{1-p}{1-\alpha}}\Bigl{)}},$ (70)
$\mu_{1}(p,\alpha)=\dfrac{\sqrt{\dfrac{p}{\alpha}}\log\dfrac{p}{\alpha}-\sqrt{\dfrac{1-p}{1-\alpha}}\log\dfrac{1-p}{1-\alpha}}{\sqrt{\dfrac{p}{\alpha}}-\sqrt{\dfrac{1-p}{1-\alpha}}}.$ (71)
The following lemmas lead to the proof of the main theorem.
Lemma 2.
The critical points $t_{*}=(t_{*,1},...,t_{*,m})$ of the function
$F(t)=H(t)-\Bigl{(}\mu_{0}(p)\Bigl{(}\sum\limits_{k=1}^{m}\sqrt{t_{k}}\Bigl{)}^{2}-\mu_{1}(p)\Bigl{)}$
have the form
$t_{*}=\Bigl{(}q/k,...,q/k,(1-q)/(m-k),...,(1-q)/(m-k)\Bigl{)}$
up to permutations, for some $1\leq k\leq m$ and $q\in[0,1]$ (for $k=m$ it is assumed that $q=1/m$). One of the critical points is equal to $(p,(1-p)/(m-1),...,(1-p)/(m-1))$, in which the value of the function is equal to 00.
Proof.
Instead of $F(t)$ we consider a function of the form $u(s)=F(t+sa)$, where $t=(t_{1},...,t_{m}),\ a=(a_{1},...,a_{m}),\ \sum\limits_{i}a_{i}=0$, on such a range of values of $s$ that $t+as$ defines a probability distribution.
$\dfrac{d}{ds}u(s)$ $=-\sum\limits_{i=1}^{m}a_{i}\log t_{i}-\mu_{0}(p)\Bigl{(}\sum\limits_{i=1}^{m}\dfrac{a_{i}}{\sqrt{t_{i}}}\Bigl{)}\Bigl{(}\sum\limits_{i=1}^{m}\sqrt{t_{i}}\Bigl{)}$
$=\sum\limits_{i=2}^{m}a_{i}\Bigg{[}-\log t_{i}+\log t_{1}-\mu_{0}(p)\Bigl{(}\sum\limits_{k=1}^{m}\sqrt{t_{k}}\Bigl{)}\Bigl{(}\dfrac{1}{\sqrt{t_{i}}}-\dfrac{1}{\sqrt{t_{1}}}\Bigl{)}\Bigg{]}.$
The condition $\dfrac{d}{ds}u(0)=0$ for a point $t_{*}=t$ and an arbitrary vector $a$ (that is, the point $t_{*}$ is critical) is satisfied if and only if for any $2\leq i\leq m$ it holds
$\log t_{*,1}-\log t_{*,i}=\mu_{0}(p)\Bigl{(}\sum\limits_{k=1}^{m}\sqrt{t_{*,k}}\Bigl{)}\Bigl{(}\dfrac{1}{\sqrt{t_{*,i}}}-\dfrac{1}{\sqrt{t_{*,1}}}\Bigl{)}.$
Let us denote $a=\ln t_{*,1}+\dfrac{\mu_{0}(p)\ln(2)}{\sqrt{t_{*,1}}}\Bigl{(}\sum\limits_{k=1}^{m}\sqrt{t_{*,k}}\Bigl{)},\ C=\mu_{0}(p)\ln(2)\Bigl{(}\sum\limits_{k=1}^{m}\sqrt{t_{*,k}}\Bigl{)}$. The equation
$a-\ln(s)=C/\sqrt{s},\quad s>0,$ (72)
has no more than two roots, since the derivatives of the right and left parts with respect to $s$ coincide only when $s=C^{2}/4$. But all the numbers $t_{*,1},...,t_{*,m}$ are solutions of the equation (72). Thus, among the values of $\{t_{*,i}\}$ there are no more than two different values, and the lemma is proved. ∎
Lemma 3.
Let $A\geq 0,\ B,\ C\in\mathbb{R}$. Then the function
$f(t)=h(t)-(A\sqrt{t(1-t)}+Bt+C)$ (73)
reaches a local minimum on $[0,1]$ at no more than two points $t_{1}\leq 1/2\leq t_{2}$ which are critical. In this case, if $B>0$ and $t_{*}>1/2$ is critical ($B<0$ and $t_{*}<1/2$), then $t_{*}$ is a point of the global minimum.
Proof.
$f^{\prime}(0)=-\infty,\quad f^{\prime}(1)=+\infty,$
therefore local minima are on the interval $(0,1)$.
In the calculations we will move on to the natural logarithm.
$f^{\prime}(t)=-\ln\dfrac{t}{1-t}-\dfrac{A\sqrt{t(1-t)}}{2}\Bigl{(}\dfrac{1}{t}-\dfrac{1}{1-t}\Bigl{)}-B=$
$=-\ln\dfrac{t}{1-t}-\dfrac{A(1-2t)}{2\sqrt{t(1-t)}}-B;$
$f^{\prime\prime}(t)=-\dfrac{1}{t(1-t)}+\dfrac{A}{4(t(1-t))^{3/2}}.$
$f^{\prime\prime}(t)=0\quad\Leftrightarrow\quad A=4\sqrt{t(1-t)}.$
Therefore, there are two inflection points located on different sides of $t=1/2$. It follows that there are 3 critical points $t_{1}\leq t_{0}\leq t_{2}$, where $t_{1},\ t_{2}$ are local minimum points.
[…]
The variables $\alpha,\ p,\beta,\ t$ used in what follows, unless otherwise stated, are such that
$\alpha\in[0,1/2],\quad p\in[1-\alpha,1],\quad\beta\in[\alpha,1/2],\quad t\in[0,1].$
Lemma 4.
For introduced $\alpha,\ p,\ \beta,\ t$ the inequality
$\mu_{0}(p,\alpha)\Bigl{(}\sqrt{\beta t}+\sqrt{(1-\beta)(1-t)}\Bigl{)}^{2}-\mu_{1}(p,\alpha)$ (74)
$\leq$ $-t\log\dfrac{t}{\beta}-(1-t)\log\dfrac{1-t}{1-\beta}$
holds.
Proof.
Let
$f(t)=-t\log\dfrac{t}{\beta}-(1-t)\log\dfrac{1-t}{1-\beta}-\Bigl{(}\mu_{0}(p,\alpha)\Bigl{(}\sqrt{\beta t}+\sqrt{(1-\beta)(1-t)}\Bigl{)}^{2}-\mu_{1}(p,\alpha)\Bigl{)},$
for which $A=2\mu_{0}(p,\alpha)\sqrt{\beta(1-\beta)}$, $B=\log\dfrac{1-\beta}{\beta}+\mu_{0}(p,\alpha)(2\beta-1)$ in the form (73).
The derivative of $f(t)$ has form
$f^{\prime}(t)=\log\dfrac{1-t}{t}+\log\dfrac{\beta}{1-\beta}-\mu_{0}(p,\alpha)\Bigg{(}\sqrt{\dfrac{\beta}{t}}-\sqrt{\dfrac{1-\beta}{1-t}}\Bigg{)}\Bigl{(}\sqrt{\beta t}+\sqrt{(1-\beta)(1-t)}\Bigl{)}.$
Consider the following cases:
  1. 1.
    $B\leq 0$.
    In this case there is minimum point $t_{*}=\beta$ ($f^{\prime}(\beta)=0$) of the function $f(t)$ on $[0,1/2]$ due to the Lemma 3. One has
    $f(\beta)=\mu_{1}(p,\alpha)-\mu_{0}(p,\alpha)\geq 0.$ (75)
    Indeed, the inequality (75) is equivalent to
    $\Bigl{(}\sqrt{\dfrac{p}{\alpha}}\log\dfrac{p}{\alpha}-\sqrt{\dfrac{1-p}{1-\alpha}}\log\dfrac{1-p}{1-\alpha}\Bigl{)}\Bigl{(}\sqrt{p\alpha}+\sqrt{(1-p)(1-\alpha)}\Bigl{)}-$
    $-\sqrt{\dfrac{p(1-p)}{\alpha(1-\alpha)}}\Bigl{(}\log\dfrac{p}{\alpha}-\log\dfrac{1-p}{1-\alpha}\Bigl{)}\geq 0\quad\Leftrightarrow$
    $p\log\dfrac{p}{\alpha}\Bigg{(}1-\sqrt{\dfrac{\alpha(1-p)}{p(1-\alpha)}}\Bigg{)}-(1-p)\log\dfrac{1-p}{1-\alpha}\Bigg{(}1-\sqrt{\dfrac{p(1-\alpha)}{\alpha(1-p)}}\Bigg{)}\geq 0\quad\Leftrightarrow$
    $\sqrt{\alpha p}\log\dfrac{p}{\alpha}+\sqrt{(1-\alpha)(1-p)}\log\dfrac{1-p}{1-\alpha}\geq 0.\quad$ (76)
    Let $L(p)=\sqrt{\alpha p}\ln\dfrac{p}{\alpha}+\sqrt{(1-\alpha)(1-p)}\ln\dfrac{1-p}{1-\alpha}$, then
    1. (a)
      $(d/dp)^{2}L(p)=-\dfrac{\sqrt{\alpha}\ln(p/\alpha)}{4p^{3/2}}-\dfrac{\sqrt{1-\alpha}\ln((1-p)/(1-\alpha))}{4(1-p)^{3/2}}\geq 0$
      ($p\geq 1-\alpha\geq 1/2$);
    2. (b)
      $L(1-\alpha)=0$,
      $L^{\prime}(1-\alpha)=\sqrt{\dfrac{\alpha}{1-\alpha}}-1+\dfrac{1}{2}\ln\dfrac{\alpha}{1-\alpha}+\dfrac{1}{2}\sqrt{\dfrac{\alpha}{1-\alpha}}\ln\dfrac{1-\alpha}{\alpha}\geq 0\quad\Leftrightarrow$
      $\dfrac{\alpha}{1-\alpha}\Bigl{(}1+\dfrac{1}{2}\ln\dfrac{1-\alpha}{\alpha}\Bigl{)}-\Bigl{(}1-\dfrac{1}{2}\ln\dfrac{1-\alpha}{\alpha}\Bigl{)}\geq 0\quad\Leftrightarrow$
      $\ln\dfrac{1-\alpha}{\alpha}\geq 2-4\alpha,\quad\alpha\in(0,1/2].$
    Therefore, $L(p)\geq 0$ and the inequality (76) is satisfied.
  2. 2.
    $B\geq 0$.
    This condition implies that the function $f(t)$ has the minimum point $t_{*}$ on $[1/2,1]$. This point is characterized by the equation
    $\log\dfrac{1-t_{*}}{t_{*}}+\log\dfrac{\beta}{1-\beta}-\mu_{0}(p,\alpha)\Bigg{(}\sqrt{\dfrac{\beta}{t_{*}}}-\sqrt{\dfrac{1-\beta}{1-t_{*}}}\Bigg{)}\Bigl{(}\sqrt{\beta t_{*}}+\sqrt{(1-\beta)(1-t_{*})}\Bigl{)}=0.$
    The value of $t_{*}$ depends on $\beta$, and
    $\dfrac{d}{d\beta}f(t_{*})=\dfrac{1}{\ln 2}\Bigl{(}\dfrac{t_{*}}{\beta}-\dfrac{1-t_{*}}{1-\beta}\Bigl{)}-\mu_{0}(p,\alpha)\Bigg{(}\sqrt{\dfrac{t_{*}}{\beta}}-\sqrt{\dfrac{1-t_{*}}{1-\beta}}\Bigg{)}\times$
    ×$\times\Bigl{(}\sqrt{\beta t_{*}}+\sqrt{(1-\beta)(1-t_{*})}\Bigl{)}.$
    Then
    $\ln 2\ \dfrac{d}{d\beta}f(t_{*})=\Bigl{(}\dfrac{t_{*}}{\beta}-\dfrac{1-t_{*}}{1-\beta}\Bigl{)}+\sqrt{\dfrac{t_{*}(1-t_{*})}{\beta(1-\beta)}}\Bigl{(}\ln\dfrac{1-t_{*}}{t_{*}}-\ln\dfrac{1-\beta}{\beta}\Bigl{)}\geq 0$
    since
    $\sqrt{\dfrac{1-\beta}{\beta}\dfrac{t_{*}}{1-t_{*}}}-\sqrt{\dfrac{\beta}{1-\beta}\dfrac{1-t_{*}}{t_{*}}}+\ln\Bigl{(}\dfrac{\beta}{1-\beta}\dfrac{1-t_{*}}{t_{*}}\Bigl{)}\equiv C-1/C-2\ln C\geq 0$
    for $C\geq 1$.
    It is simple to calculate that for $\beta=\alpha$, we have $t_{*}=p$, and $f(p)=0$. Therefore, $f(t_{*})\geq 0$ for all $\alpha\leq\beta\leq 1/2$, and $f(t)\geq 0$.
In the particular case of (76), where $\alpha=1/2$, one has an interesting inequality
$h_{1/2}(p)\equiv\sqrt{2p}\log(2p)+\sqrt{2(1-p)}\log(2(1-p))\geq 0,$
which takes place for $p\in(0,1)$. The function $h_{1/2}(p)$ is nonnegative, convex and $h_{1/2}(p)\leq\sqrt{2}$.
Acknowledgment. This work is supported by Russian Science Foundation under the grant No 24-11-00145. The authors are grateful to S.V. Grishin and D.A. Kronberg for providing numerical simulations.

### v2

@@
 The recently obtained optimality criterion [7] when applied to specific ensembles of states leads to nontrivial tight entropy inequalities that are discrete relatives of the famous log-Sobolev inequality.
-In this light, the hypothesis of globally information-optimal measurement for an ensemble of equiangular equiprobable states [2] is reconsidered and the corresponding entropy inequalities are proposed.
-We prove these inequalities in the cases of state ensembles corresponding to acute or flat pyramids, thus providing the first proof of the hypotheses concerning globally information-optimal observables put forward in the paper [2].
+In this light, the hypothesis of globally information-optimal measurement for an ensemble of equiangular equiprobable states [2] is reconsidered and the corresponding entropy inequalities are proposed and proved.
+Via the optimality criterion, this provides also the first proof of the conjecture concerning globally information-optimal observables for quantum pyramids put forward in [2].
 1
@@
 The problem was solved in some special (notably, symmetric) cases but in general remains open.
-In this paper we describe a general approach following [7] which gives a criterion for optimality and ties up the problem with entropy inequalities – the tight lower bounds for the classical Shannon information.
-It turns out that the optimality criterion [7] when applied to specific ensembles of states leads to new nontrivial entropy inequalities that are discrete “distant relatives” of the famous log-Sobolev inequality.
-In this light, the hypothesis of an information-optimal measurement for an ensemble of equiangular equiprobable states – quantum pyramids formulated and numerically substantiated by Englert and Řeháček in [2] – is reconsidered in the present paper.
-In the case of acute pyramids the corresponding parametric family of entropy inequalities (23) is proposed in the section 4.1.
+In section (2) of this paper we describe a general approach following [7] which gives a criterion for optimality and ties up the problem with entropy inequalities – the tight lower bounds for the classical Shannon information.
+It turns out that the optimality criterion when applied to specific ensembles of states leads to new nontrivial entropy inequalities that are discrete “distant relatives” of the famous log-Sobolev inequality, as explained in section 3.
+In this light, the conjecture of an information-optimal measurement for an ensemble of equiprobable equiangular states – quantum pyramid, formulated and numerically substantiated by Englert and Řeháček in [2] – is reconsidered in section 4 of the present paper.
+In the case of acute pyramids the corresponding parametric family of entropy inequalities (27) is proposed in subsections 4.1-4.3.
 The inequalities suggest the new tight lower bounds for the Shannon entropy of a discrete probability distribution as compared e.g.
 to the known bound from [11], theorem 3.11.
-The proof of the inequality (23), given in section 5, was obtained by the second author as a response to the conjecture put forward by the first author.
-Via the optimality criterion of section 2, it also implies the validity of a hypothesis concerning accessible information for state ensembles of acute quantum pyramids, formulated and numerically substantiated in [2] (corollary 1 to theorem 2).
-In the section 4.4 the case of obtuse pyramids is considered to get further insight and to formulate the hypotheses concerning the corresponding entropy inequalities (60), (64).
+In the subsection 4.4 the case of obtuse pyramids is considered to get further insight and to formulate the hypotheses concerning the corresponding entropy inequalities (66), (72).
 Our approach allows also to derive the nonlinear equations (54) (resp.
 (70)) serving to determine the critical values of the parameter $t$ of the optimal observable corresponding to transition from moderately to strongly acute (resp, obtuse) pyramids.
-In the section 4.5 we derive and prove the tight entropy inequality (66) implying the global optimality of the hypothetical observable from [2] (proposition 2 and corollary 2).
-The findings in sections 4.4, 4.5 also emphasize that for obtuse and flat pyramids in the dimensions $m\geq 7$ the sharp “square root” observable (SRM) outperforms the “unambiguous discrimination” observable (UDM) and is the optimal one in the flat case (tentatively also in the obtuse case).
+In the subsection 4.5 we derive and prove the tight entropy inequality (75) implying the global optimality of the hypothetical observable from [2] (proposition 2 and corollary 3) in the case of flat pyramids.
+The findings in subsections 4.4, 4.5 also emphasize that for obtuse and flat pyramids in the dimensions $m\geq 7$ the sharp “square root” observable (SRM) outperforms the “unambiguous discrimination” observable (UDM) and is the optimal one.
+The proofs of the inequality (27), given in subsection 5.1, as well as of the inequalities (66), (72) in subsection 5.2 was obtained by the second author as a response to the hypotheses put forward by the first author.
+Via the optimality criterion of section 2, it also implies the validity of the conjectures concerning accessible information for quantum pyramids, formulated and numerically substantiated in [2] (corollary 1 to theorem 2, corollary 2 to theorem 3).
 2
@@
 $A(\mathcal{E})=\sup_{\mathcal{M}}I(\mathcal{E},\mathcal{M}).$
-It was shown [1] that the supremum is attained on an observable of the form $M_{k}=\left|\varphi_{k}\rangle\langle\varphi_{k}\right|,\,k=1,\dots,n$ with $d\leq n\leq d^{2}$ and linearly independent $M_{k}$.
+It was shown [1] that the supremum is attained on an observable of the form $M_{k}=\left|\varphi_{k}\rangle\langle\varphi_{k}\right|,\,k=1,\dots,n$ with $d\leq n\leq d^{2}$ and linearly independent $M_{k}$ (if the matrices of $\rho_{j}$ in some basis are real then the number of components is reduced to $\frac{d(d+1)}{2}$).
 In particular, the supremum can be replaced with the maximum.
@@
 and observable $\mathcal{M}^{\prime}$ with the components [^1] In the case $\overline{\rho}$ is degenerate, one should use generalized inverse in the following formula.
-$M_{j}^{\prime}=\overline{\rho}^{-1/2}\left(\pi_{j}\rho_{j}\right)\overline{\rho}^{-1/2}$
+$M_{j}^{\prime}=\overline{\rho}^{-1/2}\left(\pi_{j}\rho_{j}\right)\overline{\rho}^{-1/2}$ (1)
 we have the same average state
@@
 The value of the accessible information (4) is
-$A(\mathcal{E})=H(\pi)-\mathrm{Tr\,}\overline{\rho}\Lambda_{0}.$
+$A(\mathcal{E})=H(\pi)-\mathrm{Tr\,}\overline{\rho}\Lambda_{0}.$ (9)
 Proof.
@@
 ∎
+Especially important for us is the case of pure-states ensemble with $\rho_{j}=|\psi_{j}\rangle\langle\psi_{j}|$ where $|\psi_{j}\rangle;j=1,\dots,d$, are linearly independent vectors.
+In this case the resolution of the identity (1) is
+$M_{j}^{\prime}=|e_{j}\rangle\langle e_{j}|,$
+where the vectors
+$|e_{j}\rangle=\sqrt{\pi_{j}}\bar{\rho}^{-1/2}|\psi_{j}\rangle;\quad j=1,\dots,d,$ (15)
+by necessity form an orthonormal basis. Then the entropy inequality (7) takes the form
+$-\sum_{j=1}^{d}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}\geq\sum_{j=1}^{d}\lambda_{jk}\overline{z_{j}}z_{k},$ (16)
+where $\lambda_{jk}=\langle e_{j}|\Lambda_{0}|e_{k}\rangle$ and $z_{j}=\langle e_{j}|\psi\rangle$ are complex variables subject only to the constraint $\sum_{j=1}^{d}\left|z_{j}\right|^{2}=1$.
+The left-hand side is just the Shannon entropy of the p.d.
+$\left\{\left|z_{j}\right|^{2};j=1,\dots,d\right\}$, and such an inequality can be approached with the tools of classical analysis.
 The optimality conditions can be useful in two ways:
@@
 The inequality (27) arises in connection with the problem of accessible information for ensemble of $m$ equiprobable equiangular pure states discussed in the next subsection.
-The proof of (23) will be given in section 5.
+The proof of (27) will be given in section 5.1.
 Note that in the case $m=2$ the inequality (27) reduces to (23).
@@
 We will consider the ensemble =$\mathcal{E}=\left\{\pi_{j},\rho_{j}\right\}_{j=1,\dots,m}$ with $\pi_{j}=\frac{1}{m},\,\rho_{j}=|\psi_{j}\rangle\langle\psi_{j}|$ with the average state
-$\bar{\rho}\equiv\frac{1}{m}\sum_{k\neq j}^{m}|\psi_{j}\rangle\langle\psi_{j}|=r_{1}I+\left(r_{0}-r_{1}\right)|e_{0}\rangle\langle e_{0}|.$ (34)
+$\bar{\rho}\equiv\frac{1}{m}\sum_{j=1}^{m}|\psi_{j}\rangle\langle\psi_{j}|=r_{1}I+\left(r_{0}-r_{1}\right)|e_{0}\rangle\langle e_{0}|.$ (38)
 Remark 2.
@@
 In what follows we consider the solution for the accessible information $A(\mathcal{E})$ and its optimizing observable conjectured and supported numerically in [2].
-In the rest of this section and in the next section we consider acute pyramids for which $r_{0}\geq r_{1}$ or, equivalently, $b\geq 0.$ According to [2] the conjectured optimal observable belongs to the family depending on the parameter $0\leq t\leq 1$ and has the form =$\mathcal{M}=\left\{M_{k};k=0,1,\dots,m\right\}$ where $M_{k}=|\tilde{e}_{k}\rangle\langle\tilde{e}_{k}|$
+In the rest of this subsection and in the next subsection we consider acute pyramids for which $r_{0}\geq r_{1}$ or, equivalently, $b\geq 0.$ According to [2] the conjectured optimal observable belongs to the family depending on the parameter $0\leq t\leq 1$ and has the form =$\mathcal{M}=\left\{M_{k};k=0,1,\dots,m\right\}$ where $M_{k}=|\tilde{e}_{k}\rangle\langle\tilde{e}_{k}|$
 $|\tilde{e}_{k}\rangle=|e_{k}\rangle+\frac{t-1}{\sqrt{m}}|e_{0}\rangle;\quad k=1,\dots,m,$ (41)
@@
 giving the accessible information (eq. (36) in [2])
-$A(\mathcal{E})=\left\{\begin{array}[]{cc}I^{(SRM)},&r_{0}\leq\frac{4(m-1)}{m-2}\\ \frac{m\left(1-r_{0}\right)}{m-2}\log\sqrt{\frac{r_{1}}{r_{0}}},&r_{0}>\frac{4(m-1)}{m-2}\end{array}\right.,$ (39)
+$A(\mathcal{E})=\left\{\begin{array}[]{cc}I^{(SRM)},&r_{0}\leq\frac{4(m-1)}{m^{2}}\\ \frac{m\left(1-r_{0}\right)}{m-2}\log(m-1),&r_{0}>\frac{4(m-1)}{m^{2}}\end{array}\right.,$ (43)
 where
 $I^{(SRM)}$ $=$ $\frac{1}{m}\left(\sqrt{r_{0}}+\left(m-1\right)\sqrt{r_{1}}\right)^{2}\log\left(\sqrt{r_{0}}+\left(m-1\right)\sqrt{r_{1}}\right)^{2}$ (44)
--- $\frac{m-1}{m}\left(\sqrt{r_{0}}-\sqrt{r_{1}}\right)^{2}\log\left(\sqrt{r_{0}}-\sqrt{r_{1}}\right)^{2}$
+$+$ $\frac{m-1}{m}\left(\sqrt{r_{0}}-\sqrt{r_{1}}\right)^{2}\log\left(\sqrt{r_{0}}-\sqrt{r_{1}}\right)^{2}$
 $=$ $\log m+a^{2}\log a^{2}+\left(m-1\right)b^{2}\log b^{2}.$ (45)
@@
 $\sigma_{k}=p_{k}^{-1}|\tilde{\phi}_{k}\rangle\langle\tilde{\phi}_{k}|\,,\quad p_{k}=\langle\tilde{\phi}_{k}|\tilde{\phi}_{k}\rangle$
-and
+where
 $|\tilde{\phi}_{k}\rangle$ $=$ $\overline{\rho}^{1/2}|\tilde{e}_{k}\rangle=\sqrt{r_{1}}|e_{k}\rangle+\frac{\sqrt{r_{0}}t-\sqrt{r_{1}}}{\sqrt{m}}|e_{0}\rangle;$
@@
 $K\left(\sigma_{0}\right)|\tilde{\phi}_{0}\rangle=\sqrt{r_{0}\left(1-t^{2}\right)}\log m\,|e_{0}\rangle.$
-Motivated by the symmetry of the problem (discrete rotations around the axis $|e_{0}\rangle$) and by the remark 2 in previous section we look for $\Lambda_{0}$ in the form
+Motivated by the symmetry of the problem (discrete rotations around the axis $|e_{0}\rangle$) and by the remark 2 in previous subsection we look for $\Lambda_{0}$ in the form
 $\Lambda_{0}=c_{1}I+c_{0}|e_{0}\rangle\langle e_{0}|.$ (50)
@@
 $c_{1}=\log m-\frac{m\log\left(m-1\right)}{m-2},\quad c_{0}=\frac{m\log\left(m-1\right)}{m-2},$ (55)
-the values which do not depend on the parameters of the pyramid $r_{0},r_{1}.$ The criterion for the applicability of this second case is $t_{a}(m)\equiv\sqrt{\frac{r_{1}}{r_{0}}}\tau_{a}(m)=\sqrt{\frac{r_{1}}{r_{0}}}\frac{2(m-1)}{m-2}<1$ which, via (30), is equivalent to $r_{0}>\frac{4(m-1)}{m^{2}}$, and, via (33), to $a^{2}<\frac{m-1}{m}.$
-Correspondingly, the condition for applicability of the solution (48) is $a^{2}\geq\frac{m-1}{m}.$ By introducing the parameter $p=a^{2}$ we have $b=\sqrt{\frac{1-p}{m-1}},$ leading to
+the values which do not depend on the parameters of the pyramid $r_{0},r_{1}.$ The criterion for the applicability of this second case is
+$t_{a}(m)\equiv\sqrt{\frac{r_{1}}{r_{0}}}\tau_{a}(m)=\sqrt{\frac{r_{1}}{r_{0}}}\frac{2(m-1)}{m-2}<1$
+which, via (34), is equivalent to $r_{0}>\frac{4(m-1)}{m^{2}}$, and, via (37), to $a^{2}<\frac{m-1}{m}.$
+Correspondingly, the condition for applicability of the solution (52) is $a^{2}\geq\frac{m-1}{m}.$ By introducing the parameter $p$ such that
+$p=a^{2},\quad\frac{1-p}{m-1}=b^{2},$ (56)
+we have $b=\sqrt{\frac{1-p}{m-1}},$ leading to
 $\Lambda_{0}=c_{1}I+c_{0}|e_{0}\rangle\langle e_{0}|=-\mu_{1}(p)I+\mu_{0}(p)m|e_{0}\rangle\langle e_{0}|$
@@
 which correspond to the values $c_{0},c_{1},$ given by (55) and obtained by solving (51) in the case $t_{a}(m)<1.$
-Finally, from (46) and (34)
-$\mathrm{Tr\,}\Lambda_{0}\bar{\rho}=c_{1}+c_{0}r_{0}$
-and
-$A(\mathcal{E})=\log m-\mathrm{Tr\,}\Lambda_{0}\bar{\rho}=\log m+p\log p+\left(1-p\right)\log\frac{1-p}{m-1}$ (53)
-agrees with the expression (39). Thus we have the following corollary to the theorem 2:
+Finally, from (50) and (38) $\mathrm{Tr\,}\Lambda_{0}\bar{\rho}=c_{1}+c_{0}r_{0}$ so that
+$A(\mathcal{E})=\log m-c_{1}-c_{0}r_{0}.$ (58)
+Thus we have the following corollary to the theorem 2:
 Corollary 1.
-For ensemble of of acute pyramids the information-optimal observable is SRM given by the equations (37), (38) as conjectured in [2]. The accessible information is given by the expression (53).
+For ensemble of acute pyramids the information-optimal observable is given by the equations (41), (42) as conjectured in [2]. The accessible information is given by the expressions
+$A(\mathcal{E})=\left\{\begin{array}[]{cc}\log m+p\log p+\left(1-p\right)\log\frac{1-p}{m-1},&p\in\left[\frac{m-1}{m},1\right]\\ \frac{m-1}{m-2}\log(m-1)\left(\sqrt{p}-\sqrt{\frac{1-p}{m-1}}\right)^{2},&p\in\left[\frac{1}{m},\frac{m-1}{m}\right]\end{array}\right.$ (59)
+which agree with (43).
+Proof.
+Taking into account theorem 2, the first statement was in fact proved above in this section.
+For accessible information we use the relation (58).
+A routine calculation using (52), (37), (56) gives (45) which is the first line in (59) for moderately acute pyramids.
+For strongly acute pyramid, using (55), one gets
+$A(\mathcal{E})=\frac{m\left(1-r_{0}\right)}{m-2}\log(m-1),$
+which is the second line, taking into account that
+$1-r_{0})=(m-1)r_{1}=\frac{m-1}{m}(a-b)^{2}$
+by (34), (37). ∎
 4.4 The case of obtuse pyramids
@@
 $-\sum_{j=1}^{3}t_{j}\log t_{j}\geq 1;\quad t_{j}=\frac{2}{3}\cos^{2}\left(\alpha+\frac{2\pi j}{3}\right),$ (60)
-saturated for $\alpha=\pi/2+\pi k$.
+which is saturated for $\alpha=\pi/2+\pi k$.
 Most interesting is the case of “nearly flat” trine when $06140<r_{0}<\bar{r}_{0}\approx 0,$ 0614 .
@@
 Shor in [13], the conjectured optimal observable has 6 outcomes and is a statistical mixture of the MUD described above and the SRM $M_{k}=|{e}_{k}\rangle\langle{e}_{k}|,\,k=1,2,3.$ The SRM becomes optimal for obtuse pyramids with $\bar{r}_{0}\leq r_{0}\leq r_{1}=1/3.$
-The generalization of this picture to the case $m\geq 3$ based on extensive numerical studies was suggested and elaborated in [2].
+The generalization of this picture to the case $m\geq 3$ based on numerical studies was suggested and elaborated in [2].
 Let us apply our optimality criterion to derive the corresponding entropy inequalities and to obtain additional insight.
@@
 where
-$|\tilde{\phi}_{k}\rangle$ $=$ $t^{-1}\overline{\rho}^{1/2}|\tilde{e}_{k}\rangle=t^{-1}\left(\sqrt{r_{1}}|e_{k}\rangle+\frac{\sqrt{r_{0}}t-\sqrt{r_{1}}}{\sqrt{m}}|e_{0}\rangle\right);$
-$p_{k}$ $=$ $t^{-2}\left(r_{1}+\frac{r_{0}t^{2}-r_{1}}{m}\right);\quad k=1,\dots,m;$
+$|\tilde{\phi}_{k}\rangle$ $=$ $t^{-1}\overline{\rho}^{1/2}|\tilde{e}_{k}\rangle=t^{-1}\left(\sqrt{r_{1}}|e_{k}\rangle+\frac{\sqrt{r_{0}}t-\sqrt{r_{1}}}{\sqrt{m}}|e_{0}\rangle\right),$
+$p_{k}$ $=$ $t^{-2}\left(r_{1}+\frac{r_{0}t^{2}-r_{1}}{m}\right);\quad k=1,\dots,m,$
 and
@@
 where
-$\tilde{\mu}_{0}(p)=\frac{\sqrt{\frac{p\left(1-p\right)}{m-1}}\left(\log p-\log\frac{1-p}{m-1}\right)}{\left(\left(m-1\right)\sqrt{\frac{1-p}{m-1}}-\sqrt{p}\right)\left(\sqrt{p}+\sqrt{\frac{1-p}{m-1}}\right)},$
-$\tilde{\mu}_{1}(p)=\frac{\sqrt{p}\log p+\sqrt{\frac{1-p}{m-1}}\log\frac{1-p}{m-1}}{\left(\sqrt{p}+\sqrt{\frac{1-p}{m-1}}\right)}.$
+$\tilde{\mu}_{0}(p)=\frac{\sqrt{\frac{p\left(1-p\right)}{m-1}}\left(\log p-\log\frac{1-p}{m-1}\right)}{\left(\left(m-1\right)\sqrt{\frac{1-p}{m-1}}-\sqrt{p}\right)\left(\sqrt{p}+\sqrt{\frac{1-p}{m-1}}\right)},$ (67)
+$\tilde{\mu}_{1}(p)=\frac{\sqrt{p}\log p+\sqrt{\frac{1-p}{m-1}}\log\frac{1-p}{m-1}}{\left(\sqrt{p}+\sqrt{\frac{1-p}{m-1}}\right)}.$ (68)
 The equality in (66) is attained for $z_{1}=\sqrt{p},\,z_{j}=-\sqrt{\frac{1-p}{m-1}},\,j=2,\dots,m,$ and all the permutations of these $z_{j}$.
-Note that here $\tilde{\mu}_{0}(p)\leq 0,$ which prevents us from passing to the probabilities $t_{j}=\left|z_{j}\right|^{2}.$ In (60) the parameter $p$ varies in the range $[\max\left\{\frac{m-1}{m},p(m)\right\},1],$ where $p(m)$ is found below.
+Note that here $\tilde{\mu}_{0}(p)\leq 0,$ which prevents us from passing to the probabilities $t_{j}=\left|z_{j}\right|^{2}$.
+In (66) the parameter $p$ varies in the range $[\max\left\{\frac{m-1}{m},p(m)\right\},1],$ where $p(m)$ is found below.
 The accessible information in this case is given by the same expression (45) as in the case of moderately acute pyramids.
@@
 $r_{0}<\left(1+\frac{m-1}{\tau_{o}(m)^{2}}\right)^{-1}\equiv r_{0}(m),$
-and, via (33), to
+and, via (37) and (56), to
 $p\equiv a^{2}<\left[1+\left(m-1\right)\left(1-\frac{m}{1-\tau_{o}(m)}\right)^{-2}\right]^{-1}\equiv p(m).$ (71)
-The numerically found values:
+The numerically found values [^5] By using equation (70) it is possible to express the function $p(m)$ via the Lambert $W$-function.:
 $m$ 2 3 4 5 6 7
@@
 (39).
-In the case $m=7$, as well as for greater values of $m$, the solution of the equation (62) becomes negative and also $p(m)<\frac{m-1}{m}.$ This makes plausible the hypothesis that for $m=7,\dots$ the optimal observable is SRM, and the corresponding entropy inequality is (60) for all $p\geq\frac{m-1}{m}.$
+In the case $m=7$, as well as for greater values of $m$, the solution of the equation (70) becomes negative and also $p(m)<\frac{m-1}{m}.$ This makes plausible the hypothesis that for $m\geq 7$ the optimal observable is SRM, and the corresponding entropy inequality is (66) for all $p\geq\frac{m-1}{m}.$
 Then, for $m\leq 6$, under the condition (71) the equation (65) implies $c_{1}=1$ while from (69)
@@
 $-\sum_{j=1}^{m}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}\geq\frac{\left(\tau_{o}(m)^{-1}-1\right)}{m}\log\frac{2\left(\tau_{o}(m)-1\right)^{2}}{m\left(m-1+\tau_{o}(m)^{2}\right)}\left|\sum_{j=1}^{m}z_{j}\right|^{2}+1.$ (72)
-The equality is attained if $z_{1}=-z_{2}=1/\sqrt{2},z_{j}=0$ for $j\geq 3$ and for all permutations of these $z_{j}.$
-In this case the accessible information is
-$A(\mathcal{E})=\log m-\mathrm{Tr\,}\Lambda_{0}\bar{\rho}=\log\frac{m}{2}-r_{0}\left(\tau_{o}(m)^{-1}-1\right)\log\frac{2\left(\tau_{o}(m)-1\right)^{2}}{m\left(m-1+\tau_{o}(m)^{2}\right)}.$
+The equality is attained for $z_{1}=-z_{2}=1/\sqrt{2},z_{j}=0$, $j\geq 3$; for $z_{1}=\sqrt{p(m)},\,z_{j}=-\sqrt{\frac{1-p(m)}{m-1}},\,j\geq 2$; and for all permutations of these $z_{j}.$
+In this case the accessible information, by using (58), is
+$A(\mathcal{E})$ $=$ $\log\frac{m}{2}-r_{0}\left(\tau_{o}(m)^{-1}-1\right)\log\frac{2\left(\tau_{o}(m)-1\right)^{2}}{m\left(m-1+\tau_{o}(m)^{2}\right)}$
+$=$ $\log\dfrac{m}{2}-r_{0}m\tilde{\mu}_{0}(p(m)).$
+The following theorem summarizes the content of this section.
+Theorem 3.
+If $m\geq 7$ then $p(m)<\frac{m-1}{m}$ and the entropy inequality (66) holds for $p\in\left[\frac{m-1}{m},1\right]$.
+If $m\leq 6$ then $p(m)>\frac{m-1}{m}$ and the inequality (66) holds for $p\in\left[p(m),1\right]$, yielding the inequality (72) for $p\in\left[\frac{m-1}{m},p(m)\right]$.
+The proof of the inequalities is given in section 5.2.
+Corollary 2.
+In the case of obtuse pyramids the information-optimal observable is given by the relations (61) with $t=1$ for $m\geq 7$, and $t=\sqrt{\frac{r_{1}}{r_{0}}}\tau_{o}(m)$ for $m\leq 6$.
+The accessible information is given by the expression (45) if $p\in[\max\left\{\frac{m-1}{m},p(m)\right\},1],$ otherwise by (4.4) if $m\leq 6$ and $p\in\left[\frac{m-1}{m},p(m)\right]$.
 4.5 The case of flat pyramid
@@
 ∎
-The theorem 1 then implies
+Theorem 1 then implies
 Corollary 3.
 For the ensemble of flat pyramid, information-optimal observable is given by (61) with $t=1$ for $m\geq 7$ and $t=\infty$ for $m\leq 6$.
-The corresponding accessible information is [^5] Cf. [2], second line in eq. (33).
+The corresponding accessible information is
 $A(\mathcal{E})=\log m-\mathrm{Tr\,}\Lambda_{0}\bar{\rho}=\left\{\begin{array}[]{cc}\log\frac{m}{2},&m\leq 6;\\ \frac{m-2}{m}\log(m-1),&m\geq 7.\end{array}\right.$
@@
 and by assumption (75) holds for real $z_{j}$. ∎
-5 Proof of theorem 2
-Let $m\geq 2,\ p\geq\dfrac{m-1}{m}$ and let $\mu_{0}(p),\mu_{1}(p)$ be given by (24), (25).
-For $p\in[(m-1)/m,1]$ and a probability distribution $(t_{1},...,t_{m})$, it holds
+5 Proofs
+5.1 Proof of theorem 2
+Let $m\geq 2,\ \dfrac{m-1}{m}\leq p<1$ and let $\mu_{0}(p),\mu_{1}(p)$ be given by (28), (29).
+For $p\in[(m-1)/m,1)$ and a probability distribution $(t_{1},...,t_{m})$, it holds
 $F(t):=H(t)-\Bigl{(}\mu_{0}(p)\Bigl{(}\sum\limits_{k=1}^{m}\sqrt{t_{k}}\Bigl{)}^{2}-\mu_{1}(p)\Bigl{)}\geq 0,$ (78)
 where $H(t)=-\sum\limits_{k=1}^{m}t_{k}\log t_{k}$ is the Shannon entropy.
+In the special case $p=1$, the inequality (78) is trivially satisfied by substituting $\mu_{0}(1):=0,\ \mu_{1}(1)=0$.
 Proof.
@@
 $\mu_{0}(p,\alpha)=\dfrac{\sqrt{\dfrac{p(1-p)}{\alpha(1-\alpha)}}\Bigl{(}\log\dfrac{p}{\alpha}-\log\dfrac{1-p}{1-\alpha}\Big{)}}{\Bigl{(}\sqrt{\alpha p}+\sqrt{(1-\alpha)(1-p)}\Bigl{)}\Bigl{(}\sqrt{\dfrac{p}{\alpha}}-\sqrt{\dfrac{1-p}{1-\alpha}}\Bigl{)}},$ (79)
-$\mu_{1}(p,\alpha)=\dfrac{\sqrt{\dfrac{p}{\alpha}}\log\dfrac{p}{\alpha}-\sqrt{\dfrac{1-p}{1-\alpha}}\log\dfrac{1-p}{1-\alpha}}{\sqrt{\dfrac{p}{\alpha}}-\sqrt{\dfrac{1-p}{1-\alpha}}}.$ (71)
+$\mu_{1}(p,\alpha)=\dfrac{\sqrt{\dfrac{p}{\alpha}}\log\dfrac{p}{\alpha}-\sqrt{\dfrac{1-p}{1-\alpha}}\log\dfrac{1-p}{1-\alpha}}{\sqrt{\dfrac{p}{\alpha}}-\sqrt{\dfrac{1-p}{1-\alpha}}}$ (80)
+for real variables $p$ and $\alpha$ varying in the range $(0,1)$.
 The following lemmas lead to the proof of the main theorem.
@@
 Lemma 3.
-Let $A\geq 0,\ B,\ C\in\mathbb{R}$. Then the function
+Let $A>0,\ B,\ C\in\mathbb{R}$. Then the function
 $f(t)=h(t)-(A\sqrt{t(1-t)}+Bt+C)$ (82)
-reaches a local minimum on $[0,1]$ at no more than two points $t_{1}\leq 1/2\leq t_{2}$ which are critical.
-In this case, if $B>0$ and $t_{*}>1/2$ is critical ($B<0$ and $t_{*}<1/2$), then $t_{*}$ is a point of the global minimum.
+[…]
 Proof.
@@
 The variables $\alpha,\ p,\beta,\ t$ used in what follows, unless otherwise stated, are such that
-$\alpha\in[0,1/2],\quad p\in[1-\alpha,1],\quad\beta\in[\alpha,1/2],\quad t\in[0,1].$
+$\alpha\in{(0,1/2],\quad p\in[1-\alpha,1},\quad\beta\in[\alpha,1/2],\quad t\in[0,1].$
 Lemma 4.
@@
       $(d/dp)^{2}L(p)=-\dfrac{\sqrt{\alpha}\ln(p/\alpha)}{4p^{3/2}}-\dfrac{\sqrt{1-\alpha}\ln((1-p)/(1-\alpha))}{4(1-p)^{3/2}}\geq 0$
-      ($p\geq 1-\alpha\geq 1/2$);
+      (since $p\geq 1-\alpha\geq 1/2$);
     2. (b)
-      $L(1-\alpha)=0$,
+      $L(1-\alpha)=0$ and
       $L^{\prime}(1-\alpha)=\sqrt{\dfrac{\alpha}{1-\alpha}}-1+\dfrac{1}{2}\ln\dfrac{\alpha}{1-\alpha}+\dfrac{1}{2}\sqrt{\dfrac{\alpha}{1-\alpha}}\ln\dfrac{1-\alpha}{\alpha}\geq 0\quad\Leftrightarrow$
       $\dfrac{\alpha}{1-\alpha}\Bigl{(}1+\dfrac{1}{2}\ln\dfrac{1-\alpha}{\alpha}\Bigl{)}-\Bigl{(}1-\dfrac{1}{2}\ln\dfrac{1-\alpha}{\alpha}\Bigl{)}\geq 0\quad\Leftrightarrow$
-      $\ln\dfrac{1-\alpha}{\alpha}\geq 2-4\alpha,\quad\alpha\in(0,1/2].$
+      $\ln\dfrac{1-\alpha}{\alpha}\geq 2-4\alpha.$
     Therefore, $L(p)\geq 0$ and the inequality (85) is satisfied.
@@
     for $C\geq 1$.
-    It is simple to calculate that for $\beta=\alpha$, we have $t_{*}=p$, and $f(p)=0$. Therefore, $f(t_{*})\geq 0$ for all $\alpha\leq\beta\leq 1/2$, and $f(t)\geq 0$.
+    It is simple to calculate that for $\beta=\alpha$, we have $t_{*}=p$, and $f(p)=0$.
+Therefore, $f(t_{*})\geq 0$ for all $\alpha\leq\beta\leq 1/2$.
+We obtained the desired inequality $f(t)\geq 0$ for the minimizer $t_{*}$ and it is even more so for the other points $t\in[0,1]$.
 In the particular case of (85), where $\alpha=1/2$, one has an interesting inequality
@@
 which takes place for $p\in(0,1)$. The function $h_{1/2}(p)$ is nonnegative, convex and $h_{1/2}(p)\leq\sqrt{2}$.
+5.2 Proof of theorem 3
+This section is devoted to proving the inequality (66) which verificates the condition (ii.a) in case of obtuse pyramid.
+Namely, if non-negative parameter $p<1$ is larger than the maximum of $\dfrac{m-1}{m}$ and $p(m)$ that is defined in (71), then for complex numbers $z_{1},...,z_{m}$ forming probability distribution $(|z_{1}|^{2},...,|z_{m}|^{2})$, the inequality
+$-\sum\limits_{j=1}^{m}|z_{j}|^{2}\log|z_{j}|^{2}\geq\tilde{\mu}_{0}(p)\Bigl{|}\sum\limits_{j=1}^{m}z_{j}\Bigl{|}^{2}-\tilde{\mu}_{1}(p).$
+holds with $\tilde{\mu}_{0}(p),\tilde{\mu}_{1}(p)$ defined in (67), (68).
+For $p=1$, which corresponds to the orthonormal case, the inequality reduces to a trivial one: $H(|z_{1}|^{2},...,|z_{m}|^{2})\geq 0$.
+Proof.
+Denote $z_{j}=x_{j}+iy_{j}$ and consider the Lagrange function
+$L(x_{1},...,x_{m},y_{1},...,y_{m})=-\sum\limits_{k=1}^{m}(x_{k}^{2}+y_{k}^{2})\log(x_{k}^{2}+y_{k}^{2})-\tilde{\mu}_{0}(p)\Bigl{(}\sum\limits_{k=1}^{m}x_{k}\Bigl{)}^{2}-$ (86)
+$-\tilde{\mu}_{0}(p)\Bigl{(}\sum\limits_{k=1}^{m}y_{k}\Bigl{)}^{2}+\tilde{\mu}_{1}(p)-\lambda\Bigl{(}\sum\limits_{k=1}^{m}(x_{k}^{2}+y_{k}^{2})-1\Bigl{)}.$ (87)
+The necessary conditions are:
+$\begin{cases}-\dfrac{1}{2}\dfrac{\partial L}{\partial x_{k}}=x_{k}\Bigl{(}\log(x_{k}^{2}+y_{k}^{2})+\log e\Bigl{)}+\tilde{\mu}_{0}(p)\sum\limits_{k=1}^{m}x_{j}+\lambda x_{k},\\ -\dfrac{1}{2}\dfrac{\partial L}{\partial y_{k}}=y_{k}\Bigl{(}\log(x_{k}^{2}+y_{k}^{2})+\log e\Bigl{)}+\tilde{\mu}_{0}(p)\sum\limits_{k=1}^{m}y_{j}+\lambda y_{k}.\end{cases}$ (88)
+for all $1\leq k\leq m$.
+Hence, we get
+$\dfrac{1}{2}\Bigl{(}y_{k}\dfrac{\partial L}{\partial x_{k}}-x_{k}\dfrac{\partial L}{\partial y_{k}}\Bigl{)}=\tilde{\mu}_{0}(p)\Bigl{(}x_{k}\sum\limits_{j=1}^{m}y_{j}-y_{k}\sum\limits_{j=1}^{m}x_{j}\Bigl{)}=0.$ (89)
+Case 1. $\sum\limits_{j=1}^{m}z_{j}=0$ or $\tilde{\mu}_{0}(p)=0$.
+We get the inequality $-\sum\limits_{j=1}^{m}|z_{j}|^{2}\log|z_{j}|^{2}\geq-\tilde{\mu}_{1}(p)$. Taking into account proposition 2 for the flat pyramid, it is sufficient to prove the lemma
+Lemma 5.
+Let $m\geq 2$ be a natural number.
+  1. 1.
+    If $\dfrac{m-1}{m}\leq p<1$, then $-\tilde{\mu}_{1}(p)\leq-\dfrac{m-2}{m}\log(m-1)+\log m$;
+  2. 2.
+    if $m\leq 6$ and $p(m)\leq p<1$, then $-\tilde{\mu}_{1}(p)\leq 1$.
+Case 2.
+$\sum\limits_{j=1}^{m}z_{j}\neq 0$.
+Then the condition (89) reduces to proportionality of two vectors $(x_{1},...,x_{m})$ and $(y_{1},...,y_{m})$.
+Due to the symmetry of (86) under multiplication by a scalar $e^{i\varphi},\ \varphi\in\mathbb{R}$, the problem reduces to the case of real variables $\{z_{j}\}$.
+For the real variables $\{z_{j}\}$ let us consider the new Lagrange function
+$L_{(1)}(z_{1},...,z_{k})=-\sum\limits_{k=1}^{m}(z_{k}^{2})\log(z_{k}^{2})-\tilde{\mu}_{0}(p)\Bigl{(}\sum\limits_{k=1}^{m}z_{k}\Bigl{)}^{2}+\tilde{\mu}_{1}(p)-\lambda\Bigl{(}\sum\limits_{k=1}^{m}z_{k}^{2}-1\Bigl{)}.$
+Thus, the necessary condition is
+$-\dfrac{1}{2}\dfrac{\partial L_{(1)}}{\partial z_{k}}=z_{k}(\log z_{k}^{2}+\log e)+\tilde{\mu}_{0}(p)\Bigl{(}\sum\limits_{j=1}^{m}z_{j}\Bigl{)}+\lambda z_{k}=0,\quad 1\leq k\leq m,$
+and the set $\{z_{j}\}$ has at most $3$ different values $c_{1},c_{2},c_{3}$ with multiplicities $k_{1},k_{k},k_{3}$, since the equation
+$z(\log z^{2}+\log e)+\tilde{\mu}_{0}(p)\Bigl{(}\sum\limits_{j=1}^{m}z_{j}\Bigl{)}+\lambda z=0$
+for a variable $z$ has at most $3$ real solutions.
+The fact the the equation of the form $z\log z^{2}+az+b=0$ for real $a$ and $b$ has no more then $3$ real roots was actually used in proof of proposition 2.
+To reduce the number of different values to two, we will introduce a new Lagrange function $L_{(2)}$.
+The critical points of $L_{(1)}$ are related with critical points of the Lagrange function
+$L_{(2)}(c_{1},c_{2},c_{3},k_{1},k_{2},k_{3})=-\sum\limits_{j=1}^{3}k_{j}c_{j}^{2}\log c_{j}^{2}-\tilde{\mu}_{0}(p)\Bigl{(}\sum\limits_{j=1}^{m}k_{j}c_{j}\Bigl{)}^{2}-$
+$-\lambda\Bigl{(}\sum\limits_{j=1}^{3}k_{j}c_{j}^{2}-1\Bigl{)}-\nu\Bigl{(}\sum\limits_{j=1}^{3}k_{j}-m\Bigl{)}.$
+New necessary conditions are
+$-\dfrac{1}{2}\dfrac{\partial L_{(2)}}{\partial c_{l}}=k_{l}c_{l}(\log c_{l}^{2}+\log e)+\tilde{\mu}_{0}(p)k_{l}\Bigl{(}\sum\limits_{j=1}^{3}k_{j}c_{j}\Bigl{)}+\lambda k_{l}c_{l}=0,$
+$-\dfrac{\partial L_{(2)}}{\partial k_{l}}=c_{l}^{2}\log c_{l}^{2}+2\tilde{\mu}_{0}(p)c_{l}\Bigl{(}\sum\limits_{j=1}^{3}k_{j}c_{j}\Bigl{)}+\lambda c_{l}^{2}+\nu=0$
+for all $1\leq l\leq 3$.
+By transformation, we have
+$-\dfrac{1}{2}c_{l}\dfrac{\partial L_{(2)}}{\partial c_{l}}+k_{l}\dfrac{\partial L_{(2)}}{\partial k_{l}}=k_{l}c_{l}\Bigl{(}c_{l}\log e-\tilde{\mu}_{0}(p)\sum\limits_{j=1}^{3}k_{j}c_{j}\Bigl{)}-\nu k_{l}=0.$
+Consider the next subcases.
+Case 2.1. $k_{1},k_{2},k_{3}\neq 0$.
+For all $1\leq l\leq 3$ it holds $c_{l}^{2}\log e=\tilde{\mu}_{0}(p)\Bigl{(}\sum\limits_{j=1}^{3}k_{j}c_{j}\Bigl{)}+\nu$, which can take place, if the set $\{c_{1},c_{2},c_{3}\}$ has no more than $2$ different values.
+This can be regarded as the next case.
+Case 2.2. $k_{3}=0$.
+Up to permutation, $(z_{j})_{j=1}^{m}=(\underbrace{c_{1},...,c_{1}}_{k_{1}\text{ times}},\underbrace{c_{2},...,c_{2}}_{k_{2}\text{ times}})$, $c_{1}\geq 0,c_{2}\in\mathbb{R}$.
+Denote
+$\tilde{\mu}_{0}(p,\alpha)=\dfrac{\sqrt{\dfrac{p(1-p)}{\alpha(1-\alpha)}}\Bigl{(}\log\dfrac{p}{\alpha}-\log\dfrac{1-p}{1-\alpha}\Bigl{)}}{\Bigl{(}\sqrt{(1-p)(1-\alpha)}-\sqrt{p\alpha}\Bigl{)}\Bigl{(}\sqrt{\dfrac{p}{\alpha}}+\sqrt{\dfrac{1-p}{1-\alpha}}\Bigl{)}},$
+$\tilde{\mu}_{1}(p,\alpha)=\dfrac{\sqrt{\dfrac{p}{\alpha}}\log\dfrac{p}{\alpha}+\sqrt{\dfrac{1-p}{1-\alpha}}\log\dfrac{1-p}{1-\alpha}}{\sqrt{\dfrac{p}{\alpha}}+\sqrt{\dfrac{1-p}{1-\alpha}}},$
+where $\alpha$ and $p$ lie on $(0,1)$.
+Relation with $\tilde{\mu}_{0}(p)$ and $\tilde{\mu}_{1}(p)$ is given by
+$\tilde{\mu}_{0}(p)=\dfrac{1}{m}\tilde{\mu}_{0}(p,1/m),$
+$\tilde{\mu}_{1}(p)=\tilde{\mu_{1}}(p,1/m)-\log m.$
+It is easy to see that for $\alpha$ and $p$ such that $10<\alpha\leq 1/2\leq 1-\alpha\leq p<$ 1 one has $\log\dfrac{p}{\alpha}-\log\dfrac{1-p}{1-\alpha}\geq 0$ and ${\sqrt{(1-p)(1-\alpha)}-\sqrt{p\alpha}\leq 0}$, therefore, $\tilde{\mu}_{0}(p,\alpha)\leq 0$.
+The initial inequality (66) for $k_{1}c_{1}^{2}=t$ is equivalent to
+$-t\log\dfrac{t}{k_{1}/m}-(1-t)\log\dfrac{1-t}{k_{2}/m}\geq\tilde{\mu}_{0}(p,1/m)\Bigl{(}\sqrt{t(k_{1}/m)}\pm\sqrt{(1-t)(k_{2}/m)}\Bigl{)}^{2}-\tilde{\mu}_{1}(p,1/m),$ (90)
+if ±$\pm c_{2}\geq 0$.
+Since $\tilde{\mu}_{0}(p,\alpha)\leq 0$,
+$\tilde{\mu}_{0}(p,1/m)\Bigl{(}\sqrt{t(k_{1}/m)}+\sqrt{(1-t)(k_{2}/m)}\Bigl{)}^{2}-\tilde{\mu}_{1}(p,1/m)\leq$
+$\leq\tilde{\mu}_{0}(p,1/m)\Bigl{(}\sqrt{t(k_{1}/m)}-\sqrt{(1-t)(k_{2}/m)}\Bigl{)}^{2}-\tilde{\mu}_{1}(p,1/m),$
+and the statement follows from lemma 6.
+Proof of lemma 5.
+First we will prove that the function $\tilde{\mu}_{1}(p)$ is increasing for $p$ in the range $\Bigl{[}\dfrac{m-1}{m},1\Bigr{)}$. The derivative is
+$\tilde{\mu}_{1}^{\prime}(p)=\log e\dfrac{2(m-2)\sqrt{\dfrac{p(1-p)}{m-1}}-4p+2+\ln p-\ln\dfrac{1-p}{m-1}}{2(m-1)\sqrt{\dfrac{p(1-p)}{m-1}}\Bigg{(}\sqrt{p}+\sqrt{\dfrac{1-p}{m-1}}\Bigg{)}^{2}}.$
+Here $2(m-2)\sqrt{\dfrac{p(1-p)}{m-1}}\geq 0$ and $\ln\dfrac{p}{1-p}-4p+2\geq 0$, therefore, $\tilde{\mu}_{1}^{\prime}(p)\geq 0$.
+So, for $\frac{m-1}{m}\leq p<1$ we have
+$-\tilde{\mu}_{1}(p)\leq-\tilde{\mu}_{1}(1-1/m)=-\dfrac{m-2}{m}\log(m-1)+\log m.$
+For $p(m)\leq p<1$ we have
+$-\tilde{\mu}_{1}(p)\leq-\tilde{\mu}_{1}(p(m))=1,$
+because the condition (71) is equivalent to $t>1$, under which $\tilde{\mu}_{1}\equiv-c_{1}=-1$.
+Lemma 6.
+Let $0\leq\alpha\leq\beta\leq\dfrac{1}{2}\leq 1-\alpha\leq p\leq 1$, $0\leq t\leq 1$. Then
+$-t\log\dfrac{t}{\beta}-(1-t)\log\dfrac{1-t}{1-\beta}\geq\tilde{\mu}_{0}(p,\alpha)\Bigl{(}\sqrt{\beta t}-\sqrt{(1-\beta)(1-t)}\Bigl{)}^{2}-\tilde{\mu}_{1}(p,\alpha).$
+Proof.
+We will prove that the inequality
+$\mu_{0}(p,\alpha)\Bigl{(}\sqrt{\beta t}+\sqrt{(1-\beta)(1-t)}\Bigl{)}^{2}-\mu_{1}(p,\alpha)\geq\tilde{\mu}_{0}(p,\alpha)\Bigl{(}\sqrt{\beta t}-\sqrt{(1-\beta)(1-t)}\Bigl{)}^{2}-\tilde{\mu}_{1}(p,\alpha)$
+holds in this range. So, the statement of the lemma follows from lemma 4.
+Collecting similar terms,
+$f(t):=2\sqrt{\beta(1-\beta)}\Bigl{(}\mu_{0}(p,\alpha)+\tilde{\mu}_{0}(p,\alpha)\Bigl{)}\sqrt{t(1-t)}-(1-2\beta)\Bigl{(}\mu_{0}(p,\alpha)-\tilde{\mu}_{0}(p,\alpha)\Bigl{)}t+$
+$+\Bigl{[}(1-\beta)\Bigl{(}\mu_{0}(p,\alpha)-\tilde{\mu}_{0}(p,\alpha)\Bigl{)}-\Bigl{(}\mu_{1}(p,\alpha)-\tilde{\mu}_{1}(p,\alpha)\Bigl{)}\Bigl{]}\geq 0.$
+One can see that
+$\mu_{0}(p,\alpha)-\tilde{\mu}_{0}(p,\alpha)=2(2p-1)\dfrac{\sqrt{p(1-p)\alpha(1-\alpha)}\Bigl{(}\log\dfrac{p}{\alpha}-\log\dfrac{1-p}{1-\alpha}\Bigl{)}}{(p-\alpha)(p+\alpha-1)},$
+$\mu_{0}(p,\alpha)+\tilde{\mu}_{0}(p,\alpha)=-2(1-2\alpha)p(1-p)\dfrac{\log\dfrac{p}{\alpha}-\log\dfrac{1-p}{1-\alpha}}{(p-\alpha)(p+\alpha-1)}\leq 0,$
+$\mu_{1}(p,\alpha)-\tilde{\mu}_{1}(p,\alpha)=2\dfrac{\sqrt{p(1-p)\alpha(1-\alpha)}\Bigl{(}\log\dfrac{p}{\alpha}-\log\dfrac{1-p}{1-\alpha}\Bigl{)}}{p-\alpha}.$
+The function $f(t)$ has the form $A\sqrt{t(1-t)}+Bt+C$ with $A\leq 0$ and $B\leq 0$, then the unique minimum point $t_{*}\in[1/2,1]$.
+it has the value $t_{*}=\dfrac{1}{2}\Bigl{(}1-\dfrac{B}{\sqrt{A^{2}+B^{2}}}\Bigl{)}$ and the minimum of the function $f(t)$ is equal to
+$f(t_{*})=\dfrac{B}{2}+C-\dfrac{1}{2}\sqrt{A^{2}+B^{2}},$
+where
+$A=2\sqrt{\beta(1-\beta)}\Bigl{(}\mu_{0}(p,\alpha)+\tilde{\mu}_{0}(p,\alpha)\Bigl{)},$
+$B=-(1-2\beta)\Bigl{(}\mu_{0}(p,\alpha)-\tilde{\mu}_{0}(p,\alpha)\Bigl{)},$
+$C=\Bigl{[}(1-\beta)\Bigl{(}\mu_{0}(p,\alpha)-\tilde{\mu}_{0}(p,\alpha)\Bigl{)}-\Bigl{(}\mu_{1}(p,\alpha)-\tilde{\mu}_{1}(p,\alpha)\Bigl{)}\Bigl{]}.$
+One has
+$\dfrac{B}{2}+C=\dfrac{\sqrt{p(1-p)\alpha(1-\alpha)}\Bigl{(}\log\dfrac{p}{\alpha}-\log\dfrac{1-p}{1-\alpha}\Bigl{)}}{(p-\alpha)(p+\alpha-1)}\times$
+×$\times\Bigl{(}(2p-1)-2(p+\alpha-1)\Bigl{)}=$
+$=(1-2\alpha)\dfrac{\sqrt{p(1-p)\alpha(1-\alpha)}\Bigl{(}\log\dfrac{p}{\alpha}-\log\dfrac{1-p}{1-\alpha}\Bigl{)}}{(p-\alpha)(p+\alpha-1)}\geq 0,$
+$\dfrac{1}{2}\sqrt{A^{2}+B^{2}}=(1-2\alpha)\dfrac{\sqrt{p(1-p)\alpha(1-\alpha)}\Bigl{(}\log\dfrac{p}{\alpha}-\log\dfrac{1-p}{1-\alpha}\Bigl{)}}{(p-\alpha)(p+\alpha-1)}\times$
+×$\times\sqrt{4\ \dfrac{p(1-p)\beta(1-\beta)}{\alpha(1-\alpha)}+\dfrac{(1-2p)^{2}(1-2\beta)^{2}}{(1-2\alpha)^{2}}}$
+Thus,
+$\dfrac{\sqrt{A^{2}+B^{2}}}{B+2C}=\sqrt{4\ \dfrac{p(1-p)\beta(1-\beta)}{\alpha(1-\alpha)}+\dfrac{(1-2p)^{2}(1-2\beta)^{2}}{(1-2\alpha)^{2}}}\leq 1$ (91)
+Let us denote $u(t)=4t(1-t)$. The inequality (91) can be rewritten in the form
+$\dfrac{u(\beta)u(p)}{u(\alpha)}+\dfrac{(1-u(\beta))(1-u(p))}{1-u(\alpha)}\leq 1,$
+where $u(p)\leq u(\alpha)\leq u(\beta)$. The linear function
+$t\mapsto g(t)=\dfrac{u(\beta)t}{u(\alpha)}+\dfrac{(1-u(\beta))(1-t)}{1-u(\alpha)}$
+has non-negative derivative $\dfrac{u(\beta)}{u(\alpha)}-\dfrac{1-u(\beta)}{1-u(\alpha)}=\dfrac{u(\beta)-u(\alpha)}{u(\alpha)(1-u(\alpha))}$ and $g(u(\alpha))=1$.
+Therefore, the inequality (91) holds, and lemma has been proven.
+∎
 Acknowledgment. This work is supported by Russian Science Foundation under the grant No 24-11-00145. The authors are grateful to S.V. Grishin and D.A. Kronberg for providing numerical simulations.

### v3

@@
 The recently obtained optimality criterion [7] when applied to specific ensembles of states leads to nontrivial tight entropy inequalities that are discrete relatives of the famous log-Sobolev inequality.
-In this light, the hypothesis of globally information-optimal measurement for an ensemble of equiangular equiprobable states [2] is reconsidered and the corresponding entropy inequalities are proposed and proved.
-Via the optimality criterion, this provides also the first proof of the conjecture concerning globally information-optimal observables for quantum pyramids put forward in [2].
+In this light, the hypothesis of globally information-optimal measurement for an ensemble of equiangular equiprobable states (quantum pyramids) [2] is reconsidered and the corresponding entropy inequalities are proposed.
+Via the optimality criterion, this suggests also a proof of the conjecture concerning globally information-optimal observables for quantum pyramids put forward in [2].
 1
 Introduction
-The problem of accessible information – the maximal information that can be extracted from measurements over an ensemble of quantum states – is a basic problem in quantum information and communication theory, coming back to its origin, and important in particular for applications in quantum cryptography, see e.g.
-the survey [14].
+The problem of accessible information – the maximal information that can be extracted from measurements over an ensemble of quantum states – is a basic problem in quantum information and communication theory, coming back to its origin, see e.g.
+[6], [16], [9], [11], and important in particular for applications in quantum cryptography.
 The problem was solved in some special (notably, symmetric) cases but in general remains open.
@@
 (76)) serving to determine the critical values of the parameter $t$ of the optimal observable corresponding to transition from moderately to strongly acute (resp, obtuse) pyramids.
-In the subsection 4.5 we derive and prove the tight entropy inequality (75) implying the global optimality of the hypothetical observable from [2] (proposition 2 and corollary 3) in the case of flat pyramids.
-The findings in subsections 4.4, 4.5 also emphasize that for obtuse and flat pyramids in the dimensions $m\geq 7$ the sharp “square root” observable (SRM) outperforms the “unambiguous discrimination” observable (UDM) and is the optimal one.
-The proofs of the inequality (27), given in subsection 5.1, as well as of the inequalities (66), (72) in subsection 5.2 was obtained by the second author as a response to the hypotheses put forward by the first author.
-Via the optimality criterion of section 2, it also implies the validity of the conjectures concerning accessible information for quantum pyramids, formulated and numerically substantiated in [2] (corollary 1 to theorem 2, corollary 2 to theorem 3).
+In the subsection 4.5 we derive the tight entropy inequality (82) implying the global optimality of the hypothetical observable from [2] (proposition 2 and corollary 3) in the case of flat pyramids.
+The findings in subsections 4.4, 4.5 also emphasize that for obtuse and flat pyramids in the dimensions $m\geq 7$ the sharp “square root” observable (SRM) outperforms the “unambiguous discrimination” observable (UDM) and tentatively is the optimal one.
+The proposed entropy inequalities may be more convenient for numerical check than the direct optimization, however proving them analytically is a challenge.
+In section 5, the suggestions for proofs proposed by the second author as a response to the hypotheses (27), (67), (79) put forward by the first author, are presented.
+Via the optimality criterion of section 2, this suggests also a proof of the conjecture concerning globally information-optimal observables for quantum pyramids formulated and numerically substantiated in [2].
 2
@@
 Let $\mathcal{H}$ be a finite dimensional Hilbert space, $\mathfrak{S}=\mathfrak{S}(\mathcal{H})$ the convex set of quantum states (density operators on $\mathcal{H}$).
-Ensemble of quantum states is a collection =$\mathcal{E}=\left\{\pi_{j},\rho_{j}\right\}_{j=1,\dots,m}$ where $\rho_{j}\in\mathfrak{S,}$ and $\pi=\left\{\pi_{j}\right\}$ is a probability distribution.
-Let =$\mathcal{M}=\left\{M_{k}\right\}_{k=1,\dots,n}$ be an observable i.e.
+Ensemble of quantum states is a collection $\mathcal{E}=\left\{\pi_{j},\rho_{j}\right\}_{j=1,\dots,m}$ where $\rho_{j}\in\mathfrak{S,}$ and $\pi=\left\{\pi_{j}\right\}$ is a probability distribution.
+Let $\mathcal{M}=\left\{M_{k}\right\}_{k=1,\dots,n}$ be an observable i.e.
 a collection of operators on $\mathcal{H}$ such that $M_{k}\geq 0,\,\sum_{k=1}^{n}M_{k}=I$ (unit operator).
@@
 In particular, the supremum can be replaced with the maximum.
-Introducing the dual ensemble =$\mathcal{E}^{\prime}=\left\{p_{k},\sigma_{k}\right\}$ with
+Introducing the dual ensemble $\mathcal{E}^{\prime}=\left\{p_{k},\sigma_{k}\right\}$ with
 $\sigma_{k}=p_{k}^{-1}\overline{\rho}^{1/2}M_{k}\overline{\rho}^{1/2}=\left|\phi_{k}\rangle\langle\phi_{k}\right|,$
@@
 The entropy inequality
-$-\sum_{j}\left\langle\psi|M_{j}^{\prime}|\psi\right\rangle\log\left\langle\psi|M_{j}^{\prime}|\psi\right\rangle\mathrm{\,}\,\geq\left\langle\psi|\Lambda_{0}|\psi\right\rangle,$ (7)
+[…]
 holds for all unit vectors $\psi\in\mathcal{H};$
@@
 The unit vectors $|\phi_{k}^{0}\rangle$ turn (7) into equality:
-$-\sum_{j}\left\langle\phi_{k}^{0}|M_{j}^{\prime}|\phi_{k}^{0}\right\rangle\log\left\langle\phi_{k}^{0}|M_{j}^{\prime}|\phi_{k}^{0}\right\rangle=\left\langle\phi_{k}^{0}|\Lambda_{0}|\phi_{k}^{0}\right\rangle,\quad k=1,\dots.n.$ (8)
+[…]
 The value of the accessible information (4) is
@@
 Taking into account (3), the condition (ii.a) becomes
-$-\sum_{j}\left\langle\psi|M_{j}^{\prime}|\psi\right\rangle\log\mathrm{Tr\,}\,\sigma M_{j}^{\prime}\geq\left\langle\psi|\Lambda_{0}|\psi\right\rangle,\quad\psi\in\mathcal{H},$
+$\sum_{j}\left\langle\psi|M_{j}^{\prime}|\psi\right\rangle\log\mathrm{Tr\,}\,\sigma M_{j}^{\prime}\geq\left\langle\psi|\Lambda_{0}|\psi\right\rangle,\quad\psi\in\mathcal{H},$
 for all $\sigma\in\mathfrak{S.}$ Without loss of generality we can assume that $\left\|\psi\right\|=1$ here. Then for fixed $\psi$ the minimal value of the lefthand side is
-$-\sum_{j}\left\langle\psi|M_{j}^{\prime}|\psi\right\rangle\log\left\langle\psi|M_{j}^{\prime}|\psi\right\rangle$
+[…]
 since the difference is just the relative entropy of two probability distributions.
@@
 by necessity form an orthonormal basis. Then the entropy inequality (7) takes the form
-$-\sum_{j=1}^{d}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}\geq\sum_{j=1}^{d}\lambda_{jk}\overline{z_{j}}z_{k},$ (16)
+$\sum_{j=1}^{d}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}\geq\sum_{j=1}^{d}\lambda_{jk}\overline{z_{j}}z_{k},$ (16)
 where $\lambda_{jk}=\langle e_{j}|\Lambda_{0}|e_{k}\rangle$ and $z_{j}=\langle e_{j}|\psi\rangle$ are complex variables subject only to the constraint $\sum_{j=1}^{d}\left|z_{j}\right|^{2}=1$.
@@
     for a problem which has a hypothetical solution, the conditions can be used to verify its actual optimality.
-One inserts the conjectured $p_{k}^{0},\sigma_{k}^{0}$ into (11) to compute Hermitean $\Lambda_{0}$.
+One inserts the conjectured $p_{k}^{0},\sigma_{k}^{0}$ into (11) to compute Hermitean $\Lambda_{0}\$.
 Then the main difficulty will be proof of the entropy inequality (7).
@@
 3 Example: two pure states
-[…]
+The case of ensemble of two states was treated previously in [12], [10].
+For simplicity of demonstration we take two pure equiprobable states.
+Let $|e_{0}\rangle,|e_{1}\rangle$ be an orthonormal basis in two-dimensional Hilbert space.
+Consider ensemble $\mathcal{E}=\left\{\pi_{\pm},\rho_{\pm}\right\},$ where $\pi_{\pm}=\frac{1}{2},\rho_{\pm}=\left|\psi_{\pm}\rangle\langle\psi_{\pm}\right|$ with $|\psi_{\pm}\rangle=\cos\alpha/2|e_{0}\rangle\pm\sin\alpha/2|e_{1}\rangle,\,0\leq\alpha\leq\pi.$ As argued in [12] the accesible information
 $A(\mathcal{E})=1-h\left(\frac{1+\sin\alpha}{2}\right)$ (17)
-is attained by the optimal observable =$\mathcal{M=}\left\{M_{+},M_{-}\right\}$ where $M_{\pm}=\left|e_{\pm}\rangle\langle e_{\pm}\right|$ and $|e_{\pm}\rangle=\frac{1}{\sqrt{2}}\left(|e_{0}\rangle\pm|e_{1}\rangle\right)$ is an orthonormal basis [^3] Strictly speaking, the paper [10] contains the check of local optimality..
+is attained by the optimal observable $\mathcal{M=}\left\{M_{+},M_{-}\right\}$ where $M_{\pm}=\left|e_{\pm}\rangle\langle e_{\pm}\right|$ and $|e_{\pm}\rangle=\frac{1}{\sqrt{2}}\left(|e_{0}\rangle\pm|e_{1}\rangle\right)$ is an orthonormal basis [^3] Strictly speaking, the paper [12] contains the check of local optimality..
 Here
@@
 Incerting $\sigma_{\pm}^{0}=\left|\psi_{\pm}\rangle\langle\psi_{\pm}\right|$ into
-$K(\sigma)=-\sum_{\pm}\left|e_{\pm}\rangle\langle e_{\pm}\right|\log\left\langle e_{\pm}|\sigma|e_{\pm}\right\rangle$
+[…]
 we obtain
@@
 $\frac{\lambda_{1}(p)-\lambda_{0}(p)}{2}=\frac{p\log p-\left(1-p\right)\log\left(1-p\right)}{2p-1}.$
-The inequality (23) becomes equality for $t=p$ and $t=1-p.$
+The inequality (23) becomes equality for $t=p$ and t=1p.t=1-p.
 In these relations $p>1/2;$ taking the limit $p\rightarrow 1/2$ the inequality (23) becomes
@@
 is monotonously decreasing on $[1/2,1],$ which can be seen by showing its derivative is negative on $(1/2,1]$ (for this it is convenient to make a monotonic change of variable $u=\sqrt{t/\left(1-t\right)}$).
-Thus $h(t)-g(t,p)=0$ for $t=p,$ and $\frac{d}{dt}\left[h(t)-g(t,p)\right]$ has the same sign as $t-p.$ Hence $h(t)-g(t,p)\geq 0.$
-Theorem 1 then implies the global optimality of the observable =$\mathcal{M=}\left\{M_{+},M_{-}\right\}.$ Also
+Thus h(t)g(t,p)=0h(t)-g(t,p)=0 for $t=p,$ and $\frac{d}{dt}\left[h(t)-g(t,p)\right]$ has the same sign as tp.t-p. Hence $h(t)-g(t,p)\geq 0.$
+Theorem 1 then implies the global optimality of the observable $\mathcal{M=}\left\{M_{+},M_{-}\right\}.$ Also
 $\mathrm{Tr\,}\overline{\rho}\Lambda_{0}=\lambda_{0}(p)\cos^{2}\alpha/2-\lambda_{1}(p)\sin^{2}\alpha/2=h(p)$
@@
 For integer $m\geq 2$ and $p>\frac{m-1}{m}$
-$-\sum_{j=1}^{m}t_{j}\log t_{j}\geq\mu_{0}(p)\left(\sum_{j=1}^{m}\sqrt{t_{j}}\right)^{2}-\mu_{1}(p),$ (27)
+$\sum_{j=1}^{m}t_{j}\log t_{j}\geq\mu_{0}(p)\left(\sum_{j=1}^{m}\sqrt{t_{j}}\right)^{2}-\mu_{1}(p),$ (27)
 where $t_{j}\geq 0,\sum_{j=1}^{m}t_{j}=1$ and
@@
 The inequality (27) arises in connection with the problem of accessible information for ensemble of $m$ equiprobable equiangular pure states discussed in the next subsection.
-The proof of (27) will be given in section 5.1.
+The suggestion for a proof of (27) will be given in section 5.1.
 Note that in the case $m=2$ the inequality (27) reduces to (23).
@@
 It can be compared with the lower bound of theorem 3.11 in [13]:
-$-\sum_{j=1}^{m}t_{j}\log t_{j}\geq\log m-\frac{m\log m}{2\left(m-1\right)}D(P;P_{U}),$ (32)
+$\sum_{j=1}^{m}t_{j}\log t_{j}\geq\log m-\frac{m\log m}{2\left(m-1\right)}D(P;P_{U}),$ (32)
 where $D(P;P_{U})=\sum_{j=1}^{m}\left|t_{j}-1/m\right|$ is the total variation distance. There is well-known relation for any two distributions $P,P^{\prime}$:
@@
 The extreme cases are $r_{1}=0\,\left(a=b=\sqrt{\frac{1}{m}}\right)$ – the ort $|\psi_{j}\rangle\equiv|e_{0}\rangle,$ and $r_{0}=0\,\left(a=\sqrt{\frac{m-1}{m}},b=-\sqrt{\frac{1}{m\left(m-1\right)}}\right)$ – the flat pyramid.
-We will consider the ensemble =$\mathcal{E}=\left\{\pi_{j},\rho_{j}\right\}_{j=1,\dots,m}$ with $\pi_{j}=\frac{1}{m},\,\rho_{j}=|\psi_{j}\rangle\langle\psi_{j}|$ with the average state
+We will consider the ensemble $\mathcal{E}=\left\{\pi_{j},\rho_{j}\right\}_{j=1,\dots,m}$ with $\pi_{j}=\frac{1}{m},\,\rho_{j}=|\psi_{j}\rangle\langle\psi_{j}|$ with the average state
 $\bar{\rho}\equiv\frac{1}{m}\sum_{j=1}^{m}|\psi_{j}\rangle\langle\psi_{j}|=r_{1}I+\left(r_{0}-r_{1}\right)|e_{0}\rangle\langle e_{0}|.$ (38)
@@
 In what follows we consider the solution for the accessible information $A(\mathcal{E})$ and its optimizing observable conjectured and supported numerically in [2].
-In the rest of this subsection and in the next subsection we consider acute pyramids for which $r_{0}\geq r_{1}$ or, equivalently, $b\geq 0.$ According to [2] the conjectured optimal observable belongs to the family depending on the parameter $0\leq t\leq 1$ and has the form =$\mathcal{M}=\left\{M_{k};k=0,1,\dots,m\right\}$ where $M_{k}=|\tilde{e}_{k}\rangle\langle\tilde{e}_{k}|$
+In the rest of this subsection and in the next subsection we consider acute pyramids for which $r_{0}\geq r_{1}$ or, equivalently, $b\geq 0.$ According to [2] the conjectured optimal observable belongs to the family depending on the parameter $0\leq t\leq 1$ and has the form $\mathcal{M}=\left\{M_{k};k=0,1,\dots,m\right\}$ where $M_{k}=|\tilde{e}_{k}\rangle\langle\tilde{e}_{k}|$
 $|\tilde{e}_{k}\rangle=|e_{k}\rangle+\frac{t-1}{\sqrt{m}}|e_{0}\rangle;\quad k=1,\dots,m,$ (41)
@@
 hence
-$K\left(\sigma\right)=-\sum_{j}|e_{j}\rangle\langle e_{j}|\,\log\mathrm{\,}\,\langle e_{j}|\sigma|e_{j}\rangle.$
-From (41), (39) we obtain the dual ensemble =$\mathcal{E}^{\prime}=\left\{p_{k},\sigma_{k}\right\}$ with
+[…]
+From (41), (39) we obtain the dual ensemble $\mathcal{E}^{\prime}=\left\{p_{k},\sigma_{k}\right\}$ with
 $\sigma_{k}=p_{k}^{-1}|\tilde{\phi}_{k}\rangle\langle\tilde{\phi}_{k}|\,,\quad p_{k}=\langle\tilde{\phi}_{k}|\tilde{\phi}_{k}\rangle$
@@
 $f(s)$ $\equiv$ $\ln\left[1+\left(m-1\right)s^{2}\right]$
--- $0\displaystyle 2\frac{1+\left(m-1\right)s}{m}\ln\left[1+\left(m-1\right)s\right]-2\frac{m-1}{m}\left(1-s\right)\ln\left|1-s\right|=$ 0 .
+- $2\frac{1+\left(m-1\right)s}{m}\ln\left[1+\left(m-1\right)s\right]-2\frac{m-1}{m}\left(1-s\right)\ln\left|1-s\right|=0.$
 At the point $s=0$ (corresponding to $\tau\rightarrow\infty$) this function and its first and second derivatives vanish, while $f^{\prime\prime\prime}(s)=2\left(m-1\right)\left(m-2\right)>0$ hence $f$ is increasing and positive for small $s$.
@@
 Introducing complex numbers $z_{j}=\langle e_{j}|\psi\rangle,$ so that $\sum_{j=1}^{m}\left|z_{j}\right|^{2}=1,$ the optimality condition (ii.a) amounts to the inequality
-$-\sum_{j=1}^{m}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}\geq\mu_{0}(p)\left|\sum_{j=1}^{m}z_{j}\right|^{2}-\mu_{1}(p).$ (57)
-Since $\mu_{0}(p)\geq 0,$ this is equivalent to (27) where $t_{j}=\left|z_{j}\right|^{2}$
+$\sum_{j=1}^{m}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}\geq\mu_{0}(p)\left|\sum_{j=1}^{m}z_{j}\right|^{2}-\mu_{1}(p).$ (57)
+Since $\mu_{0}(p)\geq 0,$ this is equivalent to (27) where $t_{j}=\left|z_{j}\right|^{2}$ .
 Remarkably, inserting the limiting value $p=\frac{m-1}{m}$ we obtain for $m>2$
@@
 For ensemble of acute pyramids the information-optimal observable is given by the equations (41), (42) as conjectured in [2]. The accessible information is given by the expressions
-$A(\mathcal{E})=\left\{\begin{array}[]{cc}\log m+p\log p+\left(1-p\right)\log\frac{1-p}{m-1},&p\in\left[\frac{m-1}{m},1\right]\\ \frac{m-1}{m-2}\log(m-1)\left(\sqrt{p}-\sqrt{\frac{1-p}{m-1}}\right)^{2},&p\in\left[\frac{1}{m},\frac{m-1}{m}\right]\end{array}\right.$ (59)
+$A(\mathcal{E})=\left\{\begin{array}[]{cc}\log m+p\log p+\left(1-p\right)\log\frac{1-p}{m-1},&p\in\left[\frac{m-1}{m},1\right]\\ \frac{m-1}{m-2}\log(m-1)\left(\sqrt{p}-\sqrt{\frac{1-p}{m-1}}\right)^{2},&p\in\left[\frac{1}{m},\frac{m-1}{m}\right]\end{array}\right$.
+(59)
 which agree with (43).
 Proof.
-Taking into account theorem 2, the first statement was in fact proved above in this section.
 For accessible information we use the relation (58).
@@
 which is the second line, taking into account that
-$1-r_{0})=(m-1)r_{1}=\frac{m-1}{m}(a-b)^{2}$
+$1-r_{0}=(m-1)r_{1}=\frac{m-1}{m}(a-b)^{2}$
 by (34), (37). ∎
@@
 [5].
-Then the optimal observable is =$\mathcal{M}=\left\{M_{k};k=1,2,3\right\},$ where $M_{k}=\frac{2}{3}|\psi_{k}^{\perp}\rangle\langle\psi_{k}^{\perp}|,$ and $|\psi_{k}^{\perp}\rangle$ are the unit vectors in the plane such that $\langle\psi_{k}^{\perp}|\psi_{k}\rangle=0;k=1,2,3$ [12].
+Then the optimal observable is $\mathcal{M}=\left\{M_{k};k=1,2,3\right\},$ where $M_{k}=\frac{2}{3}|\psi_{k}^{\perp}\rangle\langle\psi_{k}^{\perp}|,$ and $|\psi_{k}^{\perp}\rangle$ are the unit vectors in the plane such that $\langle\psi_{k}^{\perp}|\psi_{k}\rangle=0;k=1,2,3$ [14].
 In other words, $\mathcal{M}$ is “unambiguous discrimination” measurement (MUD) for the states $\rho_{j}=|\psi_{j}\rangle\langle\psi_{j}|,\,j=1,2,3.$ The analysis based on our optimality criterion shows that the conditions (ii) of the theorem are fulfilled with $\Lambda_{0}=I,$ with the condition (ii.a) equivalent to the entropy inequality
-$-\sum_{j=1}^{3}t_{j}\log t_{j}\geq 1;\quad t_{j}=\frac{2}{3}\cos^{2}\left(\alpha+\frac{2\pi j}{3}\right),$ (60)
+$\sum_{j=1}^{3}t_{j}\log t_{j}\geq 1;\quad t_{j}=\frac{2}{3}\cos^{2}\left(\alpha+\frac{2\pi j}{3}\right),$ (60)
 which is saturated for $\alpha=\pi/2+\pi k$.
-Most interesting is the case of “nearly flat” trine when $06140<r_{0}<\bar{r}_{0}\approx 0,$ 0614 .
-Then, as shown by P.
+Most interesting is the case of “nearly flat” trine when $0<r_{0}<\bar{r}_{0}\approx 0,0614.$ Then, as shown by P.
 Shor in [15], the conjectured optimal observable has 6 outcomes and is a statistical mixture of the MUD described above and the SRM $M_{k}=|{e}_{k}\rangle\langle{e}_{k}|,\,k=1,2,3.$ The SRM becomes optimal for obtuse pyramids with $\bar{r}_{0}\leq r_{0}\leq r_{1}=1/3.$
@@
 Let us apply our optimality criterion to derive the corresponding entropy inequalities and to obtain additional insight.
-We consider the ensemble =$\mathcal{E}=\left\{\pi_{j},\rho_{j}\right\}_{j=1,\dots,m}$ with $\pi_{j}=\frac{1}{m},\,\rho_{j}=|\psi_{j}\rangle\langle\psi_{j}|,$ where $|\psi_{j}\rangle$ are given by (33) with $r_{0}<r_{1}$ $(b<0)$ The conjectured optimal observable $\mathcal{M}$ [2] has $m+\frac{m(m-1)}{2}=\frac{m(m+1)}{2}$ components of the form:
+We consider the ensemble $\mathcal{E}=\left\{\pi_{j},\rho_{j}\right\}_{j=1,\dots,m}$ with $\pi_{j}=\frac{1}{m},\,\rho_{j}=|\psi_{j}\rangle\langle\psi_{j}|,$ where $|\psi_{j}\rangle$ are given by (33) with $r_{0}<r_{1}$ $(b<0)$ .
+The conjectured optimal observable $\mathcal{M}$ [2] has $m+\frac{m(m-1)}{2}=\frac{m(m+1)}{2}$ components of the form:
 $M_{k}$ $=$ $t^{-2}|\tilde{e}_{k}\rangle\langle\tilde{e}_{k}|;\quad k=1,\dots,m;$ (61)
@@
 Again, two cases are possible: $t=1$ and $t>1.$ In the first case the third equation disappears while the first two lead to the same formulas (52) as in the acute case, however with $a=\sqrt{p\text{ }}$, but $b=-\sqrt{\frac{1-p}{m-1}}.$ Then, denoting $z_{j}=\langle e_{j}|\psi\rangle$, the condition (ii.a) brings us to the inequality:
-For complex $z_{j}$ , $\sum_{j=1}^{m}\left|z_{j}\right|^{2}=1,\,$
-$-\sum_{j=1}^{m}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}\geq\tilde{\mu}_{0}(p)\left|\sum_{j=1}^{m}z_{j}\right|^{2}-\tilde{\mu}_{1}(p),$ (66)
+For complex $z_{j}$ satisfying
+$\mathit{\sum_{j=1}^{m}\left|z_{j}\right|^{2}=1,}$ (66)
+it holds
+$\sum_{j=1}^{m}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}\geq\tilde{\mu}_{0}(p)\left|\sum_{j=1}^{m}z_{j}\right|^{2}-\tilde{\mu}_{1}(p),$ (67)
 where
@@
 Note that here $\tilde{\mu}_{0}(p)\leq 0,$ which prevents us from passing to the probabilities $t_{j}=\left|z_{j}\right|^{2}$.
-In (66) the parameter $p$ varies in the range $[\max\left\{\frac{m-1}{m},p(m)\right\},1],$ where $p(m)$ is found below.
+However, as the following lemma shows, we can always restrict to the case of real $z_{j}$.
+Lemma 1.
+Let $\tilde{\mu}_{0},\tilde{\mu}_{1}$ be real numbers. If the inequality
+$\sum_{j=1}^{m}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}+\tilde{\mu}_{0}\left|\sum_{j=1}^{m}z_{j}\right|^{2}\leq\tilde{\mu}_{1}$ (70)
+holds for all real $z_{j}$ satisfying the condition (66), then it holds for all complex $z_{j}$ satisfying the same condition.
+If it holds for all real $z_{j}$ satisfying the condition (66) and
+$\mathit{\sum_{j=1}^{m}z_{j}=0,}$ (71)
+then it holds for all complex $z_{j}$ satisfying the conditions (66), (71).
+Proof.
+Let $z_{j}=x_{j}+iy_{j}$ for real $x_{j},y_{j}$, then (66) amounts to
+$\sum_{j=1}^{m}x_{j}^{2}+\sum_{j=1}^{m}y_{j}^{2}=1,\quad$ (72)
+and (71) amounts to
+$\sum_{j=1}^{m}x_{j}=0,\,\sum_{j=1}^{m}y_{j}=0.$ (73)
+Denote $P=\sum_{j=1}^{m}x_{j}^{2},\,1-P=\sum_{j=1}^{m}y_{j}^{2}$. The function $f(\xi)=\xi\log\xi$ is convex, hence
+$\sum_{j=1}^{m}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}+\tilde{\mu}_{0}\left|\sum_{j=1}^{m}z_{j}\right|^{2}$ $=$ $\sum_{j=1}^{m}f\left(P\left(\frac{x_{j}}{\sqrt{P}}\right)^{2}+(1-P)\left(\frac{y_{j}}{\sqrt{1-P}}\right)^{2}\right)$
+$+$ $P\tilde{\mu}_{0}\left|\sum_{j=1}^{m}\frac{x_{j}}{\sqrt{P}}\right|^{2}+(1-P)\tilde{\mu}_{0}\left|\sum_{j=1}^{m}\frac{y_{j}}{\sqrt{1-P}}\right|^{2}$
+$\leq$ $P\left[\sum_{j=1}^{m}f\left(\left(\frac{x_{j}}{\sqrt{P}}\right)^{2}\right)+\tilde{\mu}_{0}\left|\sum_{j=1}^{m}\frac{x_{j}}{\sqrt{P}}\right|^{2}\right]$
+$+$ $(1-P)\left[\sum_{j=1}^{m}f\left(\left(\frac{y_{j}}{\sqrt{1-P}}\right)^{2}\right)+\tilde{\mu}_{0}\left|\sum_{j=1}^{m}\frac{y_{j}}{\sqrt{1-P}}\right|^{2}\right]$
+$\leq$ $\tilde{\mu}_{1},$
+because
+$\sum_{j=1}^{m}\left(\frac{x_{j}}{\sqrt{P}}\right)^{2}=\sum_{j=1}^{m}\left(\frac{y_{j}}{\sqrt{1-P}}\right)^{2}=1,\quad$
+and by assumption (70) holds for real $z_{j}$ satisfying (66). If additionally the complex $z_{j}$ satisfy (71) then
+$\sum_{j=1}^{m}\frac{x_{j}}{\sqrt{P}}=\sum_{j=1}^{m}\frac{y_{j}}{\sqrt{1-P}}=0,$
+and the second statement of the lemma follows. ∎
+In (67) the parameter $p$ varies in the range $\max\left\{\frac{m-1}{m},p(m)\right\},1],$ where $p(m)$ is found below.
 The accessible information in this case is given by the same expression (45) as in the case of moderately acute pyramids.
+For future use we note that
+$\tilde{\mu}_{1}\left(\frac{m-1}{m}\right)=\log m-\frac{m-2}{m}\log(m-1).$ (74)
 If $t>1,$ then (65) imply $c_{1}=1$ and equations (64) reduce to
@@
 and, via (37) and (56), to
-$p\equiv a^{2}<\left[1+\left(m-1\right)\left(1-\frac{m}{1-\tau_{o}(m)}\right)^{-2}\right]^{-1}\equiv p(m).$ (71)
+$p\equiv a^{2}$ $<$ $\left[1+\left(m-1\right)\left(1-\frac{m}{1-\tau_{o}(m)}\right)^{-2}\right]^{-1}$ (77)
+$=$ $\frac{(m-1+\tau_{o}(m))^{2}}{m(m-1+\tau_{o}(m)^{2})}\equiv p(m).$
+One can check that equation (76) with $\tau=\tau_{o}(m)$ is equivalent to
+$\tilde{\mu}_{1}(p(m))=1.$ (78)
 The numerically found values [^5] By using equation (76) it is possible to express the function $p(m)$ via the Lambert $W$-function.:
@@
 The condition (ii.a) then amounts to the entropy inequality (67) with $\tilde{\mu}_{0}=c_{0}/m,\quad\tilde{\mu}_{1}=-1$ i.e.
-$-\sum_{j=1}^{m}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}\geq\frac{\left(\tau_{o}(m)^{-1}-1\right)}{m}\log\frac{2\left(\tau_{o}(m)-1\right)^{2}}{m\left(m-1+\tau_{o}(m)^{2}\right)}\left|\sum_{j=1}^{m}z_{j}\right|^{2}+1.$ (72)
-The equality is attained for $z_{1}=-z_{2}=1/\sqrt{2},z_{j}=0$, $j\geq 3$; for $z_{1}=\sqrt{p(m)},\,z_{j}=-\sqrt{\frac{1-p(m)}{m-1}},\,j\geq 2$; and for all permutations of these $z_{j}.$
+$\sum_{j=1}^{m}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}\geq\frac{\left(\tau_{o}(m)^{-1}-1\right)}{m}\log\frac{2\left(\tau_{o}(m)-1\right)^{2}}{m\left(m-1+\tau_{o}(m)^{2}\right)}\left|\sum_{j=1}^{m}z_{j}\right|^{2}+1.$ (79)
+Due to lemma 1 one can restrict here to real $z_{j}$ , $\sum_{j=1}^{m}\left|z_{j}\right|^{2}=1.$ The equality is attained for $z_{1}=-z_{2}=1/\sqrt{2},z_{j}=0$, $j\geq 3$; for $z_{1}=\sqrt{p(m)},\,z_{j}=-\sqrt{\frac{1-p(m)}{m-1}},\,j\geq 2$; and for all permutations of these $z_{j}.$
 In this case the accessible information, by using (58), is
@@
 If $m\leq 6$ then $p(m)>\frac{m-1}{m}$ and the inequality (67) holds for $p\in\left[p(m),1\right]$, yielding the inequality (79) for $p\in\left[\frac{m-1}{m},p(m)\right]$.
-The proof of the inequalities is given in section 5.2.
+The suggestion for a proof of the inequalities is given in section 5.2.
 Corollary 2.
@@
 Indeed, the unit vector $|\psi\rangle$ lies in the hyperplane $\langle e_{0}|\psi\rangle=0$, hence the second equality follows.
-Arguing as in the previous section, we suggest that the hypothetical optimal observable is given by (61) with $t=1$ for $m\geq 7$ and $t=\infty$ for $m\leq 6$.
-The corresponding tight entropy inequality is
-$-\sum_{j=1}^{m}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}\geq\left\{\begin{array}[]{cc}1,&m\leq 6;\\ \log m-\frac{m-2}{m}\log(m-1),&m\geq 7,\end{array}\right.$ (75)
-for complex $z_{j}$ satisfying (74).
-Indeed, for $m\leq 6$ we can use the argument leading to (72) with $\sum_{j=1}^{m}z_{j}=0.$ On the other hand, for $m\geq 7$ we should use (66) with $\sum_{j=1}^{m}z_{j}=0$ and $p=\frac{m-1}{m}.$ Then the first term in the righthand side vanish, while $\tilde{\mu}_{1}(p)$ gets the value
-$-\frac{m-1}{m}\log\frac{m-1}{m}-\frac{1}{m}\log\frac{1}{m(m-1)}=\log m-\frac{m-2}{m}\log(m-1),$ (76)
-which is greater than 1 for $m\leq 6$, but less for $m\geq 7$.
+Arguing as in the previous section, we suggest that the optimal observable is given by (61) with $t=1$ for $m\geq 7$ and $t=\infty$ for $m\leq 6$. The corresponding tight entropy inequality is
+$\sum_{j=1}^{m}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}\geq\left\{\begin{array}[]{cc}1,&m\leq 6;\\ \log m-\frac{m-2}{m}\log(m-1),&m\geq 7,\end{array}\right$. (82)
+for complex $z_{j}$ satisfying (81)  (due to the second statement of lemma 1 one can restrict here to real $z_{j}$ ).
+Indeed, for $m\leq 6$ we can use the argument leading to inequality (79) with $\sum_{j=1}^{m}z_{j}=0$ and $p=p(m)$.
+Then (78) implies the first line in the right-hand side of (82).
+On the other hand, for $m\geq 7$ we should use the inequality (67) with $\sum_{j=1}^{m}z_{j}=0$ and $p=\frac{m-1}{m}.$ Then the first term in the righthand side vanish, while $\tilde{\mu}_{1}(p)\$gets the value (74) i.e.
+the second line in the righthand side of (82), which is greater than 1 for $m\leq 6$, but less for $m\geq 7$.
+We thus arrive at the following special case of theorem 3:
 Proposition 2.
@@
 In the case $m\leq 6$ the equality in (82) is attained for $z_{1}=-z_{2}=1/\sqrt{2},\,z_{j}=0$ for $j\geq 3;$ in the case $m\geq 7$ – for $z_{1}=\sqrt{\frac{m-1}{m}},\,z_{j}=-\sqrt{\frac{1}{\left(m-1\right)m}},\,j\geq 2,$ (and for all permutations of such $z_{j}).$
-Proof.
-By lemma 1 below we can restrict to the case of real $z_{j}$. Introducing the Lagrange function
-$L\left(z_{1},\dots,z_{m}\right)=-\sum_{j=1}^{m}z_{j}^{2}\log z_{j}^{2}-\lambda\sum_{j=1}^{m}z_{j}^{2}-\mu\sum_{j=1}^{m}z_{j},$
-which is continuously differentiable, we have the necessary condition for a minimizer of the lefthand side of (75)
-$-\frac{\partial L}{\partial z_{j}}=2z_{j}\log ez_{j}^{2}+2\lambda z_{j}+\mu=0,\quad j=1,\dots,m.$
-Case 1.
-The minimizer has a zero component $z_{j}$.
-Then $\mu=0$ and the equation $02z\log ez^{2}+2\lambda z=$ 0 has the roots $0,c,-c.$ Therefore the minimizer should have the form $(\underset{k\text{ times}}{\underbrace{c,-c,\dots,c,-c,}}\,0,\dots,0),$ where $12kc^{2}=$ 1 giving the value $-2kc^{2}\log c^{2}=\log 2k$ which has the minimum equal to 1 for $k=1.$
-Case 2.
-The minimizer has no zero components.
-The equation $02z\log ez^{2}+2\lambda z+\mu=$ 0 has at most three real, nonzero roots $c_{1},c_{2},c_{3}$, which enter the minimizer with the multiplicities $k_{1},k_{2},$ $k_{3},$ such that $k_{1}+k_{2}+k_{3}=m.$ Then consider the Lagrange function
-$\tilde{L}\left(c_{1},c_{2},c_{3};k_{1},k_{2},k_{3}\right)=-\sum_{j=1}^{3}k_{j}c_{j}^{2}\log c_{j}^{2}-\lambda\sum_{j=1}^{3}k_{j}c_{j}^{2}-\mu\sum_{j=1}^{3}k_{j}c_{j}-\nu\sum_{j=1}^{3}k_{j}.$
-The necessary condition is
-$-\frac{\partial\tilde{L}}{\partial c_{j}}=2k_{j}c_{j}\log ec_{j}^{2}+2\lambda k_{j}c_{j}+\mu k_{j}=0,$
-$-\frac{\partial\tilde{L}}{\partial k_{j}}=c_{j}^{2}\log ec_{j}^{2}+\lambda c_{j}^{2}+\mu c_{j}+\nu=0,\quad j=1,2,3.$
-Taking into account that $c_{j}\neq 0,k_{j}\geq 1,$ we can multiply the first equation by $c_{j}/\left(2k_{j}\right)$ without lost of information and substract the second equation, resulting in
-$c_{j}^{2}-\left(\mu/2\right)c_{j}-\nu=0,$
-which has at most two different roots, say $c_{1},c_{2}$, entering the minimizer with the multiplicities $k_{1},k_{2},$ such that $k_{1}+k_{2}=m.$ Then $k_{1}c_{1}+k_{2}c_{2}=0,k_{1}c_{1}^{2}+k_{2}c_{2}^{2}=1$ whence $c_{1}=\sqrt{\frac{k_{2}}{k_{1}m}},c_{1}=-\sqrt{\frac{k_{1}}{k_{2}m}}$ and the value of the lefthand side in (75) is
-$-k_{1}c_{1}^{2}\log c_{1}^{2}-k_{2}c_{2}^{2}\log c_{2}^{2}=-\frac{k_{2}}{m}\log\frac{k_{2}}{k_{1}m}-\frac{k_{1}}{m}\log\frac{k_{1}}{k_{2}m}.$
-This is a concave function of $k_{1}=m-k_{2}$ symmetric on the interval $[1,m-1]$ hence it attains the minimum at the edges of the interval, which is equal to (76).
-Since (76) is greater than 1 for $m\leq 6$, but less for $m\geq 7$, we arrive to (75).
-∎
 Theorem 1 then implies
@@
 where $\||\tilde{\psi}_{j}\rangle\|=\||e_{j}\rangle-\sqrt{1/3}|e_{0}\rangle\|=\sqrt{2/3}$, hence (82) is the same as (60).
-Lemma 1.
-If the inequality (75) holds for all real $z_{j}$ satisfying the conditions (74), then it holds for all complex $z_{j}$ satisfying the same conditions.
-Proof.
-Let $z_{j}=x_{j}+iy_{j}$ for real $x_{j},y_{j}$, then (74) amounts to
-$\sum_{j=1}^{m}x_{j}^{2}+\sum_{j=1}^{m}y_{j}^{2}=1,\quad\sum_{j=1}^{m}x_{j}=0,\,\sum_{j=1}^{m}y_{j}=0.$ (77)
-[…]
-$\sum_{j=1}^{m}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}$ $=$ $\sum_{j=1}^{m}f\left(P\frac{x_{j}^{2}}{P}+(1-P)\frac{y_{j}^{2}}{1-P}\right)$
-$\leq$ $P\sum_{j=1}^{m}f\left(\frac{x_{j}^{2}}{P}\right)+(1-P)\sum_{j=1}^{m}f\left(\frac{y_{j}^{2}}{1-P}\right)$
-$=$ $P\sum_{j=1}^{m}\frac{x_{j}^{2}}{P}\log\left(\frac{x_{j}^{2}}{P}\right)$
-[…]
-because
-$\sum_{j=1}^{m}\frac{x_{j}^{2}}{P}=\sum_{j=1}^{m}\frac{y_{j}^{2}}{1-P}=1,\quad\sum_{j=1}^{m}\frac{x_{j}}{\sqrt{P}}=\sum_{j=1}^{m}\frac{y_{j}}{\sqrt{1-P}}=0,$
-and by assumption (75) holds for real $z_{j}$. ∎
-5 Proofs
-5.1 Proof of theorem 2
+5 Suggestions for proofs
+5.1 The case of theorem 2
 Let $m\geq 2,\ \dfrac{m-1}{m}\leq p<1$ and let $\mu_{0}(p),\mu_{1}(p)$ be given by (28), (29).
 For $p\in[(m-1)/m,1)$ and a probability distribution $(t_{1},...,t_{m})$, it holds
-$F(t):=H(t)-\Bigl{(}\mu_{0}(p)\Bigl{(}\sum\limits_{k=1}^{m}\sqrt{t_{k}}\Bigl{)}^{2}-\mu_{1}(p)\Bigl{)}\geq 0,$ (78)
+$F(t)\equiv H(t)-\Bigl{(}\mu_{0}(p)\Bigl{(}\sum\limits_{k=1}^{m}\sqrt{t_{k}}\Bigl{)}^{2}-\mu_{1}(p)\Bigl{)}\geq 0,$ (83)
 where $H(t)=-\sum\limits_{k=1}^{m}t_{k}\log t_{k}$ is the Shannon entropy.
-In the special case $p=1$, the inequality (78) is trivially satisfied by substituting $\mu_{0}(1):=0,\ \mu_{1}(1)=0$.
+In the special case $p=1$, the inequality (83) is trivially satisfied by substituting $\mu_{0}(1)\equiv 0,\ \mu_{1}(1)=0$.
 Proof.
@@
 $u(s)=-\sum\limits_{i=1}^{m}(t_{*,i}+sa_{i})\log(t_{*,i}+sa_{i})-\mu_{0}(p)\Bigl{(}\sum\limits_{i=1}^{m}\sqrt{t_{*,i}+sa_{i}}\Bigl{)}+\mu_{1}(p)$
-for the boundary point $t_{*}$ and the transversal direction $a$ has the derivative $\infty$- ∞ at zero:
+for the boundary point $t_{*}$ and the transversal direction $a$ has the derivative -$\infty$ at zero:
 $u^{\prime}(0)=\lim\limits_{t\to t_{*}}\Bigg{[}-\sum\limits_{i\in I_{0}}a_{i}\log t_{i}-C\Bigl{(}\sum\limits_{i\in I_{0}}\dfrac{a_{i}}{\sqrt{t_{i}}}\Bigl{)}+O(1)\Bigg{]},$
@@
 up to permutations, for some $1\leq k\leq m$ and $q\in[0,1]$ (for $k=m$ it is assumed that $q=1/m$).
-One of the critical points is equal to $(p,(1-p)/(m-1),...,(1-p)/(m-1))$, in which the value of the function is equal to 00.
+One of the critical points is equal to (p,(1p)/(m1),,(1p)/(m1))(p,(1-p)/(m-1),...,(1-p)/(m-1)), in which the value of the function is equal to 0.
 Proof.
@@
 ∎
-Lemma 3.
-Let $A>0,\ B,\ C\in\mathbb{R}$. Then the function
-$f(t)=h(t)-(A\sqrt{t(1-t)}+Bt+C)$ (82)
-[…]
-Proof.
-$f^{\prime}(0)=-\infty,\quad f^{\prime}(1)=+\infty,$
-therefore local minima are on the interval $(0,1)$.
-In the calculations we will move on to the natural logarithm.
-$f^{\prime}(t)=-\ln\dfrac{t}{1-t}-\dfrac{A\sqrt{t(1-t)}}{2}\Bigl{(}\dfrac{1}{t}-\dfrac{1}{1-t}\Bigl{)}-B=$
-$=-\ln\dfrac{t}{1-t}-\dfrac{A(1-2t)}{2\sqrt{t(1-t)}}-B;$
-$f^{\prime\prime}(t)=-\dfrac{1}{t(1-t)}+\dfrac{A}{4(t(1-t))^{3/2}}.$
-$f^{\prime\prime}(t)=0\quad\Leftrightarrow\quad A=4\sqrt{t(1-t)}.$
-Therefore, there are two inflection points located on different sides of $t=1/2$.
-It follows that there are 3 critical points $t_{1}\leq t_{0}\leq t_{2}$, where $t_{1},\ t_{2}$ are local minimum points.
-[…]
+The proof of the following lemma is based on the behavior of the difference between binary entropy and a function containing both a linear component and a square root term $\sqrt{t(1-t)}$.
+We conduct a qualitative analysis of the minimizer of this function.
+The global minimum (there can be no more than two local minima) appears on either the interval $[0,1/2]$ or $[1/2,1]$.
+Importantly, this segment contains no other critical points.
+Therefore, minimizing the function reduces to identifying the critical point within the appropriate interval and evaluating the function at that point.
+A detailed proof of this statement will be provided at the end of the section.
 The variables $\alpha,\ p,\beta,\ t$ used in what follows, unless otherwise stated, are such that
-$\alpha\in{(0,1/2],\quad p\in[1-\alpha,1},\quad\beta\in[\alpha,1/2],\quad t\in[0,1].$
+$\alpha\in{(0,1/2],\quad p\in[1-\alpha,1)},\quad\beta\in[\alpha,1/2],\quad t\in[0,1].$
 Lemma 3.
 For introduced $\alpha,\ p,\ \beta,\ t$ the inequality
-$\mu_{0}(p,\alpha)\Bigl{(}\sqrt{\beta t}+\sqrt{(1-\beta)(1-t)}\Bigl{)}^{2}-\mu_{1}(p,\alpha)$ (83)
+$\mu_{0}(p,\alpha)\Bigl{(}\sqrt{\beta t}+\sqrt{(1-\beta)(1-t)}\Bigl{)}^{2}-\mu_{1}(p,\alpha)$
 $\leq$ $-t\log\dfrac{t}{\beta}-(1-t)\log\dfrac{1-t}{1-\beta}$
@@
 $f(t)=-t\log\dfrac{t}{\beta}-(1-t)\log\dfrac{1-t}{1-\beta}-\Bigl{(}\mu_{0}(p,\alpha)\Bigl{(}\sqrt{\beta t}+\sqrt{(1-\beta)(1-t)}\Bigl{)}^{2}-\mu_{1}(p,\alpha)\Bigl{)},$
-for which $A=2\mu_{0}(p,\alpha)\sqrt{\beta(1-\beta)}$, $B=\log\dfrac{1-\beta}{\beta}+\mu_{0}(p,\alpha)(2\beta-1)$ in the form (82).
+for which $A=2\mu_{0}(p,\alpha)\sqrt{\beta(1-\beta)}$, $B=\log\dfrac{1-\beta}{\beta}+\mu_{0}(p,\alpha)(2\beta-1)$ in the form (89) (see lemma 4 below).
 The derivative of $f(t)$ has form
@@
     $B\leq 0$.
-    In this case there is minimum point $t_{*}=\beta$ ($f^{\prime}(\beta)=0$) of the function $f(t)$ on $[0,1/2]$ due to the Lemma 3. One has
+    In this case there is minimum point $t_{*}=\beta$ ($f^{\prime}(\beta)=0$) of the function $f(t)$ on $[0,1/2]$ due to the lemma 4 presented later. One has
     $f(\beta)=\mu_{1}(p,\alpha)-\mu_{0}(p,\alpha)\geq 0.$ (87)
@@
     $\sqrt{\alpha p}\log\dfrac{p}{\alpha}+\sqrt{(1-\alpha)(1-p)}\log\dfrac{1-p}{1-\alpha}\geq 0.\quad$ (88)
-    Let $L(p)=\sqrt{\alpha p}\ln\dfrac{p}{\alpha}+\sqrt{(1-\alpha)(1-p)}\ln\dfrac{1-p}{1-\alpha}$, then
+    Let $V(p)=\sqrt{\alpha p}\ln\dfrac{p}{\alpha}+\sqrt{(1-\alpha)(1-p)}\ln\dfrac{1-p}{1-\alpha}$, then
     1. (a)
-      $(d/dp)^{2}L(p)=-\dfrac{\sqrt{\alpha}\ln(p/\alpha)}{4p^{3/2}}-\dfrac{\sqrt{1-\alpha}\ln((1-p)/(1-\alpha))}{4(1-p)^{3/2}}\geq 0$
+      $(d/dp)^{2}V(p)=-\dfrac{\sqrt{\alpha}\ln(p/\alpha)}{4p^{3/2}}-\dfrac{\sqrt{1-\alpha}\ln((1-p)/(1-\alpha))}{4(1-p)^{3/2}}\geq 0$
       (since $p\geq 1-\alpha\geq 1/2$);
     2. (b)
-      $L(1-\alpha)=0$ and
-      $L^{\prime}(1-\alpha)=\sqrt{\dfrac{\alpha}{1-\alpha}}-1+\dfrac{1}{2}\ln\dfrac{\alpha}{1-\alpha}+\dfrac{1}{2}\sqrt{\dfrac{\alpha}{1-\alpha}}\ln\dfrac{1-\alpha}{\alpha}\geq 0\quad\Leftrightarrow$
+      $V(1-\alpha)=0$ and
+      $V^{\prime}(1-\alpha)=\sqrt{\dfrac{\alpha}{1-\alpha}}-1+\dfrac{1}{2}\ln\dfrac{\alpha}{1-\alpha}+\dfrac{1}{2}\sqrt{\dfrac{\alpha}{1-\alpha}}\ln\dfrac{1-\alpha}{\alpha}\geq 0\quad\Leftrightarrow$
       $\dfrac{\alpha}{1-\alpha}\Bigl{(}1+\dfrac{1}{2}\ln\dfrac{1-\alpha}{\alpha}\Bigl{)}-\Bigl{(}1-\dfrac{1}{2}\ln\dfrac{1-\alpha}{\alpha}\Bigl{)}\geq 0\quad\Leftrightarrow$
       $\ln\dfrac{1-\alpha}{\alpha}\geq 2-4\alpha.$
-    Therefore, $L(p)\geq 0$ and the inequality (85) is satisfied.
+    Therefore, $V(p)\geq 0$ and the inequality (88) is satisfied.
   2. 2.
     $B\geq 0$.
-    This condition implies that the function $f(t)$ has the minimum point $t_{*}$ on $[1/2,1]$. This point is characterized by the equation
+    This condition implies that the function $f(t)$ has the minimum point $t_{*}$ on $[1/2,1]$ (lemma 4 is applied in the same way). This point is characterized by the equation
     $\log\dfrac{1-t_{*}}{t_{*}}+\log\dfrac{\beta}{1-\beta}-\mu_{0}(p,\alpha)\Bigg{(}\sqrt{\dfrac{\beta}{t_{*}}}-\sqrt{\dfrac{1-\beta}{1-t_{*}}}\Bigg{)}\Bigl{(}\sqrt{\beta t_{*}}+\sqrt{(1-\beta)(1-t_{*})}\Bigl{)}=0.$
-    The value of $t_{*}$ depends on $\beta$, and
+    The value of $t_{*}$ depends on $\beta$, and for the function $f(\beta,t)\equiv f(t)$ it holds $\dfrac{d}{d\beta}f(\beta,t_{*}(\beta))=\dfrac{\partial}{\partial\beta}f(\beta,t_{*}(\beta))+\dfrac{dt_{*}(\beta)}{d\beta}\dfrac{\partial}{\partial t}f(\beta,t_{*}(\beta))=\dfrac{\partial}{\partial\beta}f(\beta,t_{*}(\beta))$ since $t_{*}$ is critical for $f(t)$.
+Then,
     $\dfrac{d}{d\beta}f(t_{*})=\dfrac{1}{\ln 2}\Bigl{(}\dfrac{t_{*}}{\beta}-\dfrac{1-t_{*}}{1-\beta}\Bigl{)}-\mu_{0}(p,\alpha)\Bigg{(}\sqrt{\dfrac{t_{*}}{\beta}}-\sqrt{\dfrac{1-t_{*}}{1-\beta}}\Bigg{)}\times$
-    ×$\times\Bigl{(}\sqrt{\beta t_{*}}+\sqrt{(1-\beta)(1-t_{*})}\Bigl{)}.$
+    $\times\Bigl{(}\sqrt{\beta t_{*}}+\sqrt{(1-\beta)(1-t_{*})}\Bigl{)}.$
     Then
@@
     since
-    $\sqrt{\dfrac{1-\beta}{\beta}\dfrac{t_{*}}{1-t_{*}}}-\sqrt{\dfrac{\beta}{1-\beta}\dfrac{1-t_{*}}{t_{*}}}+\ln\Bigl{(}\dfrac{\beta}{1-\beta}\dfrac{1-t_{*}}{t_{*}}\Bigl{)}\equiv C-1/C-2\ln C\geq 0$
-    for $C\geq 1$.
+    $\sqrt{\dfrac{1-\beta}{\beta}\dfrac{t_{*}}{1-t_{*}}}-\sqrt{\dfrac{\beta}{1-\beta}\dfrac{1-t_{*}}{t_{*}}}+\ln\Bigl{(}\dfrac{\beta}{1-\beta}\dfrac{1-t_{*}}{t_{*}}\Bigl{)}=R-1/R-2\ln R\geq 0$
+    where $R\equiv\sqrt{\dfrac{1-\beta}{\beta}\dfrac{t_{*}}{1-t_{*}}}\geq 1$.
     It is simple to calculate that for $\beta=\alpha$, we have $t_{*}=p$, and $f(p)=0$.
@@
 which takes place for $p\in(0,1)$. The function $h_{1/2}(p)$ is nonnegative, convex and $h_{1/2}(p)\leq\sqrt{2}$.
-5.2 Proof of theorem 3
-This section is devoted to proving the inequality (66) which verificates the condition (ii.a) in case of obtuse pyramid.
-Namely, if non-negative parameter $p<1$ is larger than the maximum of $\dfrac{m-1}{m}$ and $p(m)$ that is defined in (71), then for complex numbers $z_{1},...,z_{m}$ forming probability distribution $(|z_{1}|^{2},...,|z_{m}|^{2})$, the inequality
-$-\sum\limits_{j=1}^{m}|z_{j}|^{2}\log|z_{j}|^{2}\geq\tilde{\mu}_{0}(p)\Bigl{|}\sum\limits_{j=1}^{m}z_{j}\Bigl{|}^{2}-\tilde{\mu}_{1}(p).$
+The proof of lemma 3 is based on a simple fact of analysis:
+Lemma 4.
+Let $A>0,\ B,\ C\in\mathbb{R}$ (The notation of the coefficients is used only within the framework of the formulation of the lemma and its proof). Then the function
+$g(t)=h(t)-(A\sqrt{t(1-t)}+Bt+C)$ (89)
+[…]
+Proof.
+$g^{\prime}(0)=-\infty,\quad g^{\prime}(1)=+\infty,$
+therefore local minima are on the interval $(0,1)$.
+In the calculations we will move on to the natural logarithm.
+$g^{\prime}(t)=-\ln\dfrac{t}{1-t}-\dfrac{A\sqrt{t(1-t)}}{2}\Bigl{(}\dfrac{1}{t}-\dfrac{1}{1-t}\Bigl{)}-B=$
+$=-\ln\dfrac{t}{1-t}-\dfrac{A(1-2t)}{2\sqrt{t(1-t)}}-B;$
+$g^{\prime\prime}(t)=-\dfrac{1}{t(1-t)}+\dfrac{A}{4(t(1-t))^{3/2}}.$
+$g^{\prime\prime}(t)=0\quad\Leftrightarrow\quad A=4\sqrt{t(1-t)}.$
+Therefore, there are two inflection points located on different sides of $t=1/2$.
+It follows that there are 3 critical points $t_{1}\leq t_{0}\leq t_{2}$, where $t_{1},\ t_{2}$ are local minimum points.
+The last statement obviously follows from the symmetric case.
+Indeed, for $B<0$, it holds $g(t)\leq g(1-t)$ for $t\in[1/2,1]$, therefore the global minimum lies on the segment $[1/2,1]$.
+Let s, 1ss,\ 1-s ($0<s<1/2$) be zeros of $g^{\prime\prime}(t)$.
+On the segment [s,1s][s,1-s] the function $g(t)^{\prime}$ is decreasing, on the segments [0,s],[1s,1][0,s],\ [1-s,1] the function $g(t)^{\prime}$ increases, and due to symmetry $(g(t)+Bt)^{\prime}$ has root $1/2$.
+Therefore, on the segment $[1/2,1]$ when $B<0$ the function $g(t)$ has no more than one critical point.
+∎
+5.2 The case of theorem 3
+This section is devoted to proving the inequality (67) which verifies the condition (ii.a) in the case of obtuse pyramid.
+Namely, if the parameter $p<1$ satisfies $p\geq\max\left\{\dfrac{m-1}{m},p(m)\right\}$ with $p(m)$ defined in (77), then for complex numbers $z_{1},...,z_{m}$ satisfying (66) the inequality
+$\tilde{F}(z)\equiv-\sum\limits_{j=1}^{m}|z_{j}|^{2}\log|z_{j}|^{2}-\tilde{\mu}_{0}(p)\Bigl{|}\sum\limits_{j=1}^{m}z_{j}\Bigl{|}^{2}+\tilde{\mu}_{1}(p)\geq 0.$ (90)
 holds with $\tilde{\mu}_{0}(p),\tilde{\mu}_{1}(p)$ defined in (68), (69).
-For $p=1$, which corresponds to the orthonormal case, the inequality reduces to a trivial one: $H(|z_{1}|^{2},...,|z_{m}|^{2})\geq 0$.
-Proof.
-Denote $z_{j}=x_{j}+iy_{j}$ and consider the Lagrange function
-$L(x_{1},...,x_{m},y_{1},...,y_{m})=-\sum\limits_{k=1}^{m}(x_{k}^{2}+y_{k}^{2})\log(x_{k}^{2}+y_{k}^{2})-\tilde{\mu}_{0}(p)\Bigl{(}\sum\limits_{k=1}^{m}x_{k}\Bigl{)}^{2}-$ (86)
-$-\tilde{\mu}_{0}(p)\Bigl{(}\sum\limits_{k=1}^{m}y_{k}\Bigl{)}^{2}+\tilde{\mu}_{1}(p)-\lambda\Bigl{(}\sum\limits_{k=1}^{m}(x_{k}^{2}+y_{k}^{2})-1\Bigl{)}.$ (87)
-The necessary conditions are:
-$\begin{cases}-\dfrac{1}{2}\dfrac{\partial L}{\partial x_{k}}=x_{k}\Bigl{(}\log(x_{k}^{2}+y_{k}^{2})+\log e\Bigl{)}+\tilde{\mu}_{0}(p)\sum\limits_{k=1}^{m}x_{j}+\lambda x_{k},\\ -\dfrac{1}{2}\dfrac{\partial L}{\partial y_{k}}=y_{k}\Bigl{(}\log(x_{k}^{2}+y_{k}^{2})+\log e\Bigl{)}+\tilde{\mu}_{0}(p)\sum\limits_{k=1}^{m}y_{j}+\lambda y_{k}.\end{cases}$ (88)
-for all $1\leq k\leq m$.
-Hence, we get
-$\dfrac{1}{2}\Bigl{(}y_{k}\dfrac{\partial L}{\partial x_{k}}-x_{k}\dfrac{\partial L}{\partial y_{k}}\Bigl{)}=\tilde{\mu}_{0}(p)\Bigl{(}x_{k}\sum\limits_{j=1}^{m}y_{j}-y_{k}\sum\limits_{j=1}^{m}x_{j}\Bigl{)}=0.$ (89)
-Case 1. $\sum\limits_{j=1}^{m}z_{j}=0$ or $\tilde{\mu}_{0}(p)=0$.
-We get the inequality $-\sum\limits_{j=1}^{m}|z_{j}|^{2}\log|z_{j}|^{2}\geq-\tilde{\mu}_{1}(p)$. Taking into account proposition 2 for the flat pyramid, it is sufficient to prove the lemma
-Lemma 5.
-Let $m\geq 2$ be a natural number.
-  1. 1.
-    If $\dfrac{m-1}{m}\leq p<1$, then $-\tilde{\mu}_{1}(p)\leq-\dfrac{m-2}{m}\log(m-1)+\log m$;
-  2. 2.
-    if $m\leq 6$ and $p(m)\leq p<1$, then $-\tilde{\mu}_{1}(p)\leq 1$.
-Case 2.
-$\sum\limits_{j=1}^{m}z_{j}\neq 0$.
-Then the condition (89) reduces to proportionality of two vectors $(x_{1},...,x_{m})$ and $(y_{1},...,y_{m})$.
-Due to the symmetry of (86) under multiplication by a scalar $e^{i\varphi},\ \varphi\in\mathbb{R}$, the problem reduces to the case of real variables $\{z_{j}\}$.
-For the real variables $\{z_{j}\}$ let us consider the new Lagrange function
-$L_{(1)}(z_{1},...,z_{k})=-\sum\limits_{k=1}^{m}(z_{k}^{2})\log(z_{k}^{2})-\tilde{\mu}_{0}(p)\Bigl{(}\sum\limits_{k=1}^{m}z_{k}\Bigl{)}^{2}+\tilde{\mu}_{1}(p)-\lambda\Bigl{(}\sum\limits_{k=1}^{m}z_{k}^{2}-1\Bigl{)}.$
-Thus, the necessary condition is
-$-\dfrac{1}{2}\dfrac{\partial L_{(1)}}{\partial z_{k}}=z_{k}(\log z_{k}^{2}+\log e)+\tilde{\mu}_{0}(p)\Bigl{(}\sum\limits_{j=1}^{m}z_{j}\Bigl{)}+\lambda z_{k}=0,\quad 1\leq k\leq m,$
-and the set $\{z_{j}\}$ has at most $3$ different values $c_{1},c_{2},c_{3}$ with multiplicities $k_{1},k_{k},k_{3}$, since the equation
-$z(\log z^{2}+\log e)+\tilde{\mu}_{0}(p)\Bigl{(}\sum\limits_{j=1}^{m}z_{j}\Bigl{)}+\lambda z=0$
-for a variable $z$ has at most $3$ real solutions.
-The fact the the equation of the form $z\log z^{2}+az+b=0$ for real $a$ and $b$ has no more then $3$ real roots was actually used in proof of proposition 2.
-To reduce the number of different values to two, we will introduce a new Lagrange function $L_{(2)}$.
-The critical points of $L_{(1)}$ are related with critical points of the Lagrange function
-$L_{(2)}(c_{1},c_{2},c_{3},k_{1},k_{2},k_{3})=-\sum\limits_{j=1}^{3}k_{j}c_{j}^{2}\log c_{j}^{2}-\tilde{\mu}_{0}(p)\Bigl{(}\sum\limits_{j=1}^{m}k_{j}c_{j}\Bigl{)}^{2}-$
-$-\lambda\Bigl{(}\sum\limits_{j=1}^{3}k_{j}c_{j}^{2}-1\Bigl{)}-\nu\Bigl{(}\sum\limits_{j=1}^{3}k_{j}-m\Bigl{)}.$
+Suggestion for proof.
+As it has been established in lemma 1, a real case is enough for analysis.
+For real vectors $(z_{j})_{j=1}^{m}$ from the unit sphere let us consider the Lagrange function
+$L(z_{1},...,z_{k})=-\sum\limits_{k=1}^{m}z_{k}^{2}\log z_{k}^{2}-\tilde{\mu}_{0}(p)\Bigl{(}\sum\limits_{k=1}^{m}z_{k}\Bigl{)}^{2}+\tilde{\mu}_{1}(p)-\lambda\Bigl{(}\sum\limits_{k=1}^{m}z_{k}^{2}-1\Bigl{)}.$
+The necessary condition is
+$\dfrac{1}{2}\dfrac{\partial L}{\partial z_{k}}=z_{k}\log z_{k}+\tilde{\mu}_{0}(p)\Bigl{(}\sum\limits_{j=1}^{m}z_{j}\Bigl{)}+(\lambda+\log e)z_{k}=0,\quad 1\leq k\leq m,$
+and the set $\{z_{j}\}$ has at most three different values $s_{1},s_{2},s_{3}$ with multiplicities $k_{1},k_{2},k_{3}\geq 0$ ($k_{1}+k_{2}+k_{3}=m$), since the equation
+$z\log z^{2}+\tilde{\mu}_{0}(p)\Bigl{(}\sum\limits_{j=1}^{m}z_{j}\Bigl{)}+(\lambda+\log e)z=0$ (91)
+for a variable $z$ has at most three real solutions.
+The fact the equation of the form $z\log z^{2}+az+b=0$ for real $a$ and $b$ has no more then three real roots was actually used in proof of lemma 2 for equation (86) which is similar to (91) with $\leq 2$ roots.
+However, here we consider the entire real line, so the maximum number of real roots is three.
+Consider the following cases.
+Case 1. The minimizer has a zero component.
+Then $\lambda+\log e=0$ and the equation (91) has roots 0 and $\pm s$.
+Therefore the minimizer of the function $\tilde{F}(z)$ should have the form $(\underbrace{s,-s,...,s,-s}_{k\text{ times}},0,...,0)$, where $2ks^{2}=1$ giving the value $2ks^{2}\log s^{2}+\tilde{\mu}_{1}(p)=\log 2k+\tilde{\mu}_{1}(p)$ which has the minimum equal to $1+\tilde{\mu}_{1}(p)$ for $k=1$.
+The value $1+\tilde{\mu}_{1}(p)$ is nonnegative for the following reason.
+First we will prove that the function $\tilde{\mu}_{1}(p)$ is increasing for $p$ in the range $\Bigl{[}\dfrac{m-1}{m},1\Bigr{)}$. The derivative is
+$\tilde{\mu}_{1}^{\prime}(p)=\log e\dfrac{2(m-2)\sqrt{\dfrac{p(1-p)}{m-1}}-4p+2+\ln p-\ln\dfrac{1-p}{m-1}}{2(m-1)\sqrt{\dfrac{p(1-p)}{m-1}}\Bigg{(}\sqrt{p}+\sqrt{\dfrac{1-p}{m-1}}\Bigg{)}^{2}}.$
+Here $2(m-2)\sqrt{\dfrac{p(1-p)}{m-1}}\geq 0$ and $\ln\dfrac{p}{1-p}-4p+2\geq 0$, therefore, $\tilde{\mu}_{1}^{\prime}(p)\geq 0$.
+Thus for $\frac{m-1}{m}\leq p<1$ we have
+$\tilde{\mu}_{1}(p)\leq-\tilde{\mu}_{1}\Bigl{(}\frac{m-1}{m}\Bigl{)}=-\dfrac{m-2}{m}\log(m-1)+\log m$
+due to (74). For $p(m)\leq p<1$ we have
+$\tilde{\mu}_{1}(p)\leq-\tilde{\mu}_{1}(p(m))=1,$
+because of (78). Therefore, $1+\tilde{\mu}_{1}(p)\geq 0$ for $p>\max(p(m),(m-1)/m)$. Note that $\tilde{\mu}_{1}(p)\geq\tilde{\mu}_{1}((m-1)/m)\geq\tilde{\mu}_{1}(p(m))=-1$ if p(m)<(m1)/mp(m)<(m-1)/m.
+Case 2. The minimizer has no zero components.
+To reduce the number of different values to two, we will introduce a new Lagrange function $\hat{L}$. The critical points of $L$ are related with critical points of the Lagrange function
+$\hat{L}(s_{1},s_{2},s_{3},k_{1},k_{2},k_{3})=-\sum\limits_{j=1}^{3}k_{j}s_{j}^{2}\log s_{j}^{2}-\tilde{\mu}_{0}(p)\Bigl{(}\sum\limits_{j=1}^{m}k_{j}s_{j}\Bigl{)}^{2}$-
+$-\lambda\Bigl{(}\sum\limits_{j=1}^{3}k_{j}s_{j}^{2}-1\Bigl{)}-\nu\Bigl{(}\sum\limits_{j=1}^{3}k_{j}-m\Bigl{)}.$
 New necessary conditions are
-$-\dfrac{1}{2}\dfrac{\partial L_{(2)}}{\partial c_{l}}=k_{l}c_{l}(\log c_{l}^{2}+\log e)+\tilde{\mu}_{0}(p)k_{l}\Bigl{(}\sum\limits_{j=1}^{3}k_{j}c_{j}\Bigl{)}+\lambda k_{l}c_{l}=0,$
-$-\dfrac{\partial L_{(2)}}{\partial k_{l}}=c_{l}^{2}\log c_{l}^{2}+2\tilde{\mu}_{0}(p)c_{l}\Bigl{(}\sum\limits_{j=1}^{3}k_{j}c_{j}\Bigl{)}+\lambda c_{l}^{2}+\nu=0$
+$-\dfrac{1}{2}\dfrac{\partial\hat{L}}{\partial s_{l}}=k_{l}s_{l}\log s_{l}^{2}+\tilde{\mu}_{0}(p)k_{l}\Bigl{(}\sum\limits_{j=1}^{3}k_{j}s_{j}\Bigl{)}+(\lambda+\log e)k_{l}s_{l}=0,$
+$-\dfrac{\partial\hat{L}}{\partial k_{l}}=s_{l}^{2}\log s_{l}^{2}+2\tilde{\mu}_{0}(p)s_{l}\Bigl{(}\sum\limits_{j=1}^{3}k_{j}s_{j}\Bigl{)}+\lambda s_{l}^{2}+\nu=0$
 for all $1\leq l\leq 3$.
-By transformation, we have
-$-\dfrac{1}{2}c_{l}\dfrac{\partial L_{(2)}}{\partial c_{l}}+k_{l}\dfrac{\partial L_{(2)}}{\partial k_{l}}=k_{l}c_{l}\Bigl{(}c_{l}\log e-\tilde{\mu}_{0}(p)\sum\limits_{j=1}^{3}k_{j}c_{j}\Bigl{)}-\nu k_{l}=0.$
-Consider the next subcases.
+Thus, we have
+$\dfrac{1}{2}s_{l}\dfrac{\partial\hat{L}}{\partial s_{l}}+k_{l}\dfrac{\partial\hat{L}}{\partial k_{l}}=k_{l}s_{l}\Bigl{(}s_{l}\log e-\tilde{\mu}_{0}(p)\sum\limits_{j=1}^{3}k_{j}s_{j}\Bigl{)}-\nu k_{l}=0.$
 Case 2.1. $k_{1},k_{2},k_{3}\neq 0$.
-For all $1\leq l\leq 3$ it holds $c_{l}^{2}\log e=\tilde{\mu}_{0}(p)\Bigl{(}\sum\limits_{j=1}^{3}k_{j}c_{j}\Bigl{)}+\nu$, which can take place, if the set $\{c_{1},c_{2},c_{3}\}$ has no more than $2$ different values.
+For all $1\leq l\leq 3$ it holds $s_{l}^{2}\log e=\tilde{\mu}_{0}(p)\Bigl{(}\sum\limits_{j=1}^{3}k_{j}s_{j}\Bigl{)}+\nu$, which can take place, if the set $\{s_{1},s_{2},s_{3}\}$ has no more than two different values (which is realized, formally, for fractional multiplicities).
 This can be regarded as the next case.
 Case 2.2. $k_{3}=0$.
-Up to permutation, $(z_{j})_{j=1}^{m}=(\underbrace{c_{1},...,c_{1}}_{k_{1}\text{ times}},\underbrace{c_{2},...,c_{2}}_{k_{2}\text{ times}})$, $c_{1}\geq 0,c_{2}\in\mathbb{R}$.
+Up to permutation and the global sign, $(z_{j})_{j=1}^{m}=(\underbrace{s_{1},...,s_{1}}_{k_{1}\text{ times}},\underbrace{s_{2},...,s_{2}}_{k_{2}\text{ times}})$, $s_{1}>0$}, $s_{2}<0$}.
+The components of the minimizer should have different signs; otherwise, changing the sign of a single component would yield a lower value for $\tilde{F}$.
 Denote
@@
 $\tilde{\mu}_{1}(p)=\tilde{\mu}_{1}(p,1/m)-\log m.$
-It is easy to see that for $\alpha$ and $p$ such that $10<\alpha\leq 1/2\leq 1-\alpha\leq p<$ 1 one has $\log\dfrac{p}{\alpha}-\log\dfrac{1-p}{1-\alpha}\geq 0$ and ${\sqrt{(1-p)(1-\alpha)}-\sqrt{p\alpha}\leq 0}$, therefore, $\tilde{\mu}_{0}(p,\alpha)\leq 0$.
-The initial inequality (66) for $k_{1}c_{1}^{2}=t$ is equivalent to
-$-t\log\dfrac{t}{k_{1}/m}-(1-t)\log\dfrac{1-t}{k_{2}/m}\geq\tilde{\mu}_{0}(p,1/m)\Bigl{(}\sqrt{t(k_{1}/m)}\pm\sqrt{(1-t)(k_{2}/m)}\Bigl{)}^{2}-\tilde{\mu}_{1}(p,1/m),$ (90)
-if ±$\pm c_{2}\geq 0$.
-Since $\tilde{\mu}_{0}(p,\alpha)\leq 0$,
-$\tilde{\mu}_{0}(p,1/m)\Bigl{(}\sqrt{t(k_{1}/m)}+\sqrt{(1-t)(k_{2}/m)}\Bigl{)}^{2}-\tilde{\mu}_{1}(p,1/m)\leq$
-$\leq\tilde{\mu}_{0}(p,1/m)\Bigl{(}\sqrt{t(k_{1}/m)}-\sqrt{(1-t)(k_{2}/m)}\Bigl{)}^{2}-\tilde{\mu}_{1}(p,1/m),$
+It is easy to see that for $\alpha$ and $p$ such that $0<\alpha\leq 1/2\leq 1-\alpha\leq p<1$ one has $\log\dfrac{p}{\alpha}-\log\dfrac{1-p}{1-\alpha}\geq 0$ and $\sqrt{(1-p)(1-\alpha)}-\sqrt{p\alpha}\leq 0$}, therefore, $\tilde{\mu}_{0}(p,\alpha)\leq 0$.
+The initial inequality (67) for $k_{1}s_{1}^{2}=t$ is equivalent to
+$t\log\dfrac{t}{k_{1}/m}-(1-t)\log\dfrac{1-t}{k_{2}/m}\geq\tilde{\mu}_{0}(p,1/m)\Bigl{(}\sqrt{t(k_{1}/m)}-\sqrt{(1-t)(k_{2}/m)}\Bigl{)}^{2}-\tilde{\mu}_{1}(p,1/m),$ (92)
 and the statement follows from lemma 5.
-Proof of lemma 5.
-First we will prove that the function $\tilde{\mu}_{1}(p)$ is increasing for $p$ in the range $\Bigl{[}\dfrac{m-1}{m},1\Bigr{)}$. The derivative is
-$\tilde{\mu}_{1}^{\prime}(p)=\log e\dfrac{2(m-2)\sqrt{\dfrac{p(1-p)}{m-1}}-4p+2+\ln p-\ln\dfrac{1-p}{m-1}}{2(m-1)\sqrt{\dfrac{p(1-p)}{m-1}}\Bigg{(}\sqrt{p}+\sqrt{\dfrac{1-p}{m-1}}\Bigg{)}^{2}}.$
-Here $2(m-2)\sqrt{\dfrac{p(1-p)}{m-1}}\geq 0$ and $\ln\dfrac{p}{1-p}-4p+2\geq 0$, therefore, $\tilde{\mu}_{1}^{\prime}(p)\geq 0$.
-So, for $\frac{m-1}{m}\leq p<1$ we have
-$-\tilde{\mu}_{1}(p)\leq-\tilde{\mu}_{1}(1-1/m)=-\dfrac{m-2}{m}\log(m-1)+\log m.$
-For $p(m)\leq p<1$ we have
-$-\tilde{\mu}_{1}(p)\leq-\tilde{\mu}_{1}(p(m))=1,$
-because the condition (71) is equivalent to $t>1$, under which $\tilde{\mu}_{1}\equiv-c_{1}=-1$.
 Lemma 5.
 Let $0\leq\alpha\leq\beta\leq\dfrac{1}{2}\leq 1-\alpha\leq p\leq 1$, $0\leq t\leq 1$. Then
-$-t\log\dfrac{t}{\beta}-(1-t)\log\dfrac{1-t}{1-\beta}\geq\tilde{\mu}_{0}(p,\alpha)\Bigl{(}\sqrt{\beta t}-\sqrt{(1-\beta)(1-t)}\Bigl{)}^{2}-\tilde{\mu}_{1}(p,\alpha).$
+$\tilde{f}(t)\equiv-t\log\dfrac{t}{\beta}-(1-t)\log\dfrac{1-t}{1-\beta}-\Bigg{(}\tilde{\mu}_{0}(p,\alpha)\Bigl{(}\sqrt{\beta t}-\sqrt{(1-\beta)(1-t)}\Bigl{)}^{2}-\tilde{\mu}_{1}(p,\alpha)\Bigg{)}\geq 0.$
 Proof.
@@
 holds in this range. So, the statement of the lemma follows from lemma 3.
-Collecting similar terms,
-$f(t):=2\sqrt{\beta(1-\beta)}\Bigl{(}\mu_{0}(p,\alpha)+\tilde{\mu}_{0}(p,\alpha)\Bigl{)}\sqrt{t(1-t)}-(1-2\beta)\Bigl{(}\mu_{0}(p,\alpha)-\tilde{\mu}_{0}(p,\alpha)\Bigl{)}t+$
+Collecting similar terms and using notation of $f(t)$ from subsection 5.1,
+$\tilde{f}(t)-f(t)=2\sqrt{\beta(1-\beta)}\Bigl{(}\mu_{0}(p,\alpha)+\tilde{\mu}_{0}(p,\alpha)\Bigl{)}\sqrt{t(1-t)}-(1-2\beta)\Bigl{(}\mu_{0}(p,\alpha)-\tilde{\mu}_{0}(p,\alpha)\Bigl{)}t+$
 $+\Bigl{[}(1-\beta)\Bigl{(}\mu_{0}(p,\alpha)-\tilde{\mu}_{0}(p,\alpha)\Bigl{)}-\Bigl{(}\mu_{1}(p,\alpha)-\tilde{\mu}_{1}(p,\alpha)\Bigl{)}\Bigl{]}\geq 0.$
@@
 $\mu_{1}(p,\alpha)-\tilde{\mu}_{1}(p,\alpha)=2\dfrac{\sqrt{p(1-p)\alpha(1-\alpha)}\Bigl{(}\log\dfrac{p}{\alpha}-\log\dfrac{1-p}{1-\alpha}\Bigl{)}}{p-\alpha}.$
-The function $f(t)$ has the form $A\sqrt{t(1-t)}+Bt+C$ with $A\leq 0$ and $B\leq 0$, then the unique minimum point $t_{*}\in[1/2,1]$.
-it has the value $t_{*}=\dfrac{1}{2}\Bigl{(}1-\dfrac{B}{\sqrt{A^{2}+B^{2}}}\Bigl{)}$ and the minimum of the function $f(t)$ is equal to
-$f(t_{*})=\dfrac{B}{2}+C-\dfrac{1}{2}\sqrt{A^{2}+B^{2}},$
+The function $(\tilde{f}-f)(t)$ has the form $A\sqrt{t(1-t)}+Bt+C$ with $A\leq 0$ and $B\leq 0$, then the unique minimum point $t_{*}\in[1/2,1]$.
+It has the value $t_{*}=\dfrac{1}{2}\Bigl{(}1-\dfrac{B}{\sqrt{A^{2}+B^{2}}}\Bigl{)}$ and the minimum of the function $(\tilde{f}-f)(t)$ is equal to
+$(\tilde{f}-f)(t_{*})=\dfrac{B}{2}+C-\dfrac{1}{2}\sqrt{A^{2}+B^{2}},$
 where
@@
 $\dfrac{B}{2}+C=\dfrac{\sqrt{p(1-p)\alpha(1-\alpha)}\Bigl{(}\log\dfrac{p}{\alpha}-\log\dfrac{1-p}{1-\alpha}\Bigl{)}}{(p-\alpha)(p+\alpha-1)}\times$
-×$\times\Bigl{(}(2p-1)-2(p+\alpha-1)\Bigl{)}=$
+$\times\Bigl{(}(2p-1)-2(p+\alpha-1)\Bigl{)}=$
 $=(1-2\alpha)\dfrac{\sqrt{p(1-p)\alpha(1-\alpha)}\Bigl{(}\log\dfrac{p}{\alpha}-\log\dfrac{1-p}{1-\alpha}\Bigl{)}}{(p-\alpha)(p+\alpha-1)}\geq 0,$
 $\dfrac{1}{2}\sqrt{A^{2}+B^{2}}=(1-2\alpha)\dfrac{\sqrt{p(1-p)\alpha(1-\alpha)}\Bigl{(}\log\dfrac{p}{\alpha}-\log\dfrac{1-p}{1-\alpha}\Bigl{)}}{(p-\alpha)(p+\alpha-1)}\times$
-×$\times\sqrt{4\ \dfrac{p(1-p)\beta(1-\beta)}{\alpha(1-\alpha)}+\dfrac{(1-2p)^{2}(1-2\beta)^{2}}{(1-2\alpha)^{2}}}$
+$\times\sqrt{4\ \dfrac{p(1-p)\beta(1-\beta)}{\alpha(1-\alpha)}+\dfrac{(1-2p)^{2}(1-2\beta)^{2}}{(1-2\alpha)^{2}}}$
 Thus,
 $\dfrac{\sqrt{A^{2}+B^{2}}}{B+2C}=\sqrt{4\ \dfrac{p(1-p)\beta(1-\beta)}{\alpha(1-\alpha)}+\dfrac{(1-2p)^{2}(1-2\beta)^{2}}{(1-2\alpha)^{2}}}\leq 1$ (93)
-Let us denote $u(t)=4t(1-t)$. The inequality (91) can be rewritten in the form
+Let us denote u(t)=4t(1t)u(t)=4t(1-t). The inequality (93) can be rewritten in the form
 $\dfrac{u(\beta)u(p)}{u(\alpha)}+\dfrac{(1-u(\beta))(1-u(p))}{1-u(\alpha)}\leq 1,$
 where $u(p)\leq u(\alpha)\leq u(\beta)$. The linear function
-$t\mapsto g(t)=\dfrac{u(\beta)t}{u(\alpha)}+\dfrac{(1-u(\beta))(1-t)}{1-u(\alpha)}$
-has non-negative derivative $\dfrac{u(\beta)}{u(\alpha)}-\dfrac{1-u(\beta)}{1-u(\alpha)}=\dfrac{u(\beta)-u(\alpha)}{u(\alpha)(1-u(\alpha))}$ and $g(u(\alpha))=1$.
+$t\mapsto W(t)=\dfrac{u(\beta)t}{u(\alpha)}+\dfrac{(1-u(\beta))(1-t)}{1-u(\alpha)}$
+has non-negative derivative $\dfrac{u(\beta)}{u(\alpha)}-\dfrac{1-u(\beta)}{1-u(\alpha)}=\dfrac{u(\beta)-u(\alpha)}{u(\alpha)(1-u(\alpha))}$ and $W(u(\alpha))=1$}.
 Therefore, the inequality (93) holds, and lemma has been proven.

### v4

@@
 Computing accessible information for an ensemble of quantum states is a basic problem in quantum information theory.
-The recently obtained optimality criterion [7] when applied to specific ensembles of states leads to nontrivial tight entropy inequalities that are discrete relatives of the famous log-Sobolev inequality.
+We show that the recently obtained optimality criterion [7] when applied to specific ensembles of states leads to nontrivial tight entropy inequalities that are discrete relatives of the famous log-Sobolev inequality.
 In this light, the hypothesis of globally information-optimal measurement for an ensemble of equiangular equiprobable states (quantum pyramids) [2] is reconsidered and the corresponding entropy inequalities are proposed.
-Via the optimality criterion, this suggests also a proof of the conjecture concerning globally information-optimal observables for quantum pyramids put forward in [2].
-1
-Introduction
+Via the optimality criterion, this suggests also an approach to the proof of the conjecture concerning globally information-optimal observables for quantum pyramids put forward in [2].
+1 Introduction
 The problem of accessible information – the maximal information that can be extracted from measurements over an ensemble of quantum states – is a basic problem in quantum information and communication theory, coming back to its origin, see e.g.
@@
 In section 5, the suggestions for proofs proposed by the second author as a response to the hypotheses (31), (72), (80) put forward by the first author, are presented.
-Via the optimality criterion of section 2, this suggests also a proof of the conjecture concerning globally information-optimal observables for quantum pyramids formulated and numerically substantiated in [2].
-2
-Accessible information: a criterion for optimality
+Via the optimality criterion of section 2, this suggests also an approach to the proof of the conjecture concerning globally information-optimal observables for quantum pyramids formulated and numerically substantiated in [2].
+2 Accessible information: a criterion for optimality
 Let $\mathcal{H}$ be a finite dimensional Hilbert space, $\mathfrak{S}=\mathfrak{S}(\mathcal{H})$ the convex set of quantum states (density operators on $\mathcal{H}$).
@@
 Accessible information of the ensemble is defined as
-$A(\mathcal{E})=\sup_{\mathcal{M}}I(\mathcal{E},\mathcal{M}).$
+$A(\mathcal{E})=\sup_{\mathcal{M}}I(\mathcal{E},\mathcal{M}).$ (1)
 It was shown [1] that the supremum is attained on an observable of the form $M_{k}=\left|\varphi_{k}\rangle\langle\varphi_{k}\right|,\,k=1,\dots,n$ with $d\leq n\leq d^{2}$ and linearly independent $M_{k}$ (if the matrices of $\rho_{j}$ in some basis are real then the number of components is reduced to $\frac{d(d+1)}{2}$).
 In particular, the supremum can be replaced with the maximum.
-Introducing the dual ensemble $\mathcal{E}^{\prime}=\left\{p_{k},\sigma_{k}\right\}$ with
+From now on we assume that the average state $\overline{\rho}$ is nondegenerate. Introducing the dual ensemble $\mathcal{E}^{\prime}=\left\{p_{k},\sigma_{k}\right\}$ with
 $\sigma_{k}=p_{k}^{-1}\overline{\rho}^{1/2}M_{k}\overline{\rho}^{1/2}=\left|\phi_{k}\rangle\langle\phi_{k}\right|,$
@@
 $|\phi_{k}\rangle=p_{k}^{-1/2}\overline{\rho}^{1/2}|\varphi_{k}\rangle,$
-and observable $\mathcal{M}^{\prime}$ with the components [^1] In the case $\overline{\rho}$ is degenerate, one should use generalized inverse in the following formula.
+and observable $\mathcal{M}^{\prime}$ with the components
 $M_{j}^{\prime}=\overline{\rho}^{-1/2}\left(\pi_{j}\rho_{j}\right)\overline{\rho}^{-1/2}$ (2)
@@
 The following statements are equivalent:
-Ensemble $\mathcal{E}_{0}^{\prime}=\left\{p_{k}^{0},\sigma_{k}^{0}\right\}$ with $\sigma_{k}^{0}=\left|\phi_{k}^{0}\rangle\langle\phi_{k}^{0}\right|$ is the solution of the problem (5); $\Lambda_{0}$ is the solution of the problem (6);
+$\Lambda_{0}$ is the solution of the problem (7); ensemble $\mathcal{E}_{0}^{\prime}=\left\{p_{k}^{0},\sigma_{k}^{0}\right\}$ with $\sigma_{k}^{0}=\left|\phi_{k}^{0}\rangle\langle\phi_{k}^{0}\right|$ is the solution of the problem (6) (equivalently, observable $M_{k}^{0}=\left|\varphi_{k}^{o}\rangle\langle\varphi_{k}^{0}\right|,\,k=1,\dots,n,$ with
+$|\varphi_{k}^{0}\rangle=\left(p_{k}^{0}\right)^{1/2}\overline{\rho}^{-1/2}|\phi_{k}^{0}\rangle,$
+maximizes the accessible information (1));
 a.
@@
 b.
-$\left[K(\sigma)-\Lambda_{0}\right]\,|\phi_{k}^{0}\rangle=0,\,k=1,\dots,n$.
+$\left[K(\sigma_{k}^{0})-\Lambda_{0}\right]\,|\phi_{k}^{0}\rangle=0,\,k=1,\dots,n$.
 a.
@@
 […]
-The value of the accessible information (4) is
+The solution of the dual problem (7) is unique on the support of the operator $\overline{\rho}$ i.e.
+$\Lambda_{0}\overline{\rho}=\Lambda_{0}^{\prime}\overline{\rho}$ for any two solutions $\Lambda_{0},\Lambda_{0}^{\prime}.$ In particular, if $\overline{\rho}$ is non-degenerate, then the solution of the dual problem is unique.
+According to (5), the value of the accessible information (5) is then
 $A(\mathcal{E})=H(\pi)-\mathrm{Tr\,}\overline{\rho}\Lambda_{0}.$ (10)
@@
 $\sum_{k}K(\sigma_{k}^{0})\,p_{k}^{0}\sigma_{k}^{0}\,=\Lambda_{0}\overline{\rho},$ (13)
-where $\Lambda_{0}$ must be Hermitean operator.
+where $\Lambda_{0}$ must be Hermitean operator. This implies that $\Lambda_{0}\overline{\rho}$ is the same for all solutions of the dual problem.
 By taking into account that $\sigma_{k}^{0}=\left|\phi_{k}^{0}\rangle\langle\phi_{k}^{0}\right|,$ we can rewrite (12) as
@@
 $|e_{j}\rangle=\sqrt{\pi_{j}}\bar{\rho}^{-1/2}|\psi_{j}\rangle;\quad j=1,\dots,d,$ (16)
-by necessity form an orthonormal basis. Then the entropy inequality (7) takes the form
+by necessity form an orthonormal basis (depending on the ensemble $\mathcal{E}$). Then the entropy inequality (8) takes the form
 $\sum_{j=1}^{d}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}\geq\sum_{j=1}^{d}\lambda_{jk}\overline{z_{j}}z_{k},$ (17)
-where $\lambda_{jk}=\langle e_{j}|\Lambda_{0}|e_{k}\rangle$ and $z_{j}=\langle e_{j}|\psi\rangle$ are complex variables subject only to the constraint $\sum_{j=1}^{d}\left|z_{j}\right|^{2}=1$.
+where $\lambda_{jk}=\langle e_{j}|\Lambda_{0}|e_{k}\rangle$ and $z_{j}=\langle e_{j}|\psi\rangle$ are complex variables subject only to the constraint
+$\sum_{j=1}^{m}\left|z_{j}\right|^{2}=1.$ (18)
 The left-hand side is just the Shannon entropy of the p.d.
 $\left\{\left|z_{j}\right|^{2};j=1,\dots,d\right\}$, and such an inequality can be approached with the tools of classical analysis.
-The optimality conditions can be useful in two ways:
+We will need the following
+Lemma 1.
+Let $\lambda_{jk}$ be real numbers. If the inequality
+$\sum_{j=1}^{m}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}+\sum_{j=1}^{d}\lambda_{jk}\overline{z_{j}}z_{k}\leq 0$ (19)
+holds for all real $z_{j}$ satisfying the condition (18), then it holds for all complex $z_{j}$ satisfying the same condition.
+Proof.
+Let $z_{j}=x_{j}+iy_{j}$ for real $x_{j},y_{j}$, then (18) amounts to
+$\sum_{j=1}^{m}x_{j}^{2}+\sum_{j=1}^{m}y_{j}^{2}=1.\quad$ (20)
+Denote $P=\sum_{j=1}^{m}x_{j}^{2},\,1-P=\sum_{j=1}^{m}y_{j}^{2}$. The function $f(\xi)=\xi\log\xi$ is convex, and $\lambda_{jk}$ are real hence
+$\sum_{j=1}^{m}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}+\sum_{j=1}^{d}\lambda_{jk}\overline{z_{j}}z_{k}$ $=$ $\sum_{j=1}^{m}f\left(P\left(\frac{x_{j}}{\sqrt{P}}\right)^{2}+(1-P)\left(\frac{y_{j}}{\sqrt{1-P}}\right)^{2}\right)$
+$+$ $P\sum_{j=1}^{d}\lambda_{jk}x_{j}x_{k}+(1-P)\sum_{j=1}^{d}\lambda_{jk}y_{j}y_{k}$
+$\leq$ $P\left[\sum_{j=1}^{m}f\left(\left(\frac{x_{j}}{\sqrt{P}}\right)^{2}\right)+\sum_{j=1}^{d}\lambda_{jk}x_{j}x_{k}\right]$
+$+$ $(1-P)\left[\sum_{j=1}^{m}f\left(\left(\frac{y_{j}}{\sqrt{1-P}}\right)^{2}\right)+\sum_{j=1}^{d}\lambda_{jk}y_{j}y_{k}\right]\leq 0,$
+because
+$\sum_{j=1}^{m}\left(\frac{x_{j}}{\sqrt{P}}\right)^{2}=\sum_{j=1}^{m}\left(\frac{y_{j}}{\sqrt{1-P}}\right)^{2}=1,\quad$
+and by assumption (19) holds for real $z_{j}$ satisfying (18). ∎
+The optimality conditions can be useful in the following ways:
   1. 1.
@@
 In the next section we illustrate this point with a simple example resulting in the yet nontrivial tight lower bounds for the binary Shannon entropy.
+  3. 3.
+    The uniqueness of the solution of the dual problem (assuming $\bar{\rho}$ non-degenerate) has the following useful consequence.
+Assume that the ensemble $\mathcal{E}=\left\{\pi_{j},\rho_{j}\right\}_{j=1,\dots,m}$ is symmetric in that $\pi_{j}\equiv 1/m$ and there is a group of *-automorphisms $\left\{\alpha_{g};g\in G\right\}$ of the algebra of operators on $\mathcal{H}$ acting transitively on the set of states $\left\{\rho_{j}\right\}_{j=1,\dots,m}$.
+Then $\Lambda_{0}$ is an invariant of this group: $\alpha_{g}^{\ast}\left[\Lambda_{0}\right]=\Lambda_{0};g\in G.$
 3 Example: two pure states
@@
 $K(\sigma_{-}^{0})$ $=$ $-\left|e_{+}\rangle\langle e_{+}\right|\log(1-p)-\left|e_{-}\rangle\langle e_{-}\right|\log p.$
-Multiplying (14) by $\overline{\rho}^{-1/2}$ from the right and taking into account (19) we obtain
-$\Lambda_{0}\overline{\rho}^{1/2}$ $=$ $-\frac{1}{\sqrt{2}}\left(\left|e_{+}\rangle\langle e_{+}\right|+\left|e_{-}\rangle\langle e_{-}\right|\right)\sqrt{p}\log p$
-$-\frac{1}{\sqrt{2}}\left(\left|e_{+}\rangle\langle e_{-}\right|+\left|e_{-}\rangle\langle e_{+}\right|\right)\sqrt{1-p}\log(1-p)$
-$=$ $-\frac{1}{\sqrt{2}}\left|e_{0}\rangle\langle e_{0}\right|\left(\sqrt{p}\log p+\sqrt{1-p}\log(1-p)\right)$
-$-\frac{1}{\sqrt{2}}\left|e_{1}\rangle\langle e_{1}\right|\left(\sqrt{p}\log p-\sqrt{1-p}\log(1-p)\right).$
-Taking into account (18) and noticing that
-$\cos\alpha/2$ $=$ $\frac{1}{\sqrt{2}}\left(\sqrt{p}+\sqrt{1-p}\right)$
-$\sin\alpha/2$ $=$ $\frac{1}{\sqrt{2}}\left(\sqrt{p}-\sqrt{1-p}\right),$
-we obtain
-$\Lambda_{0}=\lambda_{0}(p)\left|e_{0}\rangle\langle e_{0}\right|-\lambda_{1}(p)\left|e_{1}\rangle\langle e_{1}\right|,$ (22)
-where
+The ensemble $\mathcal{E}$ is symmetric with respect to the reflection relative to the axis $|e_{0}\rangle;$ according to the remark 3 at the end of previous section, $\Lambda_{0}$ should also obey this symmetry, hence
+$\Lambda_{0}=\lambda_{0}(p)\left|e_{0}\rangle\langle e_{0}\right|-\lambda_{1}(p)\left|e_{1}\rangle\langle e_{1}\right|.$ (26)
+Solving the conditions (ii.b)
+$\left[K(\sigma_{\pm}^{0})-\Lambda_{0}\right]\,|\phi_{\pm}^{0}\rangle=0$
+and taking into account that
+$\cos\alpha/2=\frac{1}{\sqrt{2}}\left(\sqrt{p}+\sqrt{1-p}\right),\quad\sin\alpha/2=\frac{1}{\sqrt{2}}\left(\sqrt{p}-\sqrt{1-p}\right),$
+results in
 $\lambda_{0}(p)=-\frac{\sqrt{p}\log p+\sqrt{1-p}\log\left(1-p\right)}{\sqrt{p}+\sqrt{1-p}},$
@@
 $\mathrm{Tr\,}\overline{\rho}\Lambda_{0}=\lambda_{0}(p)\cos^{2}\alpha/2-\lambda_{1}(p)\sin^{2}\alpha/2=h(p)$
-in accordance with (17).
+and (10) implies $A(\mathcal{E})=1-h(p)$ in accordance with (21).
 4 Multidimensional case
@@
 The relation between the two parametrizations is
-$a=\sqrt{mr_{1}}+\frac{\sqrt{r_{0}}-\sqrt{r_{1}}}{\sqrt{m}},\quad b=\frac{\sqrt{r_{0}}-\sqrt{r_{1}}}{\sqrt{m}}.$ (37)
+$a=\sqrt{mr_{1}}+\frac{\sqrt{r_{0}}-\sqrt{r_{1}}}{\sqrt{m}},\quad b=\frac{\sqrt{r_{0}}-\sqrt{r_{1}}}{\sqrt{m}},$ (41)
+$1-r_{0}=(m-1)r_{1}=\frac{m-1}{m}(a-b)^{2}.$ (42)
 The cosine between the edges is $\langle\psi_{j}|\psi_{k}\rangle=r_{0}-r_{1}\,,\,j\neq k,$ thus the case $r_{0}>r_{1}\,(b>0)$ corresponds to the acute, $r_{0}=r_{1}\,(b=0)$ – to the orthogonal, and $r_{0}<r_{1}\,(b<0)$ – to the obtuse pyramids.
@@
 Remark 2.
-Operators of the form $c_{1}I+c_{0}|e_{0}\rangle\langle e_{0}|$ constitute an Abelian algebra.
+Operators of the form $c_{1}I+c_{0}|e_{0}\rangle\langle e_{0}|$ constitute Abelian algebra of operators invariant under the group of all *-automorphisms generated by the orthogonal transformations leaving the quantum pyramid invariant.
 One can then easily check that
@@
 (45).
+Remark 3.
+An upper bound and a benchmark for $A(\mathcal{E})$ is the classical capacity of the c-q channel $j\rightarrow|\psi_{j}\rangle\langle\psi_{j}|$:
+$C=\max_{\pi}H\left(\sum_{j=1}^{m}\pi_{j}|\psi_{j}\rangle\langle\psi_{j}|\right)=H(\bar{\rho}).$ (51)
+The last equality follows from the symmetry of the quantum pyramid and concavity of the von Neumann entropy.
+According to (43) the density operator $\bar{\rho}$ has the eigenvalues $r_{0},r_{1}$, the last of multiplicity m1m-1, hence
+$C$ $=$ $-r_{0}\log r_{0}-(m-1)r_{1}\log r_{1}$
+$=$ $-\left(1-\frac{m-1}{m}(a-b)^{2}\right)\log\left(1-\frac{m-1}{m}(a-b)^{2}\right)$
+- $\frac{m-1}{m}(a-b)^{2}\log\frac{(a-b)^{2}}{m},$
+where in the second equality we used (42).
+Unlike the case of accessible information, this expression is the same for all values of the parameters, for acute, obtuse and flat pyramids.
+The ratio $C/A(\mathcal{E})$ gives the gain due to the measurement entanglement at the channel’s output.
+In particular, it tends to the infinity as the pyramid collapses to ort (weak signal) and equals to 1 for rectangular pyramid (ideal channel).
 4.3 Checking the optimality conditions: the case of acute pyramids
@@
 $K\left(\sigma_{0}\right)|\tilde{\phi}_{0}\rangle=\sqrt{r_{0}\left(1-t^{2}\right)}\log m\,|e_{0}\rangle.$
-Motivated by the symmetry of the problem (discrete rotations around the axis $|e_{0}\rangle$) and by the remark 2 in previous subsection we look for $\Lambda_{0}$ in the form
+Motivated by the symmetry of the problem (discrete rotations around the axis $|e_{0}\rangle$) and by the remark 2 we look for $\Lambda_{0}$ in the form
 $\Lambda_{0}=c_{1}I+c_{0}|e_{0}\rangle\langle e_{0}|.$ (56)
@@
 $A(\mathcal{E})=\frac{m\left(1-r_{0}\right)}{m-2}\log(m-1),$
-which is the second line, taking into account that
-$1-r_{0}=(m-1)r_{1}=\frac{m-1}{m}(a-b)^{2}$
-by (34), (37). ∎
+which is the second line, taking into account (42). ∎
 4.4 The case of obtuse pyramids
@@
 Again, two cases are possible: $t=1$ and $t>1.$ In the first case the third equation disappears while the first two lead to the same formulas (58) as in the acute case, however with $a=\sqrt{p\text{ }}$, but $b=-\sqrt{\frac{1-p}{m-1}}.$ Then, denoting $z_{j}=\langle e_{j}|\psi\rangle$, the condition (ii.a) brings us to the inequality:
-For complex $z_{j}$ satisfying
-$\mathit{\sum_{j=1}^{m}\left|z_{j}\right|^{2}=1,}$ (66)
-it holds
+For complex $z_{j}$ satisfying $\mathit{\sum_{j=1}^{m}\left|z_{j}\right|^{2}=1,}$ it holds
 $\sum_{j=1}^{m}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}\geq\tilde{\mu}_{0}(p)\left|\sum_{j=1}^{m}z_{j}\right|^{2}-\tilde{\mu}_{1}(p),$ (72)
@@
 Note that here $\tilde{\mu}_{0}(p)\leq 0,$ which prevents us from passing to the probabilities $t_{j}=\left|z_{j}\right|^{2}$.
-However, as the following lemma shows, we can always restrict to the case of real $z_{j}$.
-Lemma 1.
-Let $\tilde{\mu}_{0},\tilde{\mu}_{1}$ be real numbers. If the inequality
-$\sum_{j=1}^{m}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}+\tilde{\mu}_{0}\left|\sum_{j=1}^{m}z_{j}\right|^{2}\leq\tilde{\mu}_{1}$ (70)
-holds for all real $z_{j}$ satisfying the condition (66), then it holds for all complex $z_{j}$ satisfying the same condition.
-If it holds for all real $z_{j}$ satisfying the condition (66) and
-$\mathit{\sum_{j=1}^{m}z_{j}=0,}$ (71)
-then it holds for all complex $z_{j}$ satisfying the conditions (66), (71).
-Proof.
-Let $z_{j}=x_{j}+iy_{j}$ for real $x_{j},y_{j}$, then (66) amounts to
-$\sum_{j=1}^{m}x_{j}^{2}+\sum_{j=1}^{m}y_{j}^{2}=1,\quad$ (72)
-and (71) amounts to
-$\sum_{j=1}^{m}x_{j}=0,\,\sum_{j=1}^{m}y_{j}=0.$ (73)
-Denote $P=\sum_{j=1}^{m}x_{j}^{2},\,1-P=\sum_{j=1}^{m}y_{j}^{2}$. The function $f(\xi)=\xi\log\xi$ is convex, hence
-$\sum_{j=1}^{m}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}+\tilde{\mu}_{0}\left|\sum_{j=1}^{m}z_{j}\right|^{2}$ $=$ $\sum_{j=1}^{m}f\left(P\left(\frac{x_{j}}{\sqrt{P}}\right)^{2}+(1-P)\left(\frac{y_{j}}{\sqrt{1-P}}\right)^{2}\right)$
-$+$ $P\tilde{\mu}_{0}\left|\sum_{j=1}^{m}\frac{x_{j}}{\sqrt{P}}\right|^{2}+(1-P)\tilde{\mu}_{0}\left|\sum_{j=1}^{m}\frac{y_{j}}{\sqrt{1-P}}\right|^{2}$
-$\leq$ $P\left[\sum_{j=1}^{m}f\left(\left(\frac{x_{j}}{\sqrt{P}}\right)^{2}\right)+\tilde{\mu}_{0}\left|\sum_{j=1}^{m}\frac{x_{j}}{\sqrt{P}}\right|^{2}\right]$
-$+$ $(1-P)\left[\sum_{j=1}^{m}f\left(\left(\frac{y_{j}}{\sqrt{1-P}}\right)^{2}\right)+\tilde{\mu}_{0}\left|\sum_{j=1}^{m}\frac{y_{j}}{\sqrt{1-P}}\right|^{2}\right]$
-$\leq$ $\tilde{\mu}_{1},$
-because
-$\sum_{j=1}^{m}\left(\frac{x_{j}}{\sqrt{P}}\right)^{2}=\sum_{j=1}^{m}\left(\frac{y_{j}}{\sqrt{1-P}}\right)^{2}=1,\quad$
-and by assumption (70) holds for real $z_{j}$ satisfying (66). If additionally the complex $z_{j}$ satisfy (71) then
-$\sum_{j=1}^{m}\frac{x_{j}}{\sqrt{P}}=\sum_{j=1}^{m}\frac{y_{j}}{\sqrt{1-P}}=0,$
-and the second statement of the lemma follows. ∎
+However, as the lemma 1 shows, we can always restrict to the case of real $z_{j}$.
 In (72) the parameter $p$ varies in the range $\max\left\{\frac{m-1}{m},p(m)\right\},1],$ where $p(m)$ is found below.
@@
 $\sum_{j=1}^{m}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}\geq\left\{\begin{array}[]{cc}1,&m\leq 6;\\ \log m-\frac{m-2}{m}\log(m-1),&m\geq 7,\end{array}\right$. (83)
-for complex $z_{j}$ satisfying (81)  (due to the second statement of lemma 1 one can restrict here to real $z_{j}$ ).
+for complex $z_{j}$ satisfying (82) (due to a simple modification of lemma 1 one can restrict here to real $z_{j}$ ).
 Indeed, for $m\leq 6$ we can use the argument leading to inequality (80) with $\sum_{j=1}^{m}z_{j}=0$ and $p=p(m)$.
@@
 Proof.
-On the boundary of the simplex of probability distributions the function $F(t)$ cannot reach a minimum, since the function $s\mapsto u(t_{*}+sa)$,
+The function $F(t)$ cannot reach a minimum on the boundary of the simplex of probability distributions, since the function
 $u(s)=-\sum\limits_{i=1}^{m}(t_{*,i}+sa_{i})\log(t_{*,i}+sa_{i})-\mu_{0}(p)\Bigl(\sum\limits_{i=1}^{m}\sqrt{t_{*,i}+sa_{i}}\Bigl)+\mu_{1}(p)$
-for the boundary point $t_{*}$ and the transversal direction $a$ has the derivative -$\infty$ at zero:
+for a boundary point $t_{*}$ and a transversal direction $a$ has the derivative -$\infty$ at zero:
 $u^{\prime}(0)=\lim\limits_{t\to t_{*}}\Bigg[-\sum\limits_{i\in I_{0}}a_{i}\log t_{i}-C\Bigl(\sum\limits_{i\in I_{0}}\dfrac{a_{i}}{\sqrt{t_{i}}}\Bigl)+O(1)\Bigg],$
-where $I_{0}=\{1\leq j\leq m;\ t_{*,j}=0\}$, $C>0$ is some constant. In this case, $a_{i}>0,\ i\in I_{0}$ (direction inside the simplex), so $u^{\prime}(0)=$-$\infty$.
-It is necessary to check the inequality (83) at the critical points (see Lemma 2), which leads to the inequality
-$\mu_{0}(p,1/m)\Bigl{(}\sqrt{q(k/m)}+\sqrt{(1-q)(1-k/m)}\Bigl{)}^{2}-\mu_{1}(p,1-1/m)\quad$
-$\leq$ $-q\log\dfrac{q}{k/m}-(1-q)\log\dfrac{1-q}{1-k/m},$
-where $\mu_{0}(p,\alpha)$ and $\mu_{1}(p,\alpha)$ are defined in (84) and (85) below. Lemma 3 states the validity of this inequality. ∎
-Consider the functions
-$\mu_{0}(p,\alpha)=\dfrac{\sqrt{\dfrac{p(1-p)}{\alpha(1-\alpha)}}\Bigl{(}\log\dfrac{p}{\alpha}-\log\dfrac{1-p}{1-\alpha}\Big{)}}{\Bigl{(}\sqrt{\alpha p}+\sqrt{(1-\alpha)(1-p)}\Bigl{)}\Bigl{(}\sqrt{\dfrac{p}{\alpha}}-\sqrt{\dfrac{1-p}{1-\alpha}}\Bigl{)}},$ (84)
-$\mu_{1}(p,\alpha)=\dfrac{\sqrt{\dfrac{p}{\alpha}}\log\dfrac{p}{\alpha}-\sqrt{\dfrac{1-p}{1-\alpha}}\log\dfrac{1-p}{1-\alpha}}{\sqrt{\dfrac{p}{\alpha}}-\sqrt{\dfrac{1-p}{1-\alpha}}}$ (85)
-for real variables $p$ and $\alpha$ varying in the range $(0,1)$.
-The following lemmas lead to the proof of the main theorem.
+where $I_{0}=\{1\leq j\leq m;\ t_{*,j}=0\}$, $C>0$ is a constant. In this case, $a_{i}>0,\ i\in I_{0}$ (direction inside the simplex), hence $u^{\prime}(0)=$-$\infty$.∎
+Thus it is sufficient to check the inequality (84) only at the critical points.
 Lemma 2.
-The critical points $t_{*}=(t_{*,1},...,t_{*,m})$ of the function
-$F(t)=H(t)-\Bigl{(}\mu_{0}(p)\Bigl{(}\sum\limits_{k=1}^{m}\sqrt{t_{k}}\Bigl{)}^{2}-\mu_{1}(p)\Bigl{)}$
-have the form
-$t_{*}=\Bigl{(}q/k,...,q/k,(1-q)/(m-k),...,(1-q)/(m-k)\Bigl{)}$
-up to permutations, for some $1\leq k\leq m$ and $q\in[0,1]$ (for $k=m$ it is assumed that $q=1/m$).
-One of the critical points is equal to (p,(1p)/(m1),,(1p)/(m1))(p,(1-p)/(m-1),...,(1-p)/(m-1)), in which the value of the function is equal to 0.
+The critical points $t_{*}=(t_{*,1},...,t_{*,m})$ of the function $F(t)$ have the form
+$t_{*}=\Bigl(q/k,...,q/k,(1-q)/(m-k),...,(1-q)/(m-k)\Bigl)$ (85)
+up to permutations, for some integer $k$, $1\leq k\leq m$ and $q\in[0,1]$ (for $k=m$ it is assumed that $q=1/m$).
+One of the critical points is (p,(1p)/(m1),,(1p)/(m1))(p,(1-p)/(m-1),...,(1-p)/(m-1)), at which the value of the function $F(t)$ is equal to 0.
 Proof.
-Instead of $F(t)$ we consider a function of the form $u(s)=F(t+sa)$, where $t=(t_{1},...,t_{m}),\ a=(a_{1},...,a_{m}),\ \sum\limits_{i}a_{i}=0$, on such a range of values of $s$ that $t+as$ defines a probability distribution.
+Consider the function $u(s)=F(t+sa)$, where $t=(t_{1},...,t_{m}),\ a=(a_{1},...,a_{m}),\ \sum\limits_{i}a_{i}=0$, for a range of $s$ such that $t+as$ defines a probability distribution. Then
 $\dfrac{d}{ds}u(s)$ $=-\sum\limits_{i=1}^{m}a_{i}\log t_{i}-\mu_{0}(p)\Bigl(\sum\limits_{i=1}^{m}\dfrac{a_{i}}{\sqrt{t_{i}}}\Bigl)\Bigl(\sum\limits_{i=1}^{m}\sqrt{t_{i}}\Bigl)$
@@
 $\log t_{*,1}-\log t_{*,i}=\mu_{0}(p)\Bigl(\sum\limits_{k=1}^{m}\sqrt{t_{*,k}}\Bigl)\Bigl(\dfrac{1}{\sqrt{t_{*,i}}}-\dfrac{1}{\sqrt{t_{*,1}}}\Bigl).$
-Let us denote $a=\ln t_{*,1}+\dfrac{\mu_{0}(p)\ln(2)}{\sqrt{t_{*,1}}}\Bigl{(}\sum\limits_{k=1}^{m}\sqrt{t_{*,k}}\Bigl{)},\ C=\mu_{0}(p)\ln(2)\Bigl{(}\sum\limits_{k=1}^{m}\sqrt{t_{*,k}}\Bigl{)}$.
-The equation
-$a-\ln(s)=C/\sqrt{s},\quad s>0,$ (86)
-has no more than two roots, since the derivatives of the right and left parts with respect to $s$ coincide only when $s=C^{2}/4$.
+Let us denote $C_{1}=\mu_{0}(p)\Bigl(\sum\limits_{k=1}^{m}\sqrt{t_{*,k}}\Bigl),C_{2}=\log t_{*,1}+\dfrac{C_{1}}{\sqrt{t_{*,1}}}$. The equation
+$C_{2}-\log s=C_{1}/\sqrt{s},\quad s>0,$ (86)
+has at most two roots, since the derivatives of the right and the left parts with respect to $s$ coincide only when $s=C_{1}^{2}/4$.
 But all the numbers $t_{*,1},...,t_{*,m}$ are solutions of the equation (86).
-Thus, among the values of $\{t_{*,i}\}$ there are no more than two different values, and the lemma is proved.
+Thus, among the values of $\{t_{*,i}\}$ there are no more than two different values, hence the lemma follows.
 ∎
-The proof of the following lemma is based on the behavior of the difference between binary entropy and a function containing both a linear component and a square root term $\sqrt{t(1-t)}$.
-We conduct a qualitative analysis of the minimizer of this function.
+At the critical points (85) the inequality $F(t_{*})\geq 0$ reduces to
+$\mu_{0}(p,1/m)\Bigl(\sqrt{q(k/m)}+\sqrt{(1-q)(1-k/m)}\Bigl)^{2}-\mu_{1}(p,1-1/m)\quad$
+$\leq$ $-q\log\dfrac{q}{k/m}-(1-q)\log\dfrac{1-q}{1-k/m},$
+where $\mu_{0}(p,\alpha)$ and $\mu_{1}(p,\alpha)$ are defined as
+$\mu_{0}(p,\alpha)=\dfrac{\sqrt{\dfrac{p(1-p)}{\alpha(1-\alpha)}}\Bigl(\log\dfrac{p}{\alpha}-\log\dfrac{1-p}{1-\alpha}\Big)}{\Bigl(\sqrt{\alpha p}+\sqrt{(1-\alpha)(1-p)}\Bigl)\Bigl(\sqrt{\dfrac{p}{\alpha}}-\sqrt{\dfrac{1-p}{1-\alpha}}\Bigl)},$ (87)
+$\mu_{1}(p,\alpha)=\dfrac{\sqrt{\dfrac{p}{\alpha}}\log\dfrac{p}{\alpha}-\sqrt{\dfrac{1-p}{1-\alpha}}\log\dfrac{1-p}{1-\alpha}}{\sqrt{\dfrac{p}{\alpha}}-\sqrt{\dfrac{1-p}{1-\alpha}}}$ (88)
+for real variables $p$ and $\alpha$ varying in the range $(0,1)$.
+But this inequality is a special case of lemma 4 below, the proof of which is based on the behavior of the difference between binary entropy and a function containing both a linear component and a square root term $\sqrt{t(1-t)}$.
+We perform a qualitative analysis of the minimizer of this function.
 The global minimum (there can be no more than two local minima) appears on either the interval $[0,1/2]$ or $[1/2,1]$.
@@
 Therefore, minimizing the function reduces to identifying the critical point within the appropriate interval and evaluating the function at that point.
-A detailed proof of this statement will be provided at the end of the section.
-The variables $\alpha,\ p,\beta,\ t$ used in what follows, unless otherwise stated, are such that
+A detailed proof of this statement follows from
+Lemma 3.
+Let $A>0,\ B,\ C\in\mathbb{R}$ (the notation of the coefficients is used only within the framework of this lemma). Then the function
+$g(t)=h(t)-(A\sqrt{t(1-t)}+Bt+C)$ (89)
+[…]
+Proof.
+We have
+$g^{\prime}(0)=-\infty,\quad g^{\prime}(1)=+\infty,$
+therefore local minima are on the interval $(0,1)$.
+In the calculations we will pass to the natural logarithm.
+$g^{\prime}(t)=-\ln\dfrac{t}{1-t}-\dfrac{A\sqrt{t(1-t)}}{2}\Bigl(\dfrac{1}{t}-\dfrac{1}{1-t}\Bigl)-B=$
+$=-\ln\dfrac{t}{1-t}-\dfrac{A(1-2t)}{2\sqrt{t(1-t)}}-B;$
+$g^{\prime\prime}(t)=-\dfrac{1}{t(1-t)}+\dfrac{A}{4(t(1-t))^{3/2}}.$
+$g^{\prime\prime}(t)=0\quad\Leftrightarrow\quad A=4\sqrt{t(1-t)}.$
+Therefore, there are two inflection points located on different sides of $t=1/2$.
+It follows that there are three critical points $t_{1}\leq t_{0}\leq t_{2}$, where $t_{1},\ t_{2}$ are local minimum points.
+The last statement obviously follows from the symmetric case.
+Indeed, for $B<0$, it holds $g(t)\leq g(1-t)$ for $t\in[1/2,1]$, therefore the global minimum lies on the segment $[1/2,1]$.
+Let s, 1ss,\ 1-s ($0<s<1/2$) be zeros of $g^{\prime\prime}(t)$.
+On the segment [s,1s][s,1-s] the function $g(t)^{\prime}$ is decreasing, on the segments [0,s],[1s,1][0,s],\ [1-s,1] the function $g(t)^{\prime}$ increases, and due to symmetry $(g(t)+Bt)^{\prime}$ has root $1/2$.
+Therefore, on the segment $[1/2,1]$ when $B<0$ the function $g(t)$ has no more than one critical point.
+∎
+Let us finally proceed to the proof of lemma 4. The variables $\alpha,\ p,\ \beta,\ t$ used in what follows, unless otherwise stated, are such that
 $\alpha\in{(0,1/2],\quad p\in[1-\alpha,1)},\quad\beta\in[\alpha,1/2],\quad t\in[0,1].$
 Lemma 4.
-For introduced $\alpha,\ p,\ \beta,\ t$ the inequality
+For $\alpha,\ p,\ \beta,\ t$ specified above one has the inequality
 $\mu_{0}(p,\alpha)\Bigl(\sqrt{\beta t}+\sqrt{(1-\beta)(1-t)}\Bigl)^{2}-\mu_{1}(p,\alpha)$
-$\leq$ $-t\log\dfrac{t}{\beta}-(1-t)\log\dfrac{1-t}{1-\beta}$
-holds.
+$\leq$ $-t\log\dfrac{t}{\beta}-(1-t)\log\dfrac{1-t}{1-\beta}.$
 Proof.
 Let
-$f(t)=-t\log\dfrac{t}{\beta}-(1-t)\log\dfrac{1-t}{1-\beta}-\Bigl{(}\mu_{0}(p,\alpha)\Bigl{(}\sqrt{\beta t}+\sqrt{(1-\beta)(1-t)}\Bigl{)}^{2}-\mu_{1}(p,\alpha)\Bigl{)},$
-for which $A=2\mu_{0}(p,\alpha)\sqrt{\beta(1-\beta)}$, $B=\log\dfrac{1-\beta}{\beta}+\mu_{0}(p,\alpha)(2\beta-1)$ in the form (89) (see lemma 4 below).
-The derivative of $f(t)$ has form
+$f(t)=-t\log\dfrac{t}{\beta}-(1-t)\log\dfrac{1-t}{1-\beta}-\Bigl[\mu_{0}(p,\alpha)\Bigl(\sqrt{\beta t}+\sqrt{(1-\beta)(1-t)}\Bigl)^{2}-\mu_{1}(p,\alpha)\Bigl],$
+for which $A=2\mu_{0}(p,\alpha)\sqrt{\beta(1-\beta)}$, $B=\log\dfrac{1-\beta}{\beta}+\mu_{0}(p,\alpha)(2\beta-1)$ in the form (89) of lemma 3.
+The derivative of $f(t)$ is
 $f^{\prime}(t)=\log\dfrac{1-t}{t}+\log\dfrac{\beta}{1-\beta}-\mu_{0}(p,\alpha)\Bigg(\sqrt{\dfrac{\beta}{t}}-\sqrt{\dfrac{1-\beta}{1-t}}\Bigg)\Bigl(\sqrt{\beta t}+\sqrt{(1-\beta)(1-t)}\Bigl).$
@@
     $B\leq 0$.
-    In this case there is minimum point $t_{*}=\beta$ ($f^{\prime}(\beta)=0$) of the function $f(t)$ on $[0,1/2]$ due to the lemma 4 presented later. One has
+    In this case there is minimum point $t_{*}=\beta$ ($f^{\prime}(\beta)=0$) of the function $f(t)$ on $[0,1/2]$ due to lemma 3 presented later. One has
     $f(\beta)=\mu_{1}(p,\alpha)-\mu_{0}(p,\alpha)\geq 0.$ (90)
@@
       $\ln\dfrac{1-\alpha}{\alpha}\geq 2-4\alpha.$
-    Therefore, $V(p)\geq 0$ and the inequality (88) is satisfied.
+    Therefore, $V(p)\geq 0$ and the inequality (91) is established.
   2. 2.
@@
     $\times\Bigl(\sqrt{\beta t_{*}}+\sqrt{(1-\beta)(1-t_{*})}\Bigl).$
-    Then
+    Hence
     $\ln 2\ \dfrac{d}{d\beta}f(t_{*})=\Bigl(\dfrac{t_{*}}{\beta}-\dfrac{1-t_{*}}{1-\beta}\Bigl)+\sqrt{\dfrac{t_{*}(1-t_{*})}{\beta(1-\beta)}}\Bigl(\ln\dfrac{1-t_{*}}{t_{*}}-\ln\dfrac{1-\beta}{\beta}\Bigl)\geq 0$
@@
 Therefore, $f(t_{*})\geq 0$ for all $\alpha\leq\beta\leq 1/2$.
-We obtained the desired inequality $f(t)\geq 0$ for the minimizer $t_{*}$ and it is even more so for the other points $t\in[0,1]$.
-In the particular case of (88), where $\alpha=1/2$, one has an interesting inequality
+We obtained the desired inequality $f(t)\geq 0$ for the minimizer $t_{*}$ and hence for all the other points $t\in[0,1]$.
+In the special case of (91), where $\alpha=1/2$, one has an interesting inequality
 $h_{1/2}(p)\equiv\sqrt{2p}\log(2p)+\sqrt{2(1-p)}\log(2(1-p))\geq 0,$
 which takes place for $p\in(0,1)$. The function $h_{1/2}(p)$ is nonnegative, convex and $h_{1/2}(p)\leq\sqrt{2}$.
-The proof of lemma 3 is based on a simple fact of analysis:
-Lemma 4.
-Let $A>0,\ B,\ C\in\mathbb{R}$ (The notation of the coefficients is used only within the framework of the formulation of the lemma and its proof). Then the function
-$g(t)=h(t)-(A\sqrt{t(1-t)}+Bt+C)$ (89)
-[…]
-Proof.
-$g^{\prime}(0)=-\infty,\quad g^{\prime}(1)=+\infty,$
-therefore local minima are on the interval $(0,1)$.
-In the calculations we will move on to the natural logarithm.
-$g^{\prime}(t)=-\ln\dfrac{t}{1-t}-\dfrac{A\sqrt{t(1-t)}}{2}\Bigl{(}\dfrac{1}{t}-\dfrac{1}{1-t}\Bigl{)}-B=$
-$=-\ln\dfrac{t}{1-t}-\dfrac{A(1-2t)}{2\sqrt{t(1-t)}}-B;$
-$g^{\prime\prime}(t)=-\dfrac{1}{t(1-t)}+\dfrac{A}{4(t(1-t))^{3/2}}.$
-$g^{\prime\prime}(t)=0\quad\Leftrightarrow\quad A=4\sqrt{t(1-t)}.$
-Therefore, there are two inflection points located on different sides of $t=1/2$.
-It follows that there are 3 critical points $t_{1}\leq t_{0}\leq t_{2}$, where $t_{1},\ t_{2}$ are local minimum points.
-The last statement obviously follows from the symmetric case.
-Indeed, for $B<0$, it holds $g(t)\leq g(1-t)$ for $t\in[1/2,1]$, therefore the global minimum lies on the segment $[1/2,1]$.
-Let s, 1ss,\ 1-s ($0<s<1/2$) be zeros of $g^{\prime\prime}(t)$.
-On the segment [s,1s][s,1-s] the function $g(t)^{\prime}$ is decreasing, on the segments [0,s],[1s,1][0,s],\ [1-s,1] the function $g(t)^{\prime}$ increases, and due to symmetry $(g(t)+Bt)^{\prime}$ has root $1/2$.
-Therefore, on the segment $[1/2,1]$ when $B<0$ the function $g(t)$ has no more than one critical point.
-∎
 5.2 The case of theorem 3
@@
 ∎
-Acknowledgment. This work is supported by Russian Science Foundation under the grant No 24-11-00145. The authors are grateful to S.V. Grishin and D.A. Kronberg for providing numerical simulations.
+Acknowledgment. The authors are grateful to S.V. Grishin and D.A. Kronberg for providing numerical simulations.

### v5

@@
 Computing accessible information for an ensemble of quantum states is a basic problem in quantum information theory.
-We show that the recently obtained optimality criterion [7] when applied to specific ensembles of states leads to nontrivial tight entropy inequalities that are discrete relatives of the famous log-Sobolev inequality.
-In this light, the hypothesis of globally information-optimal measurement for an ensemble of equiangular equiprobable states (quantum pyramids) [2] is reconsidered and the corresponding entropy inequalities are proposed.
-Via the optimality criterion, this suggests also an approach to the proof of the conjecture concerning globally information-optimal observables for quantum pyramids put forward in [2].
+We show that the recently obtained optimality criterion (A.S.
+Holevo, Lobachevskii J.
+Math., 43:7 (2022), 1646-1650), when applied to specific ensembles of states leads to nontrivial tight entropy inequalities that are discrete relatives of the famous log-Sobolev inequality.
+In this light, the hypothesis of globally information-optimal measurement for an ensemble of equiangular equiprobable states (quantum pyramids) (B.-G.
+Englert and J.
+Řeháček, J.
+Mod.
+Optics 57 N3 (2010) 218-226) is reconsidered and the corresponding entropy inequalities are proposed.
+Via the optimality criterion, this suggests also an approach to the proof of the conjectures concerning globally information-optimal observables for quantum pyramids.
+Keywords and phrases: quantum state ensemble, accessible information, convex optimization, quantum pyramid, tight lower bound for Shannon entropy.
 1 Introduction
@@
 where $\delta_{\sigma_{k}}(d\sigma)$ is the $\delta$-measure concentrated in the point $\sigma_{k}\in\mathfrak{S.}$ then the problem (6) can be written as [^1] We refer to [7] for rigorous general definition of the traced integral and formulations of the optimality conditions.
-$\mathrm{Tr}\int_{\mathfrak{S}}\,K(\sigma)\Pi(d\sigma)$ $\longrightarrow$ $\min$ (11)
+$\mathrm{Tr}\int_{\mathfrak{S}}\,K(\sigma)\Pi(d\sigma)$ $\longrightarrow$ $\min$
 […]
@@
 and (12) as
-$\sum_{k}K(\sigma_{k}^{0})\,p_{k}^{0}\left|\phi_{k}^{0}\rangle\langle\phi_{k}^{0}\right|=\Lambda_{0}\overline{\rho}.$ (15)
+$\sum_{k}K(\sigma_{k}^{0})\,p_{k}^{0}\left|\phi_{k}^{0}\rangle\langle\phi_{k}^{0}\right|=\Lambda_{0}\overline{\rho}.$
 Taking into account (4), the condition (ii.a) becomes
@@
 where the vectors
-$|e_{j}\rangle=\sqrt{\pi_{j}}\bar{\rho}^{-1/2}|\psi_{j}\rangle;\quad j=1,\dots,d,$ (16)
+$|e_{j}\rangle=\sqrt{\pi_{j}}\bar{\rho}^{-1/2}|\psi_{j}\rangle;\quad j=1,\dots,d,$
 by necessity form an orthonormal basis (depending on the ensemble $\mathcal{E}$). Then the entropy inequality (8) takes the form
-$\sum_{j=1}^{d}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}\geq\sum_{j=1}^{d}\lambda_{jk}\overline{z_{j}}z_{k},$ (17)
+$\sum_{j=1}^{d}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}\geq\sum_{j=1}^{d}\lambda_{jk}\overline{z_{j}}z_{k},$
 where $\lambda_{jk}=\langle e_{j}|\Lambda_{0}|e_{k}\rangle$ and $z_{j}=\langle e_{j}|\psi\rangle$ are complex variables subject only to the constraint
@@
 Let $z_{j}=x_{j}+iy_{j}$ for real $x_{j},y_{j}$, then (14) amounts to
-$\sum_{j=1}^{m}x_{j}^{2}+\sum_{j=1}^{m}y_{j}^{2}=1.\quad$ (20)
+$\sum_{j=1}^{m}x_{j}^{2}+\sum_{j=1}^{m}y_{j}^{2}=1.$
 Denote $P=\sum_{j=1}^{m}x_{j}^{2},\,1-P=\sum_{j=1}^{m}y_{j}^{2}$. The function $f(\xi)=\xi\log\xi$ is convex, and $\lambda_{jk}$ are real hence
@@
 with
-$\overline{\rho}^{1/2}=\cos\alpha/2\left|e_{0}\rangle\langle e_{0}\right|+\sin\alpha/2\left|e_{1}\rangle\langle e_{1}\right|$ (22)
+$\overline{\rho}^{1/2}=\cos\alpha/2\left|e_{0}\rangle\langle e_{0}\right|+\sin\alpha/2\left|e_{1}\rangle\langle e_{1}\right|$
 so that
-$\frac{1}{\sqrt{2}}\overline{\rho}^{-1/2}|\psi_{\pm}\rangle=|e_{\pm}\rangle$ (23)
+$\frac{1}{\sqrt{2}}\overline{\rho}^{-1/2}|\psi_{\pm}\rangle=|e_{\pm}\rangle$
 and
@@
 Introducing the important parameter
-$p=\frac{1+\sin\alpha}{2},\quad\frac{1}{2}\leq p\leq 1,$ (25)
+$p=\frac{1+\sin\alpha}{2},\quad\frac{1}{2}\leq p\leq 1,$
 we have
@@
 provided $p>1/2$.
-Then the conditions (ii.b) are fulfilled while (ii.a) reduces to (8) which via (24), (26) amounts to the following parametric family of tight lower bounds for the classical binary entropy  [^3] See the dynamical plot of the inequality at https://www.desmos.com/Calculator/11ikk1i8fb
+Then the conditions (ii.b) are fulfilled while (ii.a) reduces to (8) which via (17), (18) amounts to the following parametric family of tight lower bounds for the classical binary entropy
 Proposition 1.
@@
 $-\sum_{j=1}^{m}t_{j}\log t_{j}$ $\geq$ $\frac{\log\left(m-1\right)}{m-2}\left(\sum_{j=1}^{m}\sqrt{t_{j}}\right)^{2}-\left[\frac{m\log\left(m-1\right)}{m-2}-\log m\right]\quad$ (26)
-$=$ $\log m-\frac{m\log\left(m-1\right)}{m-2}\left[1-B(P;P_{U})^{2}\right],$ (35)
+$=$ $\log m-\frac{m\log\left(m-1\right)}{m-2}\left[1-B(P;P_{U})^{2}\right],$
 where
@@
 $|\psi_{j}\rangle=a|e_{j}\rangle+b\sum_{k\neq j}|e_{k}\rangle,\quad j=1,\dots,m,$ (30)
-$a\geq 0,\quad a^{2}+(m-1)b^{2}=1.$ (40)
+$a^{2}+(m-1)b^{2}=1,\quad a\geq\frac{1}{\sqrt{m}}.$
 The relation between the two parametrizations is
@@
 $1-r_{0}=(m-1)r_{1}=\frac{m-1}{m}(a-b)^{2}.$ (32)
-The cosine between the edges is $\langle\psi_{j}|\psi_{k}\rangle=r_{0}-r_{1}\,,\,j\neq k,$ thus the case $r_{0}>r_{1}\,(b>0)$ corresponds to the acute, $r_{0}=r_{1}\,(b=0)$ – to the orthogonal, and $r_{0}<r_{1}\,(b<0)$ – to the obtuse pyramids.
-The extreme cases are $r_{1}=0\,\left(a=b=\sqrt{\frac{1}{m}}\right)$ – the ort $|\psi_{j}\rangle\equiv|e_{0}\rangle,$ and $r_{0}=0\,\left(a=\sqrt{\frac{m-1}{m}},b=-\sqrt{\frac{1}{m\left(m-1\right)}}\right)$ – the flat pyramid.
+The cosine of angle between any two different edges is
+$\xi\equiv\langle\psi_{j}|\psi_{k}\rangle=r_{0}-r_{1}=2ab+(m-2)b^{2}\,,\quad\,j\neq k,$ (33)
+thus the case $r_{0}>r_{1}\,(b>0)$ corresponds to the acute ($0<\xi<1$), $r_{0}=r_{1}\,(b=0)$ – to the orthogonal ($\xi=0$), and $r_{0}<r_{1}\,(b<0)$ – to the obtuse $\left(-\frac{1}{\left(m-1\right)}<\xi<0\right)$ pyramids.
+The extreme cases are $r_{1}=0\,\left(a=b=\sqrt{\frac{1}{m}}\right)$ – the ort $|\psi_{j}\rangle\equiv|e_{0}\rangle$ ($\xi=1$), and $r_{0}=0\,\left(a=\sqrt{\frac{m-1}{m}},b=-\sqrt{\frac{1}{m\left(m-1\right)}}\right)$ – the flat $\left(\xi=-\frac{1}{\left(m-1\right)}\right)$ pyramid.
 We will consider the ensemble $\mathcal{E}=\left\{\pi_{j},\rho_{j}\right\}_{j=1,\dots,m}$ with $\pi_{j}=\frac{1}{m},\,\rho_{j}=|\psi_{j}\rangle\langle\psi_{j}|$ with the average state
@@
 $+$ $\frac{m-1}{m}\left(\sqrt{r_{0}}-\sqrt{r_{1}}\right)^{2}\log\left(\sqrt{r_{0}}-\sqrt{r_{1}}\right)^{2}$
-$=$ $\log m+a^{2}\log a^{2}+\left(m-1\right)b^{2}\log b^{2}.$ (50)
+$=$ $\log m+a^{2}\log a^{2}+\left(m-1\right)b^{2}\log b^{2}.$
 Note that in the case $t_{a}(m)=1$ the conjectured information-optimal observable $M_{k}=|{e}_{k}\rangle\langle{e}_{k}|,\,k=1,\dots,m,$ is optimal for the Bayes discrimination problem for the ensemble $\mathcal{E}$ [4], sec.
@@
 An upper bound and a benchmark for $A(\mathcal{E})$ is the classical capacity of the c-q channel $j\rightarrow|\psi_{j}\rangle\langle\psi_{j}|$:
-$C=\max_{\pi}H\left(\sum_{j=1}^{m}\pi_{j}|\psi_{j}\rangle\langle\psi_{j}|\right)=H(\bar{\rho}).$ (51)
+$C=\max_{\pi}H\left(\sum_{j=1}^{m}\pi_{j}|\psi_{j}\rangle\langle\psi_{j}|\right)=H(\bar{\rho}).$
 The last equality follows from the symmetry of the quantum pyramid and concavity of the von Neumann entropy.
@@
 Having in mind checking the optimality condition (ii.b) we compute
-$K\left(\sigma_{k}\right)|\tilde{\phi}_{k}\rangle$ $=$ $-\sum_{j}|e_{j}\rangle\left(\sqrt{r_{1}}\delta_{jk}+\frac{\sqrt{r_{0}}t-\sqrt{r_{1}}}{m}\right)\,\log\,\frac{\left|\sqrt{r_{1}}\delta_{jk}+\frac{\sqrt{r_{0}}t-\sqrt{r_{1}}}{m}\right|^{2}}{r_{1}+\frac{r_{0}t^{2}-r_{1}}{m}}$ (53)
+$K\left(\sigma_{k}\right)|\tilde{\phi}_{k}\rangle$ $=$ $-\sum_{j}|e_{j}\rangle\left(\sqrt{r_{1}}\delta_{jk}+\frac{\sqrt{r_{0}}t-\sqrt{r_{1}}}{m}\right)\,\log\,\frac{\left|\sqrt{r_{1}}\delta_{jk}+\frac{\sqrt{r_{0}}t-\sqrt{r_{1}}}{m}\right|^{2}}{r_{1}+\frac{r_{0}t^{2}-r_{1}}{m}}$
 $=$ $A(t)|e_{k}\rangle+B(t)\sqrt{m}|e_{0}\rangle;\quad k=1,\dots,m,$
@@
 Introducing complex numbers $z_{j}=\langle e_{j}|\psi\rangle,$ so that $\sum_{j=1}^{m}\left|z_{j}\right|^{2}=1,$ the optimality condition (ii.a) amounts to the inequality
-$\sum_{j=1}^{m}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}\geq\mu_{0}(p)\left|\sum_{j=1}^{m}z_{j}\right|^{2}-\mu_{1}(p).$ (63)
+$\sum_{j=1}^{m}\left|z_{j}\right|^{2}\log\left|z_{j}\right|^{2}\geq\mu_{0}(p)\left|\sum_{j=1}^{m}z_{j}\right|^{2}-\mu_{1}(p).$
 Since $\mu_{0}(p)\geq 0,$ this is equivalent to (23) where $t_{j}=\left|z_{j}\right|^{2}$ .
@@
 When $r_{0}<r_{1}$ $(b<0)$ the pyramid is obtuse, with extreme flat case $r_{0}=0$ when all the state vectors $|\psi_{j}\rangle$ lie in the hyperplane orthogonal to $|e_{0}\rangle.$ The hypothetical optimal observable is then drastically different.
-To describe it let us first consider the flat pyramid in the case $m=3$ where $|\psi_{j}\rangle;j=1,2,3,$ are essentially the “trine” of equiangular unit vectors in two-dimensional real plane, cf.
+To describe it let us first briefly survey the flat pyramid in the case $m=3$ where $|\psi_{j}\rangle;j=1,2,3,$ are essentially the “trine” of equiangular unit vectors in two-dimensional real plane, cf.
 [5].
 Then the optimal observable is $\mathcal{M}=\left\{M_{k};k=1,2,3\right\},$ where $M_{k}=\frac{2}{3}|\psi_{k}^{\perp}\rangle\langle\psi_{k}^{\perp}|,$ and $|\psi_{k}^{\perp}\rangle$ are the unit vectors in the plane such that $\langle\psi_{k}^{\perp}|\psi_{k}\rangle=0;k=1,2,3$ [14].
-In other words, $\mathcal{M}$ is “unambiguous discrimination” measurement (MUD) for the states $\rho_{j}=|\psi_{j}\rangle\langle\psi_{j}|,\,j=1,2,3.$ The analysis based on our optimality criterion shows that the conditions (ii) of the theorem are fulfilled with $\Lambda_{0}=I,$ with the condition (ii.a) equivalent to the entropy inequality
+In other words, $\mathcal{M}$ is “unambiguous discrimination” measurement (MUD) for the states $\rho_{j}=|\psi_{j}\rangle\langle\psi_{j}|,\,j=1,2,3.$ In the section 4.5 we perform the analysis based on our optimality criterion showing that the conditions (ii) of the theorem are fulfilled with $\Lambda_{0}=I,$ with the condition (ii.a) equivalent to the entropy inequality
 $\sum_{j=1}^{3}t_{j}\log t_{j}\geq 1;\quad t_{j}=\frac{2}{3}\cos^{2}\left(\alpha+\frac{2\pi j}{3}\right),$ (53)
@@
 Shor in [15], the conjectured optimal observable has 6 outcomes and is a statistical mixture of the MUD described above and the SRM $M_{k}=|{e}_{k}\rangle\langle{e}_{k}|,\,k=1,2,3.$ The SRM becomes optimal for obtuse pyramids with $\bar{r}_{0}\leq r_{0}\leq r_{1}=1/3.$
-The generalization of this picture to the case $m\geq 3$ based on numerical studies was suggested and elaborated in [2].
+A generalization of this picture to the case $m\geq 3$ based on numerical studies was suggested and elaborated in [2].
 Let us apply our optimality criterion to derive the corresponding entropy inequalities and to obtain additional insight.
@@
 If $m\leq 6$ then $p(m)>\frac{m-1}{m}$ and the inequality (59) holds for $p\in\left[p(m),1\right]$, yielding the inequality (67) for $p\in\left[\frac{m-1}{m},p(m)\right]$.
-The suggestion for a proof of the inequalities is given in section 5.2.
+The suggestion for a proof of the inequalities combining analytical argument with computer-aided reasoning is given in section 5.2.
 Corollary 2.
@@
 4.5 The case of flat pyramid
-The considerations above formally do not include the case of flat pyramid with $r_{0}=0,r_{1}=1/(m-1).$ Then the average state of the ensemble $\mathcal{E}$
-$\overline{\rho}=\frac{1}{m-1}\left(I-|0\rangle\langle 0|\right)$
+The considerations above formally do not include the case of flat pyramid with $r_{0}=0,r_{1}=1/(m-1).$ Then the edges (28) of the pyramid
+$|\psi_{j}\rangle=\sqrt{\frac{m-1}{m}}|e_{j}\rangle-\sqrt{\frac{1}{m}}|e_{0}\rangle;\quad j=1,\dots,m$
+all lie in the hyperplane $\mathcal{H}_{0}=\left\{\psi:\langle e_{0}|\psi\rangle=0\right\}$.
+The vectors are linearly dependent and their convex hull is a regular symplex in $\mathcal{H}_{0}.$ The average state of the ensemble $\mathcal{E}$
+$\overline{\rho}=\frac{1}{m-1}\left(I-|e_{0}\rangle\langle e_{0}|\right)$
 is degenerate and we should use generalized inverse in the formula (41), resulting in
 $M_{j}^{\prime}=|\tilde{\psi}_{j}\rangle\langle\tilde{\psi}_{j}|,\quad j=1,\dots,m,$
-where $|\tilde{\psi}_{j}\rangle=|e_{j}\rangle-\sqrt{\frac{1}{m}}|e_{0}\rangle.$ Denoting $z_{j}=\langle e_{j}|\psi\rangle$ we have
+where
+$|\tilde{\psi}_{j}\rangle=|e_{j}\rangle-\sqrt{\frac{1}{m}}|e_{0}\rangle=\sqrt{\frac{m-1}{m}}|\psi_{j}\rangle.$ (69)
+Denoting $z_{j}=\langle e_{j}|\psi\rangle$ we have
 $\sum_{j=1}^{m}\left|z_{j}\right|^{2}=1,\quad\sum_{j=1}^{m}z_{j}=0.$ (70)
-Indeed, the unit vector $|\psi\rangle$ lies in the hyperplane $\langle e_{0}|\psi\rangle=0$, hence the second equality follows.
+Indeed, the unit vector $|\psi\rangle$ lies in the hyperplane $\langle e_{0}|\psi\rangle=0$i equality follows.
 Arguing as in the previous section, we suggest that the optimal observable is given by (54) with $t=1$ for $m\geq 7$ and $t=\infty$ for $m\leq 6$. The corresponding tight entropy inequality is
@@
 $A(\mathcal{E})=\log m-\mathrm{Tr\,}\Lambda_{0}\bar{\rho}=\left\{\begin{array}[]{cc}\log\frac{m}{2},&m\leq 6;\\ \frac{m-2}{m}\log(m-1),&m\geq 7.\end{array}\right$.
-For $m=3$ the inequality (83) reduces to (66). Indeed, as shown below, in (83) one can restrict to real $z_{j}$. Then $|\psi\rangle$ is a vector in the real plane $\langle e_{0}|\psi\rangle=0$ and
-$z_{j}=\langle e_{j}|\psi\rangle=\langle\tilde{\psi}_{j}|\psi\rangle=\|\tilde{\psi}_{j}\|\,\cos\left(\alpha+\frac{2\pi j}{3}\right),\quad j=1,2,3,$
-where $\||\tilde{\psi}_{j}\rangle\|=\||e_{j}\rangle-\sqrt{1/3}|e_{0}\rangle\|=\sqrt{2/3}$, hence (83) is the same as (66).
+For $m=3$ the inequality (71) reduces to (53). Indeed, as shown in lemma 1, one can restrict to real $z_{j}$ in (71). Then $|\psi\rangle$ is a vector in the real plane $\mathcal{H}_{0}$ and by (69)
+$z_{j}=\langle e_{j}|\psi\rangle=\langle\tilde{\psi}_{j}|\psi\rangle=\sqrt{2/3}\langle\psi_{j}|\psi\rangle=\sqrt{2/3}\cos\alpha_{j},\quad j=1,2,3,$
+where $\alpha_{j}$ is the angle between the vector $|\psi\rangle$ and the ensemble vector $|\psi_{j}\rangle$ in the plane $\mathcal{H}_{0}.$ The angle between any two different vectors of the ensemble is $2\pi/3$ because $\langle\psi_{j}|\psi_{k}\rangle=-1/2$ according to (33).
+Therefore $\alpha_{j}=\alpha+\frac{2\pi j}{3},\quad j=1,2,3,$ thus we come to (53).
+Let us also explain how (53) follows from the result of the paper [14]. A simple calculation shows that (53) is equivalent to
+$\log 3-\frac{1}{3}\sum_{j=1}^{3}\left(1+\cos\left(2\alpha+\frac{4\pi j}{3}\right)\right)\log\left(1+\cos\left(2\alpha+\frac{4\pi j}{3}\right)\right)\geq\log 2;$
+A special case of lemma 3 in [14] (with $M=3$) implies that the deductible in the lefthand side is maximized for $\alpha=\pi/2+\pi k$ (giving in fact the accessible information $\log\left(3/2\right)$).
+The proof in [14] is based on the decomposition of the function $\left(1+x\right)\log\left(1+x\right)$ into the power series and a clever manipulation with the terms of the decomposition.
 5 Suggestions for proofs
@@
 where $\mu_{0}(p,\alpha)$ and $\mu_{1}(p,\alpha)$ are defined as
-$\mu_{0}(p,\alpha)=\dfrac{\sqrt{\dfrac{p(1-p)}{\alpha(1-\alpha)}}\Bigl(\log\dfrac{p}{\alpha}-\log\dfrac{1-p}{1-\alpha}\Big)}{\Bigl(\sqrt{\alpha p}+\sqrt{(1-\alpha)(1-p)}\Bigl)\Bigl(\sqrt{\dfrac{p}{\alpha}}-\sqrt{\dfrac{1-p}{1-\alpha}}\Bigl)},$ (87)
-$\mu_{1}(p,\alpha)=\dfrac{\sqrt{\dfrac{p}{\alpha}}\log\dfrac{p}{\alpha}-\sqrt{\dfrac{1-p}{1-\alpha}}\log\dfrac{1-p}{1-\alpha}}{\sqrt{\dfrac{p}{\alpha}}-\sqrt{\dfrac{1-p}{1-\alpha}}}$ (88)
+$\mu_{0}(p,\alpha)=\dfrac{\sqrt{\dfrac{p(1-p)}{\alpha(1-\alpha)}}\Bigl(\log\dfrac{p}{\alpha}-\log\dfrac{1-p}{1-\alpha}\Big)}{\Bigl(\sqrt{\alpha p}+\sqrt{(1-\alpha)(1-p)}\Bigl)\Bigl(\sqrt{\dfrac{p}{\alpha}}-\sqrt{\dfrac{1-p}{1-\alpha}}\Bigl)},$
+$\mu_{1}(p,\alpha)=\dfrac{\sqrt{\dfrac{p}{\alpha}}\log\dfrac{p}{\alpha}-\sqrt{\dfrac{1-p}{1-\alpha}}\log\dfrac{1-p}{1-\alpha}}{\sqrt{\dfrac{p}{\alpha}}-\sqrt{\dfrac{1-p}{1-\alpha}}}$
 for real variables $p$ and $\alpha$ varying in the range $(0,1)$.
@@
 Namely, if the parameter $p<1$ satisfies $p\geq\max\left\{\dfrac{m-1}{m},p(m)\right\}$ with $p(m)$ defined in (65), then for complex numbers $z_{1},...,z_{m}$ satisfying (14) the inequality
-$\tilde{F}(z)\equiv-\sum\limits_{j=1}^{m}|z_{j}|^{2}\log|z_{j}|^{2}-\tilde{\mu}_{0}(p)\Bigl|\sum\limits_{j=1}^{m}z_{j}\Bigl|^{2}+\tilde{\mu}_{1}(p)\geq 0.$ (92)
+$\tilde{F}(z)\equiv-\sum\limits_{j=1}^{m}|z_{j}|^{2}\log|z_{j}|^{2}-\tilde{\mu}_{0}(p)\Bigl|\sum\limits_{j=1}^{m}z_{j}\Bigl|^{2}+\tilde{\mu}_{1}(p)\geq 0.$
 holds with $\tilde{\mu}_{0}(p),\tilde{\mu}_{1}(p)$ defined in (60), (61).
-Suggestion for proof.
-As it has been established in lemma 1, a real case is enough for analysis.
-For real vectors $(z_{j})_{j=1}^{m}$ from the unit sphere let us consider the Lagrange function
+Suggestion for proof with computer verification.
+As it has been established in lemma 1, it suffices to consider real vectors.
+For real vectors $(z_{j})_{j=1}^{m}$ on the unit sphere let us consider the Lagrange function
 $L(z_{1},...,z_{k})=-\sum\limits_{k=1}^{m}z_{k}^{2}\log z_{k}^{2}-\tilde{\mu}_{0}(p)\Bigl(\sum\limits_{k=1}^{m}z_{k}\Bigl)^{2}+\tilde{\mu}_{1}(p)-\lambda\Bigl(\sum\limits_{k=1}^{m}z_{k}^{2}-1\Bigl).$
-The necessary condition is
+The first-order conditions yield
 $\dfrac{1}{2}\dfrac{\partial L}{\partial z_{k}}=z_{k}\log z_{k}+\tilde{\mu}_{0}(p)\Bigl(\sum\limits_{j=1}^{m}z_{j}\Bigl)+(\lambda+\log e)z_{k}=0,\quad 1\leq k\leq m,$
-and the set $\{z_{j}\}$ has at most three different values $s_{1},s_{2},s_{3}$ with multiplicities $k_{1},k_{2},k_{3}\geq 0$ ($k_{1}+k_{2}+k_{3}=m$), since the equation
+and the set $\{z_{j}\}$ has at most three different values $s_{1},s_{2},s_{3}$ with multiplicities $k_{1},k_{2},k_{3}\geq 0$ ($k_{1}+k_{2}+k_{3}=m$), since equation
 $z\log z^{2}+\tilde{\mu}_{0}(p)\Bigl(\sum\limits_{j=1}^{m}z_{j}\Bigl)+(\lambda+\log e)z=0$ (78)
-for a variable $z$ has at most three real solutions.
-The fact the equation of the form $z\log z^{2}+az+b=0$ for real $a$ and $b$ has no more then three real roots was actually used in proof of lemma 2 for equation (86) which is similar to (93) with $\leq 2$ roots.
+in a variable $z$ has at most three real solutions.
+The fact that the equation of the form $z\log z^{2}+az+b=0$ for real $a$ and $b$ has no more than three real roots was actually used in the proof of lemma 2 for equation (74) which is similar to (78) with $\leq 2$ roots.
 However, here we consider the entire real line, so the maximum number of real roots is three.
-Consider the following cases.
+We consider two cases.
 Case 1. The minimizer has a zero component.
 Then $\lambda+\log e=0$ and the equation (78) has roots 0 and $\pm s$.
-Therefore the minimizer of the function $\tilde{F}(z)$ should have the form $(\underbrace{s,-s,...,s,-s}_{k\text{ times}},0,...,0)$, where $2ks^{2}=1$ giving the value $2ks^{2}\log s^{2}+\tilde{\mu}_{1}(p)=\log 2k+\tilde{\mu}_{1}(p)$ which has the minimum equal to $1+\tilde{\mu}_{1}(p)$ for $k=1$.
+Therefore the minimizer of $\tilde{F}(z)$ must be of the form $(\underbrace{s,-s,...,s,-s}_{k\text{ times}},0,...,0)$, where $2ks^{2}=1$ giving the value $2ks^{2}\log s^{2}+\tilde{\mu}_{1}(p)=\log 2k+\tilde{\mu}_{1}(p)$ which has the minimum equal to $1+\tilde{\mu}_{1}(p)$ for $k=1$.
 The value $1+\tilde{\mu}_{1}(p)$ is nonnegative for the following reason.
-First we will prove that the function $\tilde{\mu}_{1}(p)$ is increasing for $p$ in the range $\Bigl[\dfrac{m-1}{m},1\Bigr)$. The derivative is
+First, we prove that the function $\tilde{\mu}_{1}(p)$ is increasing for $p$ in the range $\Bigl[\dfrac{m-1}{m},1\Bigr)$. The derivative is
 $\tilde{\mu}_{1}^{\prime}(p)=\log e\dfrac{2(m-2)\sqrt{\dfrac{p(1-p)}{m-1}}-4p+2+\ln p-\ln\dfrac{1-p}{m-1}}{2(m-1)\sqrt{\dfrac{p(1-p)}{m-1}}\Bigg(\sqrt{p}+\sqrt{\dfrac{1-p}{m-1}}\Bigg)^{2}}.$
-Here $2(m-2)\sqrt{\dfrac{p(1-p)}{m-1}}\geq 0$ and $\ln\dfrac{p}{1-p}-4p+2\geq 0$, therefore, $\tilde{\mu}_{1}^{\prime}(p)\geq 0$.
+Here, $2(m-2)\sqrt{\dfrac{p(1-p)}{m-1}}\geq 0$ and $\ln\dfrac{p}{1-p}-4p+2\geq 0$, hence $\tilde{\mu}_{1}^{\prime}(p)\geq 0$.
 Thus for $\frac{m-1}{m}\leq p<1$ we have
 $\tilde{\mu}_{1}(p)\leq-\tilde{\mu}_{1}\Bigl(\frac{m-1}{m}\Bigl)=-\dfrac{m-2}{m}\log(m-1)+\log m$
-due to (75). For $p(m)\leq p<1$ we have
+by (62). For $p(m)\leq p<1$ we have
 $\tilde{\mu}_{1}(p)\leq-\tilde{\mu}_{1}(p(m))=1,$
-because of (79). Therefore, $1+\tilde{\mu}_{1}(p)\geq 0$ for $p>\max(p(m),(m-1)/m)$. Note that $\tilde{\mu}_{1}(p)\geq\tilde{\mu}_{1}((m-1)/m)\geq\tilde{\mu}_{1}(p(m))=-1$ if p(m)<(m1)/mp(m)<(m-1)/m.
+by (66). Therefore, $1+\tilde{\mu}_{1}(p)\geq 0$ for $p>\max(p(m),(m-1)/m)$. Note that $\tilde{\mu}_{1}(p)\geq\tilde{\mu}_{1}((m-1)/m)\geq\tilde{\mu}_{1}(p(m))=-1$ if p(m)<(m1)/mp(m)<(m-1)/m.
 Case 2. The minimizer has no zero components.
-To reduce the number of different values to two, we will introduce a new Lagrange function $\hat{L}$. The critical points of $L$ are related with critical points of the Lagrange function
-$\hat{L}(s_{1},s_{2},s_{3},k_{1},k_{2},k_{3})=-\sum\limits_{j=1}^{3}k_{j}s_{j}^{2}\log s_{j}^{2}-\tilde{\mu}_{0}(p)\Bigl(\sum\limits_{j=1}^{m}k_{j}s_{j}\Bigl)^{2}$-
-$-\lambda\Bigl(\sum\limits_{j=1}^{3}k_{j}s_{j}^{2}-1\Bigl)-\nu\Bigl(\sum\limits_{j=1}^{3}k_{j}-m\Bigl).$
-New necessary conditions are
-$-\dfrac{1}{2}\dfrac{\partial\hat{L}}{\partial s_{l}}=k_{l}s_{l}\log s_{l}^{2}+\tilde{\mu}_{0}(p)k_{l}\Bigl(\sum\limits_{j=1}^{3}k_{j}s_{j}\Bigl)+(\lambda+\log e)k_{l}s_{l}=0,$
-$-\dfrac{\partial\hat{L}}{\partial k_{l}}=s_{l}^{2}\log s_{l}^{2}+2\tilde{\mu}_{0}(p)s_{l}\Bigl(\sum\limits_{j=1}^{3}k_{j}s_{j}\Bigl)+\lambda s_{l}^{2}+\nu=0$
-for all $1\leq l\leq 3$.
-Thus, we have
-$\dfrac{1}{2}s_{l}\dfrac{\partial\hat{L}}{\partial s_{l}}+k_{l}\dfrac{\partial\hat{L}}{\partial k_{l}}=k_{l}s_{l}\Bigl(s_{l}\log e-\tilde{\mu}_{0}(p)\sum\limits_{j=1}^{3}k_{j}s_{j}\Bigl)-\nu k_{l}=0.$
-Case 2.1. $k_{1},k_{2},k_{3}\neq 0$.
-For all $1\leq l\leq 3$ it holds $s_{l}^{2}\log e=\tilde{\mu}_{0}(p)\Bigl(\sum\limits_{j=1}^{3}k_{j}s_{j}\Bigl)+\nu$, which can take place, if the set $\{s_{1},s_{2},s_{3}\}$ has no more than two different values (which is realized, formally, for fractional multiplicities).
-This can be regarded as the next case.
-Case 2.2. $k_{3}=0$.
-Up to permutation and the global sign, $(z_{j})_{j=1}^{m}=(\underbrace{s_{1},...,s_{1}}_{k_{1}\text{ times}},\underbrace{s_{2},...,s_{2}}_{k_{2}\text{ times}})$, $s_{1}>0$}, $s_{2}<0$}.
-The components of the minimizer should have different signs; otherwise, changing the sign of a single component would yield a lower value for $\tilde{F}$.
+The Lagrange multipliers method predicts that the minimizer has at most 3 different values, that are in our case nonzero.
+They are denoted as $s_{1}<0<s_{2}\leq s_{3}$ (without loss of generality we can assume that only $s_{1}$ is negative) with multipliers $k_{1},k_{2},k_{3}$, such that $k_{1}+k_{2}+k_{3}=m$ and $k_{1}s_{1}^{2}+k_{2}s_{2}^{2}+k_{3}s_{3}^{2}=1$.
+Our hypothesis is that, however, there is only two distinct nonzero values among $s_{1},s_{2},s_{3}$.
+This fact can be verified by computer experiments; we performed numerical simulations for $m$ from $3$ to $50$ and for the parameter $a=\tilde{\mu}_{0}(p)>0$ taking values in $\{0.1,0.5,1,5,10,100,\infty\}$ (see Appendix).
+Thus, up to permutation and an overall sign change, $(z_{j})_{j=1}^{m}=(\underbrace{s_{1},...,s_{1}}_{k_{1}\text{ times}},\underbrace{s_{2},...,s_{2}}_{k_{2}\text{ times}})$, $s_{1}<0$}, $s_{2}>0$}.
+The components of the minimizer must have opposite signs; otherwise, changing the sign of a single component would decrease $\tilde{F}$.
 Denote
@@
 The initial inequality (59) for $k_{1}s_{1}^{2}=t$ is equivalent to
-$t\log\dfrac{t}{k_{1}/m}-(1-t)\log\dfrac{1-t}{k_{2}/m}\geq\tilde{\mu}_{0}(p,1/m)\Bigl(\sqrt{t(k_{1}/m)}-\sqrt{(1-t)(k_{2}/m)}\Bigl)^{2}-\tilde{\mu}_{1}(p,1/m),$ (94)
-and the statement follows from lemma 5.
+$t\log\dfrac{t}{k_{1}/m}-(1-t)\log\dfrac{1-t}{k_{2}/m}\geq\tilde{\mu}_{0}(p,1/m)\Bigl(\sqrt{t(k_{1}/m)}-\sqrt{(1-t)(k_{2}/m)}\Bigl)^{2}-\tilde{\mu}_{1}(p,1/m),$
+and the result follows from lemma 5.
 Lemma 5.
@@
 ∎
-Acknowledgment. The authors are grateful to S.V. Grishin and D.A. Kronberg for providing numerical simulations.
+Appendix. Numerical Verification of Hypotheses for Minimization on the Sphere (Generated by DeepSeek)
+We consider the minimization problem of the function
+$\tilde{F}(z_{1},\dots,z_{m})=-\sum_{k=1}^{m}z_{k}^{2}\log_{2}(z_{k}^{2})+a\left(\sum_{k=1}^{m}z_{k}\right)^{2}$ (80)
+on the unit sphere $\sum_{k=1}^{m}z_{k}^{2}=1$, where the parameter $a>0$ equals to $\tilde{\mu}_{0}(p)$ given by (60) in the obtuse scenario; in the flat scenario (formally corresponding to the case $a=\infty$) the second term in the righthand side of (80) is absent and there is additional constraint $\sum_{k=1}^{m}z_{k}=0$.
+The main hypothesis suggests that minimizers have at most two distinct nonzero values. To verify this hypothesis numerically, we implement the following verification scheme:
+Algorithm 1 Numerical verification of the minimization hypothesis
+1:Input: Parameter $a$, dimension range $m_{\min},m_{\max}]$, tolerance $\epsilon=10^{-8}$
+2:Output: Verification result for each dimension $m$
+3:procedure VerifyHypothesis($a$, $_{\min}, mm$$_{\max}$)
+4:  for $m=m_{\min}$ to $m_{\max}$ do
+5:   Find approximate minimizer $\mathbf{z}^{*}$ on the sphere
+6:   Compute $\tilde{F}_{\min}=\tilde{F}(\mathbf{z}^{*})$
+7:   $\triangleright$ Test Hypothesis 1: Two-value structure
+8:   $\mathbf{z}_{\text{proj}}\leftarrow\text{ProjectToTwoValueStructure}(\mathbf{z}^{*},a)$
+9:   if $\mathbf{z}_{\text{proj}}\neq\text{None}$ then
+10:     $\Delta\tilde{F}_{1}\leftarrow\tilde{F}(\mathbf{z}_{\text{proj}})-\tilde{F}_{\min}$
+11:     $\text{valid}_{1}\leftarrow(|\Delta\tilde{F}_{1}|\leq\epsilon)$
+12:   else
+13:     $\text{valid}_{1}\leftarrow\text{False}$
+14:   end if
+15:   $\triangleright$ Test Hypothesis 2: Special two-component vector
+16:   $\mathbf{z}_{\text{special}}\leftarrow(1/\sqrt{2},0,\dots,0,-1/\sqrt{2})$
+17:   $\Delta\tilde{F}_{2}\leftarrow\tilde{F}(\mathbf{z}_{\text{special}})-\tilde{F}_{\min}$
+18:   $\text{valid}_{2}\leftarrow(|\Delta\tilde{F}_{2}|\leq\epsilon)$
+19:   $\triangleright$ Accept hypothesis if either projection is nearly optimal
+20:   $\text{confirmed}\leftarrow\text{valid}_{1}\ \textbf{or}\ \text{valid}_{2}$
+21:   Output result for dimension $m$
+22:  end for
+23:end procedure
+For a given vector $\mathbf{z}^{*}$, we compute the averages $u$ and $v$ of its positive and negative components, respectively, then rescale them to lie on the unit sphere.
+In the case $a=\infty$, the exact values $u=\sqrt{\ell/(km)}$, $v=-\sqrt{k/(\ell m)}$ are used to satisfy both $\|\mathbf{z}\|=1$ and $\sum_{i}z_{i}=0$.
+The projection replaces all positive entries by $u$ and all negative entries by $v$, preserving zeros.
+Verification Criterion
+The hypothesis is considered confirmed for dimension $m$ if either:
+  • The projection $\mathbf{z}_{\text{proj}}$ onto a two-value structure yields a function value within tolerance $\epsilon$ of the computed minimum: $|\tilde{F}(\mathbf{z}_{\text{proj}})-\tilde{F}_{\min}|\leq\epsilon$
+  • The special two-component vector $\mathbf{z}_{\text{special}}=(1/\sqrt{2},0,\dots,0,-1/\sqrt{2})$ yields a function value within tolerance $\epsilon$ of the computed minimum
+We use $\epsilon=10^{-8}$ throughout our computations. The optimization is performed using SciPy’s SLSQP algorithm with multiple initial points to ensure robustness.
+VERIFICATION OF THE MINIMIZATION HYPOTHESIS FOR FUNCTION F ON THE SPHERE
+Enter parameter a: 10
+Enter m_min: 3
+Enter m_max: 50
+Parameters: a = 10, m in [3, 50]
+Criterion: hypothesis holds if $\Delta\tilde{F}$_1 < 1e-8 or $\Delta\tilde{F}$_2 < 1e-8
+   m |    $\tilde{F}$_min    |    $\Delta\tilde{F}$_1     |    $\Delta\tilde{F}$_2     | Result
+   3 |   1.00000000 |     2.52e-01 |    -9.44e-15 |    YES
+   4 |   1.00000000 |     2.08e-01 |    -4.11e-15 |    YES
+   5 |   1.00000000 |     1.22e-01 |    -4.33e-15 |    YES
+   6 |   0.98734793 |    -1.44e-15 |     1.27e-02 |    YES
+   7 |   0.91479512 |     0.00e+00 |     8.52e-02 |    YES
+   8 |   0.85213438 |     1.11e-16 |     1.48e-01 |    YES
+   9 |   0.79788958 |    -8.88e-16 |     2.02e-01 |    YES
+  10 |   0.75061687 |     1.11e-16 |     2.49e-01 |    YES
+  11 |   0.70910123 |    -3.33e-16 |     2.91e-01 |    YES
+  12 |   0.67236219 |    -3.33e-16 |     3.28e-01 |    YES
+  13 |   0.63961653 |    -3.33e-16 |     3.60e-01 |    YES
+  14 |   0.61023775 |     0.00e+00 |     3.90e-01 |    YES
+  15 |   0.58372158 |    -2.22e-16 |     4.16e-01 |    YES
+  16 |   0.55965888 |    -2.22e-16 |     4.40e-01 |    YES
+  17 |   0.53771482 |    -1.67e-15 |     4.62e-01 |    YES
+  18 |   0.51761305 |     2.22e-16 |     4.82e-01 |    YES
+  19 |   0.49912361 |     5.55e-17 |     5.01e-01 |    YES
+  20 |   0.48205361 |    -1.11e-16 |     5.18e-01 |    YES
+  21 |   0.46624010 |    -5.55e-17 |     5.34e-01 |    YES
+  22 |   0.45154446 |    -7.77e-15 |     5.48e-01 |    YES
+  23 |   0.43784802 |    -2.78e-16 |     5.62e-01 |    YES
+  24 |   0.42504857 |    -2.50e-15 |     5.75e-01 |    YES
+  25 |   0.41305761 |    -2.78e-16 |     5.87e-01 |    YES
+  26 |   0.40179807 |    -8.88e-16 |     5.98e-01 |    YES
+  27 |   0.39120256 |    -3.33e-16 |     6.09e-01 |    YES
+  28 |   0.38121185 |    -2.22e-16 |     6.19e-01 |    YES
+  29 |   0.37177368 |    -2.33e-15 |     6.28e-01 |    YES
+  30 |   0.36284175 |     1.11e-16 |     6.37e-01 |    YES
+  31 |   0.35437487 |     3.89e-16 |     6.46e-01 |    YES
+  32 |   0.34633629 |    -2.78e-16 |     6.54e-01 |    YES
+  33 |   0.33869310 |    -1.11e-16 |     6.61e-01 |    YES
+  34 |   0.33141576 |    -5.55e-17 |     6.69e-01 |    YES
+  35 |   0.32447761 |    -2.05e-15 |     6.76e-01 |    YES
+  36 |   0.31785461 |     3.33e-16 |     6.82e-01 |    YES
+  37 |   0.31152497 |    -2.66e-15 |     6.88e-01 |    YES
+  38 |   0.30546889 |    -7.61e-15 |     6.95e-01 |    YES
+  39 |   0.29966835 |    -3.33e-16 |     7.00e-01 |    YES
+  40 |   0.29410691 |    -1.59e-14 |     7.06e-01 |    YES
+  41 |   0.28876954 |    -6.11e-16 |     7.11e-01 |    YES
+  42 |   0.28364247 |    -2.61e-15 |     7.16e-01 |    YES
+  43 |   0.27871305 |    -1.67e-16 |     7.21e-01 |    YES
+  44 |   0.27396965 |    -8.60e-15 |     7.26e-01 |    YES
+  45 |   0.26940154 |    -8.60e-15 |     7.31e-01 |    YES
+  46 |   0.26499883 |     3.33e-16 |     7.35e-01 |    YES
+  47 |   0.26075237 |     2.78e-16 |     7.39e-01 |    YES
+  48 |   0.25665367 |    -4.00e-15 |     7.43e-01 |    YES
+  49 |   0.25269489 |     5.55e-16 |     7.47e-01 |    YES
+  50 |   0.24886871 |    -1.67e-16 |     7.51e-01 |    YES
+FINAL RESULTS:
+ YES: Hypothesis CONFIRMED for all 48 values of m
+The computational results reveal that for large $a$ and small $m$ (<7<7), the minimizer typically contains only two nonzero entries of opposite sign.
+Otherwise, all components are nonzero and assume exactly two distinct values.